### v1

# A robust parameterized enhanced shift-splitting preconditioner for three-by-three block saddle point problems

Abstract
This paper proposes a new parameterized enhanced shift-splitting (PESS) preconditioner to solve the three-by-three block saddle point problem (SPP). In addition, necessary and sufficient criteria are established for the convergence of the proposed PESS iterative process for any random initial guess. Furthermore, we meticulously investigate the spectral bounds of the PESS preconditioned matrix. Moreover, empirical investigation has been performed for the sensitivity analysis of the system, revealing the robustness of the proposed PESS preconditioner. Numerical experiments are carried out to demonstrate the enhanced efficiency and robustness of the proposed PESS preconditioner compared to the existing block diagonal and shift-splitting preconditioners.
Keywords. Krylov subspace iterative process $\cdot$ Preconditioner $\cdot$ Shift-splitting $\cdot$ Saddle point problem $\cdot$ Robust AMS subject classification. 65F08 $\cdot$ 65F10 $\cdot$ 65F50
1 Introduction
Three-by-three block saddle point problems (SPPs) have recently received considerable attention due to their wide applicability in various computational research and technological applications. These domains include the constraint optimization problems [16], computational fluid dynamics [8, 14], the constrained least square estimations [6], optimal control [20], solving Maxwell equations [12], magma-mantle dynamics [21] and so on. The three-by-three block SPP is represented as follows [18]:
$\mathcal{A}{\bf u}={\bf d},$ (1.1)
where $\mathcal{A}=\left[\begin{matrix}A&B^{T}&0\\ -B&0&-C^{T}\\ 0&C&0\end{matrix}\right],\,{\bf u}=\left[\begin{matrix}x\\ y\\ z\end{matrix}\right],\,{\bf d}=\left[\begin{matrix}f\\ g\\ h\end{matrix}\right],$ and $A\in{\mathbb{R}}^{n\times n}$ is a symmetric positive definite (SPD) matrix, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are the full row rank matrices. Here, $B^{T}$ stands for the transpose of $B,$ and $f\in{\mathbb{R}}^{n},\,g\in{\mathbb{R}}^{m}$ and $h\in{\mathbb{R}}^{p}$ are known vectors. Under the stated assumptions on block matrices A, B, and C, $\mathcal{A}$ is nonsingular, which ensures the existence of a unique solution for the system (1.1). Due to the growing popularity of three-by-three block SPPs in a variety of applications, it is essential to determine the efficient and robust numerical solution of three-by-three block SPPs (1.1). However, the indefiniteness and poor spectral properties of $\mathcal{A}$ makes challenging to determine the solution of the system (1.1). Furthermore, $\mathcal{A}$ is generally sparse and large; therefore, iterative processes become superior to direct methods. The Krylov subspace iterative process is preferable among other iterative processes investigated recently in the literature due to their minimal storage requirement and feasible implementation [29]. Slow convergence of the Krylov subspace iterative process is a major drawback for system of linear equations of large dimensions. Moreover, the saddle point matrix $\mathcal{A}$ can be very sensitive, which leads to a significant slowdown in the solution algorithm. Therefore, it is important to develop novel, efficient, and robust preconditioners for the fast convergence of the Krylov subspace iterative process, which can handle the sensitivity of the problem (1.1) when the matrix $\mathcal{A}$ is ill-conditioned. Consider the following partitioning of $\mathcal{A}:$
$\left[\begin{array}[]{cc|c}A&B^{T}&0\\ -B&0&-C^{T}\\ \hline\cr 0&C&\hfil 0\hfil\\ \end{array}\right],\qquad\left[\begin{array}[]{c|cc}A&B^{T}&0\\ \hline\cr-B&0&-C^{T}\\ 0&C&0\\ \end{array}\right],$ (1.8)
then the linear system (1.1) can be identified as the standard two-by-two block SPP [5]. In the last decades, colossal attention has been given to the development of numerical solution methods for standard two-by-two block SPP, such as Uzawa methods [7, 13, 2], Krylov subspace iterative process [17, 26], null space methods [15, 25]. However, the properties of the submatrices in (1.8) are different from the standard two-by-two block SPP. Notice that the leading blocks are skew-symmetric and positive semi-definite in the first partitioning. The second partitioning highlights that the $(1,2)$ block is rank deficient, and the $(2,2)$ block is skew-symmetric. In contrast, in the standard two-by-two block SPP, $(1,2)$ block is full row rank, and the $(2,2)$ block is generally symmetric positive semidefinite. Therefore, the existing literature for solving the two-by-two SPPs (see [5]) may not be applied to solve (1.1) directly, and it is essential to develop new preconditioners for three-by-three block SPP that exploit the specific structure of the saddle point matrix $\mathcal{A}$, which is sensitive in nature.
Recently, various effective preconditioners have been developed in the literature for solving three-by-three block SPPs. For instance, Huang and Ma [18] has studied the exact block diagonal (BD) and inexact BD (IBD) preconditioners ${\mathscr{P}}_{BD}$ and ${\mathscr{P}}_{IBD}$ in the following forms:
${\mathscr{P}}_{BD}=\left[\begin{matrix}A&0&0\\ 0&S&0\\ 0&0&CS^{-1}C^{T}\end{matrix}\right]\quad\mbox{and}\quad{\mathscr{P}}_{IBD}=\left[\begin{matrix}\hat{A}&0&0\\ 0&\hat{S}&0\\ 0&0&C\hat{S}^{-1}C^{T}\end{matrix}\right],$ (1.15)
where $S=BA^{-1}B^{T},$ $\hat{A}$ and $\hat{S}$ are SPD approximations of $A$ and $S,$ respectively. Although the spectrum of the preconditioned matrices ${\mathscr{P}}^{-1}_{BD}\mathcal{A}$ and ${\mathscr{P}}^{-1}_{IBD}\mathcal{A}$ have good clustering properties, these preconditioners have certain shortcomings, such as they are time-consuming, expensive, require excessive number of iterations, and condition numbers are very large. For more details, see the paper [18]. To overcome these drawbacks of BD preconditioners, researchers have started exploring shift-splitting (SS) preconditioners, which was initially developed for a non-Hermitian system of linear equations by Bai et al. [3] and later for two-by-two block SPPs [10, 23, 11]. Recently, Cao [9] enhanced this idea and introduced the following SS preconditioner ${\mathscr{P}}_{SS}$ and relaxed SS (RSS) preconditioner ${\mathscr{P}}_{RSS}$ for the three-by-three block SPP (1.1)
${\mathscr{P}}_{SS}=\frac{1}{2}\left[\begin{matrix}\alpha I+A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right],\,\,\mbox{and}\,\,{\mathscr{P}}_{RSS}=\frac{1}{2}\left[\begin{matrix}A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right],$ (1.16)
where $\alpha>0$ and verified unconditionally convergence of the associated SS iterative process. Here, $I$ stands for the identity matrix of the appropriate dimension. Wang and Zhang [27] generalized the SS preconditioner by introducing a new parameter $\beta>0$ in the $(3,3).$ Further, Yin et al. [30] developed an extensive generalized SS (EGSS) preconditioner
${\mathscr{P}}_{EGSS}=\frac{1}{2}\left[\begin{matrix}\alpha P+A&B^{T}&0\\ -B&\beta Q&-C^{T}\\ 0&C&\gamma W\end{matrix}\right],$ (1.17)
for the three-by-three block SPP (1.1), where $\alpha,\beta,\gamma>0$ and $P,Q,W$ are SPD matrices and investigated its convergent properties. There are a number of other preconditioners and iterative processes to solve the three-by-three block SPP (1.1) has been studied recently in the literature  [19, 1, 24, 28, 4]. Despite exhibiting favorable performance of SS, RSS and EGSS preconditioners, they do not outperform BD and IBD preconditioners. Therefore, to improve the convergence speed and efficiency of the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS}$ and ${\mathscr{P}}_{EGSS},$ this paper presents a parameterized enhanced shift-splitting (PESS) preconditioner by introducing a parameter $s>0$ for there-by-three block SPPs. As per the direct correlation between the rate of convergence in Krylov subspace iterative processes and the spectral properties of the preconditioned matrix, we perform in-depth spectral distribution of PESS preconditioned matrix. We summarize the following main contributions of this paper:
  • We propose the PESS iterative process and associated PESS preconditioner to solve the three-by-three block SPP (1.1). The proposed preconditioner generalizes the other SS preconditioners studied in the literature see, for example, [9, 27, 30].
  • General framework is given on necessary and sufficient criteria for the convergence of the PESS iterative process. These investigations also encompass the unconditional convergence of other exiting SS preconditioners in the literature [9, 30].
  • We have conducted a comprehensive analysis of the spectral distribution for the PESS preconditioned matrix. Our framework allows us to derive the spectral distribution for the SS and EGSS preconditioned matrices.
  • We empirically show that our proposed PESS preconditioner significantly reduces the condition number of the ill-conditioned saddle point matrix $\mathcal{A}$. Thereby establishing an efficiently solvable, well-conditioned system.
  • Numerical experiments show that the proposed PESS preconditioner outperforms the existing BD and SS preconditioners. Moreover, we perform sensitivity analysis by perturbing the ill-conditioned matrix $\mathcal{A}$ in (1.1), demonstrating the robustness of the proposed preconditioner.
The structure of this article is as follows. In Section 2, we propose the PESS iterative process and the associated PESS preconditioner. Section 3 is devoted to investigating the convergence of the PESS iterative process. In Section 4, we investigate the spectral distribution of the PESS preconditioned matrix. In Section 5, we conduct numerical experiments to illustrate the computational efficiency and robustness of the proposed preconditioner. In the end, a brief concluding remark is provided in Section 6. Notation. The following notations are adopted throughout the paper. For any matrix $A\in{\mathbb{R}}^{n\times n},$ $A^{T},\sigma(A)$ and $\vartheta(A)$ denotes transpose, spectrum and spectral radius of $A,$ respectively. For any $x\in\mathbb{C}^{n},$ $x^{*}$ denotes its conjugate transpose. The symbol $\mathrm{diag}(A,B)$ represents a block diagonal matrix with diagonal blocks $A$ and $B.$ The real and imaginary parts of any complex number $z$ are represented by $\Re(z)$ and $\Im(z)$, respectively. For any matrix $A\in{\mathbb{R}}^{n\times n},$ $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$ denotes the maximum and minimum eigenvalue of $A,$ respectively. The notation $A>0$ signifies that $A$ is SPD, and $A>B$ identifies $A-B>0.$ Abbreviations used in the paper are as follows:
SPP: saddle point problem, SPD: symmetric positive definite, BD: block diagonal, IBD: inexact block diagonal, SS: shift-splitting, RSS: relaxed shift-splitting, EGSS: extensive generalized shift-spliting, PESS: parameterized enhanced shift-splitting, GMRES: generalized minimum residual process.
2 The parameterized enhanced shift-splitting (PESS) iterative process and the preconditioner
Let $s>0$ be a real number. Then, we write the matrix $\mathcal{A}$ in the form
$\mathcal{A}=(\Sigma+s\mathcal{A})-(\Sigma-(1-s)\mathcal{A}):={\mathscr{P}}_{PESS}-\mathcal{Q}_{PESS},\,\mbox{where}$ (2.1)
${\mathscr{P}}_{PESS}=\left[\begin{matrix}\Lambda_{1}+sA&sB^{T}&0\\ -sB&\Lambda_{2}&-sC^{T}\\ 0&sC&\Lambda_{3}\end{matrix}\right],\mathcal{Q}_{PESS}:=\left[\begin{matrix}\Lambda_{1}-(1-s)A&-(1-s)B^{T}&0\\ (1-s)B&\Lambda_{2}&(1-s)C^{T}\\ 0&-(1-s)C&\Lambda_{3}\end{matrix}\right],$
$\Sigma=\mathrm{diag}(\Lambda_{1},\Lambda_{2},\Lambda_{3}),$ and $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ are SPD matrices. The matrix ${\mathscr{P}}_{PESS}$ is nonsingular for $s>0.$ The special matrix splitting (2.1) leads us to the subsequent iterative process for solving the three-by-three block SPP given in (1.1): The PESS iterative process. Let $s$ be a positive constant, and let $\Lambda_{1}\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}$ $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices. For any initial guess ${\bf u}_{0}\in{\mathbb{R}}^{n+m+p}$ and $k=0,1,2\ldots,$ until a specified termination criterion is fulfilled, we compute
${\bf u}_{k+1}=\mathcal{T}{\bf u}_{k}+{\bf c},$ (2.2)
where ${\bf u}_{k}=[x_{k}^{T},y_{k}^{T},z_{k}^{T}]^{T},{\bf c}={\mathscr{P}}_{PESS}^{-1}{\bf d},$ and $\mathcal{T}={\mathscr{P}}_{PESS}^{-1}\mathcal{Q}_{PESS}$ is called the iteration matrix for the PESS iterative process. Moreover, the matrix splitting (2.2) introduces a new preconditioner ${\mathscr{P}}_{PESS}$, identified as PESS preconditioner. This preconditioner is a generalization of the previous work in the literature, see for example [9, 27, 30] for specific choices of $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ and $s,$ which are listed in Table 1.
Table 1: ${\mathscr{P}}_{PESS}$ as a generalization of the above SS preconditioners for different choices of $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ and $s.$
${\mathscr{P}}_{PESS}$ $\Lambda_{1}$ $\Lambda_{2}$ $\Lambda_{3}$ $s$ memo
${\mathscr{P}}_{SS}$ [9] $\frac{1}{2}\alpha I$ $\frac{1}{2}\alpha I$ $\frac{1}{2}\alpha I$ $s=\frac{1}{2}$ $\alpha>0$
${\mathscr{P}}_{GSS}$ [27] $\frac{1}{2}\alpha I$ $\frac{1}{2}\alpha I$ $\frac{1}{2}\beta I$ $s=\frac{1}{2}$ $\alpha,\beta>0$
${\mathscr{P}}_{EGSS}$ [30] $\frac{1}{2}\alpha P$ $\frac{1}{2}\beta Q$ $\frac{1}{2}\gamma W$ $s=\frac{1}{2}$ $P,Q,W$ are SPD
matrices and $\alpha,\beta,\gamma>0$
In the implementation of the PESS iterative process or the PESS preconditioner to enhance the rate of convergence of the Krylov subspace iterative process like GMRES, at each iterative step, we solve the following system of linear equations
${\mathscr{P}}_{PESS}w=r,$ (2.3)
where $r=[r_{1}^{T},r_{2}^{T},r_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}$ and $w=[w_{1}^{T},w_{2}^{T},w_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}.$ Specifying $\hat{X}=\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C$ and $\hat{A}=\Lambda_{1}+sA+s^{2}B^{T}\hat{X}^{-1}B,$ then ${\mathscr{P}}_{PESS}$ can be written in the following way
${\mathscr{P}}_{PESS}=\left[\begin{matrix}I&sB^{T}\hat{X}^{-1}&0\\ 0&I&-sC^{T}\Lambda_{3}^{-1}\\ 0&0&I\end{matrix}\right]\left[\begin{matrix}\hat{A}&0&0\\ 0&\hat{X}&0\\ 0&0&\Lambda_{3}\end{matrix}\right]\left[\begin{matrix}I&0&0\\ -s\hat{X}^{-1}B&I&0\\ 0&s\Lambda_{3}^{-1}C&I\end{matrix}\right].$ (2.13)
The decomposition (2.13) leads us to the following Algorithm 1 to determine the solution of the system (2.3).
Algorithm 1 Computation of $w$ from ${\mathscr{P}}_{PESS}w=r$
1:Solve $(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)v_{1}=r_{2}+sC^{T}\Lambda_{3}^{-1}r_{3}$ to find $v_{1};$
2:Compute $v=r_{1}-sB^{T}v_{1};$
3:Solve $(\Lambda_{1}+sA+s^{2}B^{T}(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)^{-1}B)w_{1}=v$ to find $w_{1}$
4:Solve $(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)v_{2}=sBw_{1}$ to find $v_{2}$;
5:Compute $w_{2}=v_{1}+v_{2};$
6:Solve $\Lambda_{3}w_{3}=r_{3}-sCw_{2}$ for $w_{3}.$
The implementation of Algorithm 1 requires to solve two linear subsystems with coefficient matrices $\hat{X}=\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C$ and $\hat{A}=\Lambda_{1}+sA+s^{2}B^{T}(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)^{-1}B$. Since $\Lambda_{1},\Lambda_{2},\Lambda_{3},\hat{X},$ and $\hat{A}$ are SPD and $s>0$, we can apply the Cholesky factorization to solve these linear subsystems exactly or inexactly by the preconditioned conjugate gradient method. Similarly, we can solve step $6$ of the Algorithm 1.
3 Convergence analysis of the PESS iterative process
In this section, we investigate the convergence of the PESS iterative process (2.2). To achieve this aim, the following definition and lemmas are crucial.
Definition 3.1.
A matrix $\mathcal{A}$ is called positive stable if $\Re(\lambda)>0$ for all $\lambda\in\sigma(\mathcal{A})$.
Lemma 3.1.
[9] Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Then, the saddle point matrix $\mathcal{A}$ is positive stable.
Lemma 3.2.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Then $\Sigma^{-1}\mathcal{A}$ (normal-(((or, $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})$ is positive stable.
Proof.
Since $\Sigma^{-\frac{1}{2}}=\mathrm{diag}(\Lambda_{1}^{-\frac{1}{2}},\Lambda_{2}^{-\frac{1}{2}},\Lambda_{3}^{-\frac{1}{2}}),$ we obtain
$\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}=\left[\begin{matrix}\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}&\Lambda_{1}^{-\frac{1}{2}}B^{T}\Lambda_{2}^{-\frac{1}{2}}&0\\ -\Lambda_{2}^{-\frac{1}{2}}B\Lambda_{1}^{-\frac{1}{2}}&0&\Lambda_{2}^{-\frac{1}{2}}C^{T}\Lambda_{3}^{-\frac{1}{2}}\\ 0&\Lambda_{3}^{-\frac{1}{2}}C\Lambda_{2}^{-\frac{1}{2}}&0\end{matrix}\right].$ (3.4)
Since the matrices $\Lambda_{1},\Lambda_{2},$ and $\Lambda_{3}$ are SPD, then $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$ is SPD, and further $\Lambda_{2}^{-\frac{1}{2}}B\Lambda_{1}^{-\frac{1}{2}}$ and $\Lambda_{3}^{-\frac{1}{2}}C\Lambda_{2}^{-\frac{1}{2}}$ are of full row rank. Consequently, the block structure of $\mathcal{A}$ and the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ are identical. Then by Lemma 3.1, $\mathcal{A}$ is positive stable implies that, the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ and consequently, $\Sigma^{-1}\mathcal{A}$ is positive stable. $\blacksquare$ By [22], the stationary iterative process (2.2) converges to the exact solution of the three-by-three block SPP (1.1) for any initial guess vector if and only if $|\vartheta(\mathcal{T})|<1,$ where $\mathcal{T}$ is the iteration matrix. Now, we establish an if and only if condition that precisely determines the convergence of the PESS iterative process.
Theorem 3.1.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0$. Then, the PESS iterative process (2.2) converges to the unique solution of the three-by-three block SPP (1.1) if and only if
$(2s-1)|\mu|^{2}+2\Re(\mu)>0,\,\,\forall\,\mu\in\sigma(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}).$ (3.5)
Proof.
From (2.2) we have
$\mathcal{T}={\mathscr{P}}_{PESS}^{-1}\mathcal{Q}_{PESS}=(\Sigma+s\mathcal{A})^{-1}(\Sigma-(1-s)\mathcal{A}),$ (3.6)
where $\Sigma=\mathrm{diag}(\Lambda_{1},\Lambda_{2},\Lambda_{3}).$ Given that $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ are SPD, then
$\mathcal{T}=\Sigma^{-\frac{1}{2}}(I+s\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{-1}(I-(1-s)\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})\Sigma^{\frac{1}{2}}.$ (3.7)
Thus, (3.7) shows that the iteration matrix $\mathcal{T}$ is similar to $\widetilde{\mathcal{T}},$ where
$\widetilde{\mathcal{T}}=(I+s\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{-1}(I-(1-s)\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}).$
Let $\theta$ and $\mu$ be the eigenvalues of the matrices $\widetilde{\mathcal{T}}$ and $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}},$ respectively. Then, it is easy to show that
$\theta=\frac{1-(1-s)\mu}{1+s\mu}.$
Since $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ is nonsingular then $\mu\neq 0,$ and we have
$|\theta|=\frac{|1-(1-s)\mu|}{|1+s\mu|}=\sqrt{\frac{(1-(1-s)\Re(\mu))^{2}+(1-s)^{2}\Im(\mu)^{2}}{(1+s\Re(\mu))^{2}+s^{2}\Im(\mu)^{2}}}.$ (3.8)
From (3.8), it is clear that $|\theta|<1$ if and only if
$(1-(1-s)\Re(\mu))^{2}+(1-s)^{2}\Im(\mu)^{2}<(1+s\Re(\mu))^{2}+s^{2}\Im(\mu)^{2}.$
This implies $(2s-1)|\mu|^{2}+2\Re(\mu)^{2}>0.$ Since, $\vartheta(\mathcal{T})=\vartheta{\widetilde{(}\mathcal{T}}),$ the proof is conclusive. $\blacksquare$ Using the condition in Theorem 3.1, next, we present two sufficient conditions that ensure the convergence of the PESS iterative process. The first one is presented as follows:
Corollary 3.1.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and if $s\geq\frac{1}{2},$ then $\vartheta(\mathcal{T})<1,$ i.e., the PESS iterative process always converges to the unique solution of the three-by-three block SPP (1.1).
Proof.
By Lemma 3.2, $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ is positive stable, implies that $\Re(\mu)>0$ and $|\mu|>0.$ Thus, the inequality (3.5) holds if $s\geq\frac{1}{2}$. This completes the proof. $\blacksquare$
Remark 3.1.
By applying Corollary 3.1, the unconditional convergence of the existing preconditioners, namely ${\mathscr{P}}_{SS}$, ${\mathscr{P}}_{GSS},$ and ${\mathscr{P}}_{EGSS}$ can be obtained from ${\mathscr{P}}_{PESS}$ by sustituting $s=\frac{1}{2}$.
Next, we present a stronger sufficient condition to Corollary 3.1 for the convergence of the PESS iterative process.
Corollary 3.2.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, if
$s>\max\left\{\frac{1}{2}\left(1-\frac{\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}\right),0\right\},$
then the PESS iterative process is convergent for any random initial guess.
Proof.
Note that, as $\Re(\mu)>0,$ the condition in (3.5) holds if and only if $\frac{1}{2}-\frac{\Re(\mu)}{|\mu|^{2}}<s.$ Let ${\bf p}$ be an eigenvector corresponding to $\mu,$ then we have
$\Re(\mu)=\frac{{\bf p}^{*}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}){\bf p}}{2{\bf p}^{*}{\bf p}}\geq\frac{1}{2}\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}).$
Again, we have $|\mu|<\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}),$ then we get the following
$\frac{1}{2}-\frac{\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{2\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}>\frac{1}{2}-\frac{\Re(\mu)}{|\mu|^{2}}.$ (3.9)
If we apply $s>0$, then we get the desired result from (3.9). $\blacksquare$ The sufficient condition on Corollary 3.2 is difficult to find for large $\mathcal{A}$ due to the involvement of computation of $\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})$ and $\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}.$ Thus, we consider $s\geq\frac{1}{2}$ for practical implementation.
4 Spectral distribution of the PESS preconditioned matrix
Solving the three-by-three block SPP (1.1) is equivalent to solving the preconditioned system of linear equations
${\mathscr{P}}_{PESS}^{-1}\mathcal{A}{\bf u}={\mathscr{P}}_{PESS}^{-1}{\bf d},$ (4.1)
where ${\mathscr{P}}_{PESS}$ serves as a preconditioner for the preconditioned GMRES (PGMRES) process. Since preconditioned matrices with clustered spectrum frequently culminate in rapid convergence for GMRES, see, for example [5, 22], the spectral distribution of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ requires careful attention. As an immediate consequence of Corollary 3.1, the following clustering property for the eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ can be established.
Theorem 4.1.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are full row rank matrices. Suppose that $\lambda$ is an eigenvalue of the PESS preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$. Then, for $s\geq\frac{1}{2},$ we have
$|\lambda-1|<1,$
i.e., the spectrum of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ entirely contained in a disk with centered at $(1,0)$ and radius strictly smaller than $1$
Proof.
Assume that $\lambda$ and $\theta$ are the eigenvalue of the PESS preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and the iteration matrix $\mathcal{T}$, respectively. On the other hand, from the matrix splitting (2.1), we have
${\mathscr{P}}_{PESS}^{-1}\mathcal{A}={\mathscr{P}}_{PESS}^{-1}({\mathscr{P}}_{PESS}-\mathcal{Q}_{PESS})=I-\mathcal{T}.$
Then, we get $\lambda=1-\theta.$ Now, from Corollary 3.1, for $s\geq\frac{1}{2},$ it holds that $|\theta|<1.$ Hence, it follows that, for $s\geq\frac{1}{2},$ $|\lambda-1|<1,$ which completes the proof. $\blacksquare$ Let $(\lambda,{\bf p}=[u^{T},v^{T},w^{T}]^{T})$ be an eigenpair of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then, we have $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p},$ which can be written as
$Au+B^{T}v=\lambda(\Lambda_{1}+sA)u+s\lambda B^{T}v,$ (4.2a)
$-Bu-C^{T}w=-s\lambda Bu+\lambda\Lambda_{2}v-s\lambda C^{T}w,$ (4.2b)
$Cv=s\lambda Cv+\lambda\Lambda_{3}w.$ (4.2c)
Remark 4.1.
Notice that from (4.2a)-(4.2c), we get $\lambda\neq\frac{1}{s},$ otherwise , ${\bf p}=[u^{T},v^{T},w^{T}]^{T}={0}\in{\mathbb{R}}^{n+m+p},$ which is impossible as ${\bf p}$ is an eigenvector.
Proposition 4.1.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0.$ Let $(\lambda,{\bf p}=[u^{T},v^{T},w^{T}]^{T})$ be an eigenpair of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then, the following holds
  1. 1.
    $u\neq 0.$
  2. 2.
    When $\lambda$ is real, $\lambda>0.$
  3. 3.
    $\Re(\mu)>0,$ where $\mu=\lambda/(1-s\lambda).$
Proof.
  1. 1.
    Let $u=0$. Then by (4.2a), we get $(1-s\lambda)B^{T}v=0,$ which shows that $v=0,$ as $B$ has full row rank and $\lambda\neq\frac{1}{s}.$ Furthermore, when combined with (4.2c) leads to $w=0,$ as $\Lambda_{3}$ is SPD. Combining the above facts, it follows that ${\bf p}=[u^{T},v^{T},w^{T}]^{T}=0,$ which is impossible as ${\bf p}$ is an eigenvector. Hence, $u\neq 0.$
  2. 2.
    Let ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ be an eigenvector corresponding to the real eigenvalue $\lambda.$ Then, we have ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}{\bf p}=\lambda{\bf p},$ or, $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p}.$ Consequently, $\lambda$ can be written as
    $\lambda$ $=$ $\frac{{\bf p}^{T}\mathcal{A}{\bf p}+{\bf p}^{T}\mathcal{A}^{T}\bf p}{2({\bf p}^{T}{\mathscr{P}}_{PESS}{\bf p}+{\bf p}^{T}{\mathscr{P}}_{PESS}^{T}\bf p)}$
    $=$ $\frac{u^{T}Au}{2(su^{T}Au+u^{T}\Lambda_{1}u+v^{T}\Lambda_{2}v+w^{T}\Lambda_{3}w)}>0.$
    The last inequality follows from the assumptions that $A,\Lambda_{1},\Lambda_{2}$, and $\Lambda_{3}$ are SPD matrices.
  3. 3.
    The system of linear equations between (4.2a)-(4.2c) can be reformulated as
    $\left\{\begin{array}[]{lcl}Au+B^{T}v=(\lambda/(1-s\lambda))\Lambda_{1}u,&\\ -Bu-C^{T}w=(\lambda/(1-s\lambda))\Lambda_{2}v,&\\ Cv=(\lambda/(1-s\lambda))\Lambda_{3}w.\end{array}\right.$ (4.3)
    Now, the systems of linear equations in (4.3) can also be written as $\mathcal{A}{\bf p}=\mu\Sigma{\bf p},$ or, $\Sigma^{-1}\mathcal{A}{\bf p}=\mu{\bf p},$ where $\mu=\lambda/(1-s\lambda).$ By Lemma (3.2), we have $\Sigma^{-1}\mathcal{A}$ is positive stable, which implies $\Re(\mu)>0,$ hence the proof is completed. $\blacksquare$
Remark 4.2.
From Proposition 4.1, it follows that when $\lambda$ is real, then $\mu>0.$ Furthermore, the values of $\lambda$ lies in $(0,1/s)$ for all $s>0$.
In the following theorem, we provide sharper bounds for real eigenvalues of the PESS preconditioned matrix.
Theorem 4.2.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0.$ Suppose $\lambda$ is a real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then
$\lambda\in\left(0,\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}\right].$ (4.4)
Proof.
Let ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ be an eigenvector corresponding to the real eigenvalue $\lambda.$ Then, the system of linear equations in (4.2a)-(4.2c) are satisfied. First, we assume that $v=0.$ Subsequently, (4.2a) and (4.2c) are simplified to
$Au=\lambda(\Lambda_{1}+sA)u,\quad\lambda\Lambda_{3}w=0,$ (4.5)
respectively. The second equation of (4.5) gives $w=0,$ as $\Lambda_{3}$ is SPD. Multiplying $u^{T}$ to the first equation of (4.5) from the left, we get
$\lambda=\frac{u^{T}Au}{u^{T}\Lambda_{1}u+su^{T}Au}=\frac{p}{1+sp},$
where $p={\displaystyle{\frac{u^{T}Au}{u^{T}\Lambda_{1}u}}}>0.$ Now, for any nonzero vector $u\in{\mathbb{R}}^{n},$ we can have
$\lambda_{\min}(\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}})\leq\frac{u^{T}Au}{u^{T}\Lambda_{1}u}=\frac{(\Lambda_{1}^{\frac{1}{2}}u)^{T}\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}\Lambda_{1}^{\frac{1}{2}}u}{(\Lambda_{1}^{\frac{1}{2}}u)^{T}\Lambda_{1}^{\frac{1}{2}}u}\leq\lambda_{\max}(\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}).$ (4.6)
Since $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$ is similar to $\Lambda_{1}^{-1}A$ and the function $f:{\mathbb{R}}\rightarrow{\mathbb{R}}$ defined by $f(p):=p/(1+sp),$ where $s>0,$ is monotonically increasing in $p>0$, we obtain the following bound for $\lambda:$
$\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\min}(\Lambda_{1}^{-1}A)}\leq\lambda\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}.$ (4.7)
[…]
$v=-\frac{1}{\mu}\Lambda_{2}^{-1}Bu.$ (4.8)
Substituting (4.8) in (4.2a) yields
$Au=\lambda(\Lambda_{1}+sA)u+\frac{(1-s\lambda)}{\mu}B^{T}\Lambda_{2}^{-1}Bu.$ (4.9)
Premultiplying both sides of (4.9) by $u^{T},$ we obtain
$\lambda u^{T}Au=\lambda^{2}u^{T}\Lambda_{1}u+s\lambda^{2}u^{T}Au+(1-s\lambda)^{2}u^{T}B^{T}\Lambda_{2}^{-1}Bu,$ (4.10)
which is equivalent to
$\lambda^{2}-\lambda\frac{2sq+p}{s^{2}q+sp+1}+\frac{q}{s^{2}q+sp+1}=0,$ (4.11)
where $q={\displaystyle{\frac{u^{T}B^{T}\Lambda_{2}^{-1}Bu}{u^{T}\Lambda_{1}u}}}.$ By solving (4.11), we obtain the following real solutions for $\lambda$
$\lambda^{\pm}=\frac{2sq+p\pm\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)}$ (4.12)
where $p^{2}-4q\geq 0.$ Consider the functions $\Phi_{1},\Phi_{2}:{\mathbb{R}}\times{\mathbb{R}}\times{\mathbb{R}}\longrightarrow{\mathbb{R}}$ defined by
$\Phi_{1}(p,q,s)$ $=$ $\frac{2sq+p+\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)},$
$\Phi_{2}(p,q,s)$ $=$ $\frac{2sq+p-\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)},$
respectively, where $p,s>0,$ $q\geq 0$ and $p^{2}>4q$. Then $\Phi_{1}$ and $\Phi_{2}$ is monotonically increasing in the argument $p>0,$ and $q\geq 0,$ and decreasing in the argument $q\geq 0$ and $p>0,$ respectively. Now, using (4.6), (4.12) and monotonicity of the functions $\Phi_{1}$ and $\Phi_{2},$ we get the following bounds
$\lambda^{+}=\Phi_{1}(p,q,s)<\Phi_{1}(\lambda_{\max}(\Lambda^{-1}_{1}A),0,s),$ (4.13)
$\Phi_{2}(\lambda_{\max}(\Lambda^{-1}_{1}A),0,s)<\lambda^{-}=\Phi_{2}(p,q,s).$ (4.14)
Combining (4.13) and (4.14), we get the following bounds for $\lambda$
$0<\lambda^{\pm}<\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}.$ (4.15)
Next, consider the case $Cv\neq 0.$ Then from (4.2c), we get $w=\frac{1}{\mu}\Lambda_{3}^{-1}Cv.$ Substituting the value of $w$ in (4.2b), we obtain
$(1-s\lambda)Bu+\frac{(1-s\lambda)}{\mu}C^{T}\Lambda_{3}^{-1}Cv=-\lambda\Lambda_{2}v.$ (4.16)
Now, we assume that
$\lambda>\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}.$ (4.17)
Since by Remark 4.2, we get $01-s\lambda>$ 0 and the above assertion yields $\mu>\lambda_{\max}(\Lambda_{1}^{-1}A).$ Hence, the matrix $\mu\Lambda_{1}-A$ is nonsingular and (4.2a) gives
$u=\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v.$ (4.18)
Substituting (4.18) into (4.16), we obtain
$(1-s\lambda)B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v+\frac{(1-s\lambda)}{\mu}C^{T}\Lambda_{3}^{-1}Cv=-\lambda\Lambda_{2}v.$ (4.19)
Note that, $v\neq 0.$ Premultiplying by $v^{T}$ on the both sides of (4.19) leads to the following identity
$(1-s\lambda)v^{T}B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v+\lambda v^{T}\Lambda_{2}v=-\frac{(1-s\lambda)}{\mu}v^{T}C^{T}\Lambda_{3}^{-1}Cv.$ (4.20)
On the other hand, $\mu\Lambda_{1}-A>\Lambda_{1}(\mu-\ \lambda_{\max}(\Lambda^{-1}A))I>0,$ and hence a SPD matrix. Consequently, the matrix $B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}$ is also a SPD matrix and this implies $v^{T}B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v>0,$ for all $v\not=0.$ This leads to a contradiction to (4.20) as $v^{T}C^{T}\Lambda_{3}^{-1}Cv\geq 0.$ Hence, the assumption (4.17) is not true and we get
$\lambda\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}.$ (4.21)
Again, from (1) of Proposition 4.1, we have $\lambda>0.$ Combining the above together with (4.7) and (4.15), we obtained the desired bounds in (4.4) for $\lambda.$ Hence, the proof is concluded. $\blacksquare$ Since the preconditioners ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{EGSS}$ are special cases of the PESS preconditioner, next, we obtain refined bounds for real eigenvalues of the SS and EGSS preconditioned matrices from Theorem 4.2.
Corollary 4.1.
Let ${\mathscr{P}}_{SS}$ be defined as in (1.16) with $\alpha>0$ and let $\lambda$ be an real eigenvalue of the SS preconditioned matrix ${\mathscr{P}}_{SS}^{-1}\mathcal{A}.$ Then
$\lambda\in\left(0,\frac{2\lambda_{\max}(A)}{\alpha+\lambda_{\max}(A)}\right].$ (4.22)
Proof.
Since ${\mathscr{P}}_{SS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, the bounds in (4.22) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha I$ and $s=\frac{1}{2}$ in Theorem 4.2. $\blacksquare$ In the next result, we discuss bounds for the real eigenvalues of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$.
Corollary 4.2.
Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.17) with $\alpha>0$ and let $\lambda$ be an real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}.$ Then,
$\lambda\in\left(0,\frac{2\lambda_{\max}(P^{-1}A)}{\alpha+\lambda_{\max}(P^{-1}A)}\right].$ (4.23)
Proof.
Since ${\mathscr{P}}_{EGSS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, and the bounds are in (4.23) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha P$ and $s=\frac{1}{2}$ in Theorem 4.2. $\blacksquare$ The following result shows the bounds when $\lambda$ is a non-real eigenvalue.
Theorem 4.3.
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Let $s>0$ and $\lambda$ be a non-real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0.$ Suppose ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the eigenvector corresponding to $\lambda.$ Then the following holds $(1).$   If $Cv=0,$ then $\lambda$ satisfies
$\frac{\lambda_{\min}(\Lambda^{-1}A)}{2+s\lambda_{\min}(\Lambda_{1}^{-1}A)}\leq|\lambda|\leq\sqrt{\frac{\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}{1+s\lambda_{\min}(\Lambda_{1}^{-1}A)+s^{2}\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}}.$
$(2).$   If $Cv\neq 0,$ then real and imaginary part of $\lambda/(1-s\lambda)$ satisfies
$\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)\,\lambda_{\min}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}{2\left(\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)+\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)\right)}\leq\Re\left(\frac{\lambda}{1-s\lambda}\right)\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2},$
$\left|\Im\left(\frac{\lambda}{1-s\lambda}\right)\right|\leq\sqrt{\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}.$
Proof.
$(1).$   Let $\lambda$ be an eigenvalue of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0$ and the corresponding eigenvector is ${\bf p}=[u^{T},v^{T},w^{T}]^{T}.$ Then, the system of linear equations in (4.2a)- (4.2c) holds. We assert that, $v\neq 0,$ otherwise from (4.2a), we get $Au=\lambda(\Lambda_{1}+sA)u,$ which further implies $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}\bar{u}=\lambda\bar{u},$ where $\bar{u}=(\Lambda_{1}+sA)^{\frac{1}{2}}u.$ Since, $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}$ is a SPD matrix and by Lemma 4.1, $u\neq 0,$ and hence $\bar{u}\neq 0,$ which implies $\lambda$ is real, thus a contradiction to $\Im(\lambda)\neq 0.$ Initially, we consider the case $Cv=0.$ The aforementioned discussion suggests that $Au\neq\lambda(\Lambda_{1}+sA)u$ for any $u\neq 0.$ Consequently, $(\mu\Lambda_{1}-A)$ is nonsingular. Again, since $Cv=0,$ from (4.2c), and similar to (4.8)-(4.10), we obtain quadratic equation (4.11) in $\lambda$, which has complex solutions
$\lambda^{\pm i}=\frac{2sq+p\pm i\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)}$ (4.24)
for $4q>p^{2}.$ From (4.24), we get
$|\lambda^{\pm i}|^{2}=\frac{q}{s^{2}q+sp+1}.$ (4.25)
Now, consider the function $\Psi_{1}:{\mathbb{R}}\times{\mathbb{R}}\times{\mathbb{R}}\longrightarrow{\mathbb{R}}$ defined by
$\Psi_{1}(p,q,s)=\frac{q}{s^{2}q+sp+1},$
where $p,q,s>0.$ Then $\Psi_{1}$ is monotonically increasing in the argument $q>0$ while decreasing in the argument $p>0.$ Due to the fact that $q\leq\lambda_{\max}(\Lambda_{1}^{-1}B^{T}\Lambda_{2}^{-1}B)=\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T}),$ we get the following bounds
$|\lambda^{\pm i}|^{2}$ $=$ $\Psi_{1}(p,q,s)\leq\Psi_{1}(\lambda_{\min}(\Lambda_{1}^{-1}A),\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T}),s)$ (4.26)
$=$ $\frac{\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}{1+s\lambda_{\min}(\Lambda_{1}^{-1}A)+s^{2}\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}.$
Again, considering the condition $q>(p^{2}/4),$ (4.25) adheres to the inequality $(p/(2+sp))\leq|\lambda^{\pm i}|.$ Notably, the function $\Psi_{2}:{\mathbb{R}}\longrightarrow{\mathbb{R}}$ defined by $\Psi_{2}(p)=p/(2+sp),$ where $p,s>0,$ is monotonically increasing in the argument $p>0,$ we obtain the following bound
$|\lambda^{\pm i}|=\Psi_{2}(p)\geq\Psi_{2}(\lambda_{\min}(\Lambda_{1}^{-1}A))=\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)}{2+s\lambda_{\min}(\Lambda_{1}^{-1}A)}.$ (4.27)
Hence, combining (4.26) and (4.27), proof for the part $(1).$ follows.
$(2).$   Next, consider the case $Cv\neq 0,$ then (4.2c) yields $w=\frac{1}{\mu}\Lambda_{3}^{-1}Cv,$ where $\mu=\lambda/(1-s\lambda).$ Substituting this in (4.2b), we get
$Bu+\frac{1}{\mu}C^{T}\Lambda_{3}^{-1}Cv=-\mu\Lambda_{2}v.$ (4.28)
On the other side, nonsingularity of $(\mu\Lambda_{1}-A)$ and (4.2a) enables us the following identity
$u=(\mu\Lambda_{1}-A)^{-1}B^{T}v.$ (4.29)
Putting (4.29) on (4.28), we get
$B(\mu\Lambda_{1}-A)^{-1}B^{T}v+\frac{1}{\mu}C^{T}\Lambda_{3}^{-1}Cv=-\mu\Lambda_{2}v.$ (4.30)
As $v\neq 0,$ premultiplying by $v^{*}$ on the both sides of (4.30) yields
$v^{*}B(\mu\Lambda_{1}-A)^{-1}B^{T}v+\frac{1}{\mu}v^{*}C^{T}\Lambda_{3}^{-1}Cv=-\mu v^{*}\Lambda_{2}v.$ (4.31)
Consider the following eigenvalue decomposition of $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$
$\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}=V\mathcal{D}V^{T},$
where $\mathcal{D}=\mathrm{diag}(\theta_{i})\in{\mathbb{R}}^{n\times n}$ and $\theta_{i}>0,\,i=1,2,\ldots,n$ are in $\sigma(\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}})$ and $V\in{\mathbb{R}}^{n\times n}$ is an orthonormal matrix. Then, (4.31) can be written as
$v^{*}B\Lambda_{1}^{-\frac{1}{2}}V(\mu I-\mathcal{D})^{-1}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v+\frac{1}{\mu}v^{*}C^{T}\Lambda_{3}^{-1}Cv=-\mu v^{*}\Lambda_{2}v.$ (4.32)
Since $(\mu I-\mathcal{D})^{-1}=\Theta_{1}-{\bf{i\mkern 1.0mu}}\Im(\mu)\Theta_{2},$ where ${\bf{i\mkern 1.0mu}}$ denotes the imaginary unit and
$\Theta_{1}={\displaystyle{\mathrm{diag}\left(\frac{\Re(\mu)-\theta_{i}}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right),\Theta_{2}=\mathrm{diag}\left(\frac{1}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right)}},$
from (4.32), and the fact $\Im(\mu)\neq 0,$ we get
$v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\Theta_{1}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v+\frac{\Re(\mu)}{|\mu|^{2}}v^{*}C^{T}\Lambda_{3}^{-1}Cv=-\Re(\mu)v^{*}\Lambda_{2}v,$ (4.33)
$v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\Theta_{2}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v+\frac{1}{|\mu|^{2}}v^{*}C^{T}\Lambda_{3}^{-1}Cv=v^{*}\Lambda_{2}v.$ (4.34)
Observed that, $\Theta_{2}\leq\frac{1}{\Im(\mu)^{2}}I,$ then from(4.34), we obtain
$1\leq\frac{1}{\Im(\mu)^{2}}\left(\frac{v^{*}B\Lambda_{1}^{-1}B^{T}v}{v^{*}\Lambda_{2}v}+\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}\right).$
Therefore, we get
$|\Im(\mu)|\leq\sqrt{\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}.$ (4.35)
Furthermore, notice that $\Theta_{1}\geq\left(\Re(\mu)-\lambda_{\max}(\Lambda_{1}^{-1}A)\right)\Theta_{2},$ then from (4.33), we deduce
$-\Re(\mu)\geq\left(\Re(\mu)-\lambda_{\max}(\Lambda_{1}^{-1}A)\right)\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\Theta_{2}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v}+\frac{\Re(\mu)}{|\mu|^{2}}\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}.$ (4.36)
Using (4.34) to (4.36), we get
$2\Re(\mu)\leq\lambda_{\max}(\Lambda_{1}^{-1}A)-\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{|\mu|^{2}}\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}\leq\lambda_{\max}(\Lambda_{1}^{-1}A).$ (4.37)
Hence, (4.37) yields the following bound
$\Re(\mu)\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2}.$ (4.38)
On the other side, from (4.33) and (4.34), we deduce that
$2\Re(\mu)=\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V(\Re(\mu)\Theta_{2}-\Theta_{1})V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v}\geq\lambda_{\min}(\Lambda_{1}^{-1}A)\,\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\mathcal{X}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v},$ (4.39)
where ${\displaystyle{\mathcal{X}=\mathrm{diag}\left(\frac{1}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right).}}$ Note that
${\displaystyle{\frac{1}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\geq\frac{1}{\displaystyle{\max_{i}}(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}.}}$
Now, using Proposition 4.1 and (4.38), we get
$-\theta_{i}<\Re(\mu)-\theta_{i}\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2}-\theta_{i},$
which further yields $(\Re(\mu)-\theta_{i})^{2}\leq\max\left\{\left(\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2}-\theta_{i}\right)^{2},\theta_{i}^{2}\right\}.$ Hence, we get
$\displaystyle{\max_{i}}(\Re(\mu)-\theta_{i})^{2}$ $=$ $\displaystyle{\max_{i}}\{(\Re(\mu)-\lambda_{\min}(\Lambda_{1}^{-1}A))^{2},(\Re(\mu)-\lambda_{\max}(\Lambda_{1}^{-1}A))^{2}\}$ (4.40)
$\leq$ $\displaystyle{\max_{i}}\left\{\left(\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2}-\lambda_{\min}(\Lambda_{1}^{-1}A)\right)^{2},\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)\right\}$
$=$ $\lambda_{\max}^{2}(\Lambda_{1}^{-1}A).$
Therefore, by (4.40) and from the bound in (4.35), we obtain
$\mathcal{X}\geq\frac{I}{\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)+\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}$ (4.41)
Combining (4.39) and (4.41) leads to the following bounds
$2\Re(\mu)$$\geq$ $\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)}{\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)+\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}\frac{v^{*}B\Lambda_{1}^{-1}B^{T}v}{v^{*}\Lambda_{2}v}$ (4.42)
$\geq$ $\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)\lambda_{\min}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}{\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)+\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}.$
Hence, the proof of part $(2).$ follows by merging the inequalities of (4.35) and (4.42). $\blacksquare$ Utilizing the established bounds in Theorem 4.3, we derive the subsequent estimations for non-real eigenvalues of SS and EGSS preconditioned matrices.
Corollary 4.3.
Let ${\mathscr{P}}_{SS}$ be defined as in (1.16) with $\alpha>0$ and let $\lambda$ be a non-real eigenvalue of the SS preconditioned matrix ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector. Then, $(1).$   If $Cv=0,$ then $\lambda$ satisfies
$\frac{2\lambda_{\min}(A)}{2\alpha+\lambda_{\min}(A)}\leq|\lambda|\leq\sqrt{\frac{4\lambda_{\max}(BB^{T})}{\alpha^{2}+\alpha\lambda_{\min}(A)+\lambda_{\max}(BB^{T})}}.$
$(2).$    If $Cv\neq 0,$ the real and imaginary part of $\lambda/(2-\lambda)$ satisfies
[…]
$\left|\Im\left(\frac{\lambda}{2-\lambda}\right)\right|\leq\frac{\sqrt{\lambda_{\max}(BB^{T})+\lambda_{\max}(C^{T}C)}}{\alpha}.$
Proof.
Since ${\mathscr{P}}_{SS}$ is a special case of ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, then desired bounds will be obtained by setting $\Lambda_{1}=\Lambda_{2}=\Lambda_{3}=\frac{1}{2}\alpha I$ and $s=\frac{1}{2}$ into Theorem 4.3. $\blacksquare$
Corollary 4.4.
Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.17), and let $P,Q$ and $W$ are SPD matrices with $\alpha,\beta,\gamma>0.$ Let $\lambda$ be a non-real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector. Then $(1).$ If $Cv=0,$ then $\lambda$ satisfies
$\frac{2\lambda_{\min}(P^{-1}A)}{2\alpha+\lambda_{\min}(P^{-1}A)}\leq|\lambda|\leq\sqrt{\frac{4\lambda_{\max}(Q^{-1}BP^{-1}B^{T})}{\alpha\beta+\beta\lambda_{\min}(P^{-1}A)+\lambda_{\max}(Q^{-1}BP^{-1}B^{T})}}.$
$(2).$   If $Cv\neq 0,$ then the real and imaginary part of $\lambda/(2-\lambda)$ satisfies
$\frac{\lambda_{\min}(P^{-1}A)\,\lambda_{\min}(Q^{-1}BP^{-1}B^{T})}{2\left(\beta\lambda_{\max}^{2}(P^{-1}A)+\alpha\lambda_{\max}(Q^{-1}BP^{-1}B^{T})+\frac{\alpha^{2}}{\gamma}\lambda_{\max}(Q^{-1}C^{T}W^{-1}C)\right)}$
[…]
$\left|\Im\left(\frac{\lambda}{2-\lambda}\right)\right|\leq\sqrt{\frac{1}{\alpha\beta}\lambda_{\max}(Q^{-1}BP^{-1}B^{T})+\frac{1}{\beta\gamma}\lambda_{\max}(\Lambda_{2}^{-1}C^{T}W^{-1}C)}.$
Proof.
Since ${\mathscr{P}}_{EGSS}$ is a special case of ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, then the desired bounds will be obtained by setting $\Lambda_{1}=\frac{1}{2}\alpha P,$ $\Lambda_{2}=\frac{1}{2}\beta Q$, $\Lambda_{3}=\frac{1}{2}\gamma W$ and $s=\frac{1}{2}$ in Theorem 4.3. $\blacksquare$
5 Numerical experiments
In this section, we conduct a few numerical experiments to showcase the superiority and efficiency of the proposed PESS preconditioner over the existing preconditioners to enhance the convergence speed of the Krylov subspace iterative process to solve three-by-three block SPPs. Our study involves a comparative analysis among GMRES process and the PGMRES process employing the proposed PESS preconditioner ${\mathscr{P}}_{PESS}$ and the existing preconditioners ${\mathscr{P}}_{BD}$, ${\mathscr{P}}_{IBD}$, ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS}$ and ${\mathscr{P}}_{EGSS}.$ The numerical results are reported from the aspect of iteration counts (abbreviated as “IT”) and elapsed CPU times in seconds (abbreviated as “CPU”). Each subsystem featured in Algorithm 1 is precisely solved by applying the Cholesky factorization of the coefficient matrices. For all iterative process, the initial guess vector is ${\bf u}_{0}=0\in{\mathbb{R}}^{n+m+p}$, and the termination criterion is
${\tt RES}:=\frac{\|\mathcal{A}{\bf u}^{k+1}-{\bf d}\|_{2}}{\|{\bf d}\|_{2}}<10^{-6}.$
The vector ${\bf d}\in{\mathbb{R}}^{n+m+p}$ is chosen so that the exact solution of the system (1.1) is ${\bf u}_{*}=[1,1,\ldots,1]^{T}\in{\mathbb{R}}^{n+m+p}.$ All numerical tests are run in MATLAB (version R2023a) on a Windows 11 operating system with Intel(R) Core(TM) i7-10700 CPU, 2.90GHz, 16 GB memory.
Table 2: Numerical results of GMRES, BD, IBD, SS, RSS, EGSS and PESS PGMRES process for Example 1.
Process $l$ $16$ $32$ $48$ $64$ $80$
size($\mathcal{A}$) $1024$ $4096$ $9216$ $16384$ $25600$
[…]
CPU $0.6607$ $101.7659$ $989.0242$ $2116.1245$ $3755.9251$
RES $078.2852e-$ 07 $079.9189e-$ 07 $079.8389e-$ 07 $031.0813e-$ 03 $031.9862e-$ 03
BD IT $4$ $4$ $4$ $4$ $4$
CPU 0.0611 $2.3135$ $17.0971$ $105.4514$ $337.4322$
RES $131.2728e-$ 13 $131.7577e-$ 13 $136.6653e-$ 13 $126.0211e-$ 12 $122.6610e-$ 12
IBD IT $22$ $22$ $21$ $21$ $21$
CPU 0.0702 $1.1711$ $9.9850$ $44.2197$ $141.1108$
RES $074.0221e-$ 07 $074.6059e-$ 07 $079.1711e-$ 07 $077.8518e-$ 07 $076.8357e-$ 07
Case I
SS IT $5$ $4$ $4$ $4$ $4$
CPU $0.2221$ $1.1664$ $6.6960$ $32.3304$ $106.8526$
RES $084.7675e-$ 08 $078.7496e-$ 07 $077.1592e-$ 07 $076.2088e-$ 07 $075.5601e-$ 07
RSS IT $5$ $4$ $4$ $4$ $4$
CPU $0.2099$ $1.1297$ $6.5668$ $32.6876$ $109.8077$
RES $083.2028e-$ 08 $079.5518e-$ 07 $077.9278e-$ 07 $076.9203e-$ 07 $076.2180e-$ 07
EGSS IT $4$ $4$ $4$ $4$ $4$
CPU $0.2592$ $1.3542$ $7.9751$ $34.3676$ $125.4022$
RES $081.8724e-$ 08 $081.7563e-$ 08 $081.7090e-$ 08 $081.6765e-$ 08 $081.6497e-$ 08
${}^{\dagger}$ IT 2 2 2 2 2
CPU $0.2078$ 0.8102 4.6363 21.1566 75.4631
$s=12$ RES $073.1630e-$ 07 $072.3239e-$ 07 $071.9486e-$ 07 $071.7260e-$ 07 $071.5754e-$ 07
${}^{\dagger}$ IT 2 2 2 2 2
CPU $0.2289$ 0.8097 4.6984 21.1552 74.09290
$s=13$ RES $072.6946e-$ 07 $071.9800e-$ 07 $071.6603e-$ 07 $071.4708e-$ 07 $071.3426e-$ 07
Case II
SS IT $10$ $9$ $9$ $9$ $9$
CPU $0.2879$ $2.0186$ $12.7565$ $63.8265$ $211.2261$
RES $071.4343e-$ 07 $077.5194e-$ 07 $076.0005e-$ 07 $075.2050e-$ 07 $074.6709e-$ 07
RSS IT $9$ $9$ $9$ $9$ $9$
CPU $0.1990$ $1.8589$ $12.7943$ $62.9284$ $211.9332$
RES $077.8940e-$ 07 $075.4297e-$ 07 $074.5044e-$ 07 $073.8357e-$ 07 $073.3306e-$ 07
EGSS IT $9$ $9$ $9$ $9$ $9$
CPU $0.2130$ $1.4008$ $8.7647$ $41.7527$ $156.1426$
RES $071.1981e-$ 07 $071.7824e-$ 07 $071.5671e-$ 07 $071.4056e-$ 07 $071.2878e-$ 07
${}^{\dagger}$ IT $3$ $3$ $3$ $3$ $3$
CPU $0.2344$ $0.9795$ 5.8979 28.3526 98.0878
$s=12$ RES $081.0992e-$ 08 $082.4637e-$ 08 $084.8787e-$ 08 $087.2624e-$ 08 $088.4769e-$ 08
${}^{\dagger}$ IT $3$ $3$ $3$ $3$ $3$
CPU $0.1929$ $0.9965$ $5.9099$ $28.0443$ $101.0273$
$s=13$ RES $098.6456e-$ 09 $081.9393e-$ 08 $083.8703e-$ 08 $085.8810e-$ 08 $087.0253e-$ 08
[…]
Example 1.
We consider the three-by-three block SPP (1.1) taken from [18], where
$A=\left[\begin{matrix}I\otimes G+G\otimes I&0\\ 0&I\otimes G+G\otimes I\end{matrix}\right]\in{\mathbb{R}}^{2l^{2}\times 2l^{2}},\,\,B=\left[\begin{matrix}I\otimes F&F\otimes I\end{matrix}\right]\in{\mathbb{R}}^{l^{2}\times 2l^{2}},$ and $C=E\otimes F\in{\mathbb{R}}^{l^{2}\times l^{2}},$ where $G=\frac{1}{(l+1)^{2}}\,\mathrm{tridiag}(-1,2,-1)\in{\mathbb{R}}^{l\times l},\quad F=\frac{1}{l+1}\,\mathrm{tridiag}(0,1,-1)\in{\mathbb{R}}^{l\times l}$ , and $E=\mathrm{diag}(1,l+1,\ldots,l^{2}-l+1)\in{\mathbb{R}}^{l\times l}.$ Here, $E\otimes F$ denotes the Kronecker product of two matrices $E$ and $F,$ and $\mathrm{tridiag}(a,b,c)\in{\mathbb{R}}^{l\times l}$ denotes the $l\times l$ tridiagonal matrix with diagonal, subdiagonal, and superdiagonal entries are $a,b,$ and $c,$ respectively. For this problem, the size of the matrix $\mathcal{A}$ is $4l^{2}.$ We consider the parameter choices for the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS},{\mathscr{P}}_{EGSS}$ and ${\mathscr{P}}_{PESS}$ in two cases.
  • In Case I: $\alpha=0.2$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=0.2,\beta=2,\gamma=0.002$ and $P=I,Q=I,W=I$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=\frac{1}{2}\alpha I,\Lambda_{2}=\frac{1}{2}\beta I,\Lambda_{3}=\frac{1}{2}\gamma I$ for ${\mathscr{P}}_{PESS}.$
  • In Case II: $\alpha=2$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=2,\beta=2,\gamma=0.002$ and $P=I,Q=I,W=CC^{T}$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=\frac{1}{2}\alpha I,\Lambda_{2}=\frac{1}{2}\beta I,\Lambda_{3}=\frac{1}{2}\gamma CC^{T}$ for ${\mathscr{P}}_{PESS}.$
The optimal values (experimentally) for $s$ in the range $[10,20]$ are determined to be $12$ and $13$ for minimal CPU times. Following [18], selection of $\hat{A}$ and $\hat{S}$ (SPD approximations of $A$ and $S,$ respectively) for IBD preconditioner are done as follows:
$\hat{A}=LL^{T},\quad\hat{S}=\mathrm{diag}(B\hat{A}^{-1}B^{T}),$
where $L$ is the incomplete Cholesky factor of $A$ produced by the Matlab function:
[…]
The test problems generated for the values of $l=16,32,48,64,80,$ and their numerical results are reported in Table 2.
(a) $l=16$
(b) $l=32$
(c) $l=48$
(d) $l=64$
(e) $l=80$
Figure 1: Convergence curves for IT versus RES of PGMRES process employing BD, IBD, SS, RSS, EGSS, and PESS $(s=13)$ preconditioners in Case II for Example 1.
(a) $\mathcal{A}$
(b) ${\mathscr{P}}_{BD}^{-1}\mathcal{A}$
(c) ${\mathscr{P}}_{IBD}^{-1}\mathcal{A}$
(d) ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$
(e) ${\mathscr{P}}_{RSS}^{-1}\mathcal{A}$
(f) ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$
(g) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=12$
Figure 2: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=16$ for Example 1.
Figure 3: Parameter $s$ vs condition number of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ for Example 1.
Figure 1 illustrates convergence curves pertaining to preconditioners BD, IBD, SS, RSS, EGSS, and PESS ($s=13$) for Case II with $l=16,32,48,64$ and $80$ These curves depict the relationship between the relative residue at each iteration step (RES) and IT counts. To further illustrate the superiority of the PESS preconditioner, spectral distributions of $\mathcal{A}$ and the preconditioned matrices ${\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{RSS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for the Case II with $l=16$ and $s=12$ are displayed in Figure 2. To investigate the robustness of the proposed PESS preconditioner, we measure the condition number of the ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ which is for any nonsingular matrix $A$ defined by $\kappa(A):=\|A^{-1}\|_{2}\|A\|_{2}.$ In Figure 3, the influence of the parameter $s$ on the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ is depicted. The parameter $s$ is chosen from the interval $[5,50]$ with step size one. The data presented in Table 2 and Figures 1-3 allow us to make the following noteworthy observations:
  • From Table 2, it is observed that the GMRES has a very slow convergence speed. In both Cases I and II, our proposed PESS preconditioner outperforms all other compared existing preconditioners from the aspects of IT and CPU times. For example, in Case II with $l=80,$ our proposed PESS preconditioner is $71\%$, $30\%$, $54\%$, $54\%$, and $37\%$ more efficient than the existing BD, IBD, SS, RSS, and EGSS preconditioner, respectively. Similar patterns are noticed for $l=16,32,48$ and $64$.
  • In Figure 1, it is evident that the PESS preconditioner has a faster convergence speed than the other baseline preconditioners when applied to the PGMRES process for all $l=16,32,48,64$ and $80$
  • According to Figure 2, the spectrum of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ has better clustering properties than the baseline preconditioned matrices, consequently improves the computational performance of our proposed PESS PGMRES process. Moreover, the real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ are contained in the interval $(0,0.08330]$ and for ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ they are in $(0,1.9983],$ which are consistent with bounds of Theorem 4.2, Corollaries 3.1, and 3.2, respectively. For non-real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ obtained bounds are consistent with Theorem 4.3.
  • For $l=32,$ the computed condition numbers of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},$ and ${\mathscr{P}}_{IBD}^{-1}\mathcal{A}$ are $5.4289e+04,$ $9.5567e+09,$ and $9.5124e+09,$ respectively, which are very large. Whereas Figure 3 illustrate a decreasing trend in the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ (for instance, when $s=50$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=3.422$) with increasing values of $s.$ This observation highlights that ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ emerges as well-conditioned, consequently the solution obtained by the proposed PESS PGMRES process is more reliable and robust.
The above discussions affirm that the proposed PESS preconditioner is efficient, robust, and better well-conditioned with respect to the baseline preconditioners.
Figure 4: Relationship of norm error of solution with increasing noise percentage, employing proposed ${\mathscr{P}}_{PESS}$ for Case II with $s=12$ for $l=16,32,48,64$ and $80$ of Example 1.
Sensitivity analysis of the solution by employing the proposed PESS preconditioner
To study the sensitivity of the solution obtained by employing the proposed PESS PGMRES process to small perturbations on the input data matrices. In the following, we consider the perturbed counterpart $(\mathcal{A}+\Delta\mathcal{A})\tilde{\textbf{u}}=\textbf{d}$ of the three-by-three block SPP (1.1), where the perturbation matrix $\Delta\mathcal{A}$ has same structure as of $\mathcal{A}.$ We construct the perturbation matrix $\Delta\mathcal{A}$ by adding noise to the block matrices $B$ and $C$ of $\mathcal{A}$ as follows:
$\Delta B=\varepsilon*N_{P}*std(B).*randn(m,n)\quad\text{and}\quad\Delta C=\varepsilon*N_{P}*std(C).*randn(p,m),$
where $\varepsilon=10^{-4}$ and $N_{P}$ is the noise percentage and $std(B)$ and $std(C)$ are the standard deviation of $B$ and $C,$ respectively. Here, $randn(m,n)$ is the is an $m$-by-$n$ normally distributed random matrix, generated by the MATLAB command $randn$. We perform the numerical test for $l=16,32,48,64$ and $80$ for Example 1 using the PESS PGMRES process (Case II, $s=12$). The calculated norm error $\|\tilde{\textbf{u}}-\textbf{u}\|_{2}$ among the solution of the perturbed system and the original system with increasing $N_{P}$ from $5\%$ to $40\%$ with step size $5\%$ are displayed in the figure 4. We noticed that with growing $N_{P},$ the norm error of the solution remains consistently less than $10^{-8},$ which demonstrates the robustness of the proposed PESS preconditioner and the solution u is not sensitive to small perturbation on $\mathcal{A}$.
Table 3: Numerical results of GMRES, BD, IBD, SS, RSS, EGSS, and PESS PGMRES processes for Example 2.
Process h $1/8$ $1/16$ $1/32$ $1/64$
size($\mathcal{A}$) $288$ $1088$ $4224$ $16640$
GMRES IT $160$ $555$ $2047$bf{--}- -
CPU $\bf{0.0279}$ $0.2233$ $55.1948$ $2097.3235$
RES $078.0113e-$ 07 $078.8023e-$ 07 $079.0887e-$ 07 $064.0481e-$ 06
IBD IT $34$ $35$ $25$ $23$
CPU $\bf{0.0066}$ $\bf 0.0697$ $\bf 1.4026$ $53.9580$
RES 8.8956e-07 6.2172e-07 9.5786e-07 7.9687e-07
Case I
SS IT $7$ $8$ $11$ $9$
CPU $0.1387$ $0.2258$ $2.4943$ $70.2884$
RES $071.4964e-$ 07 $074.2135e-$ 07 $072.2106e-$ 07 $078.7734e-$ 07
RSS IT $\bf 4$ $5$ $5$ $6$
CPU $0.1487$ $0.2089$ $1.4538$ $49.1211$
RES $079.0256e-$ 07 $083.7846e-$ 08 $071.9761e-$ 07 $081.0091e-$ 08
EGSS IT $9$ $10$ $13$ $18$
CPU $0.1437$ $0.2381$ $2.7262$ $133.7547$
RES $081.8639e-$ 08 $073.9243e-$ 07 $071.6563e-$ 07 $075.1266e-$ 07
${}^{\dagger}$ IT 4 4 5 5
CPU $0.1455$ 0.2062 $1.5472$ 44.7190
$s=25$ RES $071.4060e-$ 07 $075.5647e-$ 07 $088.6496e-$ 08 $074.3922e-$ 07
Case II
SS IT $41$ $56$ $77$ $66$
CPU $0.1717$ $0.5602$ $14.0111$ $466.4250$
RES $078.2375e-$ 07 $078.9915e-$ 07 $079.1640e-$ 07 $079.7171e-$ 07
RSS IT $17$ $18$ $15$ $15$
CPU $0.1564$ $0.2734$ $2.9773$ $109.9137$
RES $071.6292e-$ 07 $071.4261e-$ 07 $075.8149e-$ 07 $076.0840e-$ 07
EGSS IT $9$ $9$ $9$ $7$
CPU $0.1403$ $0.2274$ $2.0706$ $54.4502$
RES $084.1786e-$ 08 $091.4701e-$ 09 $092.4144e-$ 09 $072.6089e-$ 07
${}^{\dagger}$ IT $5$ 5 5 5
CPU $0.1278$ $0.2112$ 1.5175 44.0049
$s=26$ RES $082.0196e-$ 08 $081.8653e-$ 08 $081.5968e-$ 08 $085.7013e-$ 08
[…]
Example 2.
[…]
(a) $\textbf{h}=1/8$
(b) $\textbf{h}=1/16$
(c) $\textbf{h}=1/32$
(d) $\textbf{h}=1/64$
Figure 5: Convergence curves for IT versus RES of the PGMRES processes by employing IBD, SS, RSS, EGSS, and PESS preconditioners in Case II for Example 2.
[…]
  • In Case I: $\alpha=0.02$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=0.02,\beta=0.2,\gamma=0.002$ and $P=I,Q=I,W=I$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=\frac{1}{2}\alpha I,\Lambda_{2}=\frac{1}{2}\beta I,\Lambda_{3}=\frac{1}{2}\gamma I$ for ${\mathscr{P}}_{PESS}.$
  • In Case II: $\alpha=2$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=2,\beta=2,\gamma=0.002$ and $P=A,Q=I,W=CC^{T}$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=\frac{1}{2}\alpha A,\Lambda_{2}=\frac{1}{2}\beta I,\Lambda_{3}=\frac{1}{2}\gamma CC^{T}.$
For the IBD preconditioner, the matrices $\hat{A}$ and $\hat{S}$ are constructed as in Example 1. In the interval $[10,30]$, the empirical optimal choice for $s$ is found to be $25$ for Case I and $26$ for Case II. Table 3 shows that the GMRES process has a very slow convergence speed and also does not converge for $\textbf{h}=1/64$ within $7000$ iterations. The proposed PESS preconditioner requires almost five times fewer iterations compared to the IBD preconditioner for convergence. Moreover, in both cases, the PESS preconditioned process outperforms the SS, RSS, and EGSS preconditioners in terms of IT and CPU times. Notably, IT for the PESS preconditioner, in both cases, remains constant, whereas for SS, RSS, and EGSS, IT increases as the size of $\mathcal{A}$ increases. Furthermore, in Case II with $l=1/64$, our proposed PESS preconditioner is approximately $18\%$, $91\%$, $60\%$, and $19\%$ more efficient than the existing IBD, SS, RSS, and EGSS preconditioner, respectively. Similar patterns are observed for other values of $l.$
The convergence curves in Figure 5 demonstrate the rapid convergence of the proposed PESS preconditioner compared to the baseline IBD, SS, RSS, EGSS, and PESS preconditioners in terms of RES versus IT counts.
(a) $\mathcal{A}$
(b) ${\mathscr{P}}_{IBD}^{-1}\mathcal{A}.$
(c) ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$
(d) ${\mathscr{P}}_{RSS}^{-1}\mathcal{A}$
(e) ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$
(f) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=26$
Figure 6: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with ${\bf h}=1/8.$
(a) CPU versus parameters $s$ and $\alpha$
(b) IT versus parameters $s$ and $\alpha$
Figure 7: Characteristic curves for CPU (left) and IT (right) of the proposed PESS iteration process by varying $s$ in the interval $[1,30]$ with step size $1$ and $\alpha$ in the interval $[0.05,1]$ with step size $0.05$ with $\textbf{h}=1/32$ in Case I for Example 2.
Moreover, Figure 6 illustrates the spectral distributions of the original matrix $\mathcal{A},$ and the preconditioned matrices ${\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{RSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with ${\bf{h}}=1/8.$ According to Figure 6, eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ demonstrate a superior clustering compared to the other preconditioned matrices, which leads to a favorable convergence speed for the proposed PESS PGMRES process. Furthermore, the system (1.1) exhibits ill-conditioning nature with $\kappa(\mathcal{A})=5.07e+05.$ While for Case I and $\mathbf{h}=1/16$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=1.4024,$ indicating that the proposed PESS preconditioned system is well-conditioned, ensuring an efficient and robust solution. In addition, to demonstrate the relationship of the parameters and the convergence speed of the PESS preconditioner, we plot the graph of CPU times and IT counts by varying the parameters $s$ and $\alpha,$ in Figure 7. For $\textbf{h}=1/32$, $s$ varying in the interval $[1,30]$ with step size one and $\alpha$ from the interval $[0.05,1]$ with step size $0.05$ Other parameters remain consistent with Case I. We draw the following observations from Figure 7.
  • For the given value of $s$, the proposed PESS preconditioner is insensitive to the change of the parameter $\alpha.$
  • With an increase in $s$, IT and CPU times show a substantial decrease irrespective of the value of $\alpha.$
6 Conclusion
In this paper, we proposed the PESS iterative process and corresponding PESS preconditioner ${\mathscr{P}}_{PESS}$ to solve the three-by-three block SPP (1.1). For the convergence of the proposed PESS iterative process, necessary and sufficient criteria are derived. Moreover, we estimated the spectral bounds of the proposed PESS preconditioned matrices. This empowers us to derive spectral bounds for SS and EGSS preconditioned matrices, which have not yet been investigated in the literature to the best of our knowledge. Numerous experimental analyses are performed to demonstrate the effectiveness of our proposed PESS preconditioner. The key observations are as follows: $(i)$ the proposed PESS preconditioner is found to outperform the existing baseline preconditioners in terms of IT and CPU times. $(ii)$ The proposed PESS preconditioner significantly reduces the condition number of $\mathcal{A}$, consequently showing the its proficiency in solving three-by-three block SPPs. $(iii)$ The PESS preconditioned matrix has a more clustered spectral distribution than the baseline preconditioned matrices. $(iv)$ Sensitivity analysis conducted by introducing different percentages of noise on the system (1.1) showcases the robustness of the proposed PESS preconditioner.
Acknowledgements
The second author (Pinki Khatun) sincerely acknowledges the Council of Scientific &\&& Industrial Research (CSIR) in New Delhi, India, for financial assistance in the form of a fellowship (File no. 09/1022(0098)/2020-EMR-I).
Declaration of competing interest
The authors disclose that they have no conflicting interests.
Data availability
No data was used for the research described in the article.

### v2

@@
 This paper proposes a new parameterized enhanced shift-splitting (PESS) preconditioner to solve the three-by-three block saddle point problem (SPP).
-In addition, necessary and sufficient criteria are established for the convergence of the proposed PESS iterative process for any random initial guess.
-Furthermore, we meticulously investigate the spectral bounds of the PESS preconditioned matrix.
-Moreover, empirical investigation has been performed for the sensitivity analysis of the system, revealing the robustness of the proposed PESS preconditioner.
-Numerical experiments are carried out to demonstrate the enhanced efficiency and robustness of the proposed PESS preconditioner compared to the existing block diagonal and shift-splitting preconditioners.
-Keywords. Krylov subspace iterative process $\cdot$ Preconditioner $\cdot$ Shift-splitting $\cdot$ Saddle point problem $\cdot$ Robust AMS subject classification. 65F08 $\cdot$ 65F10 $\cdot$ 65F50
+Additionally, we introduce a local PESS (LPESS) preconditioner by relaxing the PESS preconditioner.
+Necessary and sufficient criteria are established for the convergence of the proposed PESS iterative process for any random initial guess.
+Furthermore, we meticulously investigate the spectral bounds of the PESS and LPESS preconditioned matrices.
+Moreover, empirical investigations have been performed for the sensitivity analysis of the proposed PESS preconditioner, which unveils its robustness.
+Numerical experiments are carried out to demonstrate the enhanced efficiency and robustness of the proposed PESS and LPESS preconditioners compared to the existing block diagonal and shift-splitting preconditioners.
+Keywords. Krylov subspace iterative process $\cdot$ Preconditioner $\cdot$ Shift-splitting $\cdot$ Saddle point problem $\cdot$ Robust $\cdot$ GMRES
+AMS subject classification. 65F08 $\cdot$ 65F10 $\cdot$ 65F50
 1 Introduction
@@
 Furthermore, $\mathcal{A}$ is generally sparse and large; therefore, iterative processes become superior to direct methods.
-The Krylov subspace iterative process is preferable among other iterative processes investigated recently in the literature due to their minimal storage requirement and feasible implementation [29].
+The Krylov subspace iterative process is preferable among other iterative processes investigated widely in the literature due to their minimal storage requirement and feasible implementation [30].
 Slow convergence of the Krylov subspace iterative process is a major drawback for system of linear equations of large dimensions.
@@
 Consider the following partitioning of $\mathcal{A}:$
-$\left[\begin{array}[]{cc|c}A&B^{T}&0\\ -B&0&-C^{T}\\ \hline\cr 0&C&\hfil 0\hfil\\ \end{array}\right],\qquad\left[\begin{array}[]{c|cc}A&B^{T}&0\\ \hline\cr-B&0&-C^{T}\\ 0&C&0\\ \end{array}\right],$ (1.8)
+[…]
 then the linear system (1.1) can be identified as the standard two-by-two block SPP [5].
@@
 However, the properties of the submatrices in (1.8) are different from the standard two-by-two block SPP.
-Notice that the leading blocks are skew-symmetric and positive semi-definite in the first partitioning.
+Notice that the leading blocks are skew-symmetric and positive semidefinite in the first partitioning.
 The second partitioning highlights that the $(1,2)$ block is rank deficient, and the $(2,2)$ block is skew-symmetric.
-In contrast, in the standard two-by-two block SPP, $(1,2)$ block is full row rank, and the $(2,2)$ block is generally symmetric positive semidefinite.
+In contrast, in the standard SPP, $(1,2)$ block is full row rank, and the $(2,2)$ block is generally symmetric positive semidefinite.
 Therefore, the existing literature for solving the two-by-two SPPs (see [5]) may not be applied to solve (1.1) directly, and it is essential to develop new preconditioners for three-by-three block SPP that exploit the specific structure of the saddle point matrix $\mathcal{A}$, which is sensitive in nature.
@@
 For instance, Huang and Ma [18] has studied the exact block diagonal (BD) and inexact BD (IBD) preconditioners ${\mathscr{P}}_{BD}$ and ${\mathscr{P}}_{IBD}$ in the following forms:
-${\mathscr{P}}_{BD}=\left[\begin{matrix}A&0&0\\ 0&S&0\\ 0&0&CS^{-1}C^{T}\end{matrix}\right]\quad\mbox{and}\quad{\mathscr{P}}_{IBD}=\left[\begin{matrix}\hat{A}&0&0\\ 0&\hat{S}&0\\ 0&0&C\hat{S}^{-1}C^{T}\end{matrix}\right],$ (1.15)
-where $S=BA^{-1}B^{T},$ $\hat{A}$ and $\hat{S}$ are SPD approximations of $A$ and $S,$ respectively.
+${\mathscr{P}}_{BD}=\left[\begin{matrix}A&0&0\\ 0&S&0\\ 0&0&CS^{-1}C^{T}\end{matrix}\right]\quad\mbox{and}\quad{\mathscr{P}}_{IBD}=\left[\begin{matrix}\widehat{A}&0&0\\ 0&\widehat{S}&0\\ 0&0&C\widehat{S}^{-1}C^{T}\end{matrix}\right],$ (1.9)
+where $S=BA^{-1}B^{T},$ $\widehat{A}$ and $\widehat{S}$ are SPD approximations of $A$ and $S,$ respectively.
 Although the spectrum of the preconditioned matrices ${\mathscr{P}}^{-1}_{BD}\mathcal{A}$ and ${\mathscr{P}}^{-1}_{IBD}\mathcal{A}$ have good clustering properties, these preconditioners have certain shortcomings, such as they are time-consuming, expensive, require excessive number of iterations, and condition numbers are very large.
@@
 Recently, Cao [9] enhanced this idea and introduced the following SS preconditioner ${\mathscr{P}}_{SS}$ and relaxed SS (RSS) preconditioner ${\mathscr{P}}_{RSS}$ for the three-by-three block SPP (1.1)
-${\mathscr{P}}_{SS}=\frac{1}{2}\left[\begin{matrix}\alpha I+A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right],\,\,\mbox{and}\,\,{\mathscr{P}}_{RSS}=\frac{1}{2}\left[\begin{matrix}A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right],$ (1.16)
+${\mathscr{P}}_{SS}=\frac{1}{2}\left[\begin{matrix}\alpha I+A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right]\,\,\mbox{and}\,\,{\mathscr{P}}_{RSS}=\frac{1}{2}\left[\begin{matrix}A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right],$ (1.10)
 where $\alpha>0$ and verified unconditionally convergence of the associated SS iterative process.
 Here, $I$ stands for the identity matrix of the appropriate dimension.
-Wang and Zhang [27] generalized the SS preconditioner by introducing a new parameter $\beta>0$ in the $(3,3).$ Further, Yin et al.
+Wang and Zhang [27] generalized the SS preconditioner by introducing a new parameter $\beta>0$ in the $(3,3)$ block.
+Further, Yin et al.
 [31] developed an extensive generalized SS (EGSS) preconditioner
@@
 We summarize the following main contributions of this paper:
-  • We propose the PESS iterative process and associated PESS preconditioner to solve the three-by-three block SPP (1.1).
-The proposed preconditioner generalizes the other SS preconditioners studied in the literature see, for example, [9, 27, 30].
+  • We propose the PESS iterative process along with its associated PESS preconditioner and its relaxed version known as the local PESS (LPESS) preconditioner to solve the three-by-three block SPP (1.1).
+The proposed preconditioners generalize the existing SS preconditioners, for example, [9, 27, 31].
   • General framework is given on necessary and sufficient criteria for the convergence of the PESS iterative process.
@@
 Our framework allows us to derive the spectral distribution for the SS and EGSS preconditioned matrices.
-  • We empirically show that our proposed PESS preconditioner significantly reduces the condition number of the ill-conditioned saddle point matrix $\mathcal{A}$.
+  • We empirically show that our proposed preconditioner significantly reduces the condition number of the ill-conditioned saddle point matrix $\mathcal{A}$.
 Thereby establishing an efficiently solvable, well-conditioned system.
-  • Numerical experiments show that the proposed PESS preconditioner outperforms the existing BD and SS preconditioners.
+  • Numerical experiments show that the proposed PESS and LPESS preconditioners outperform the existing BD and SS preconditioners.
 Moreover, we perform sensitivity analysis by perturbing the ill-conditioned matrix $\mathcal{A}$ in (1.1), demonstrating the robustness of the proposed preconditioner.
@@
 In Section 4, we investigate the spectral distribution of the PESS preconditioned matrix.
+In Section 5, we present the LPESS preconditioner.
 In Section 6, we conduct numerical experiments to illustrate the computational efficiency and robustness of the proposed preconditioner.
@@
 The symbol $\mathrm{diag}(A,B)$ represents a block diagonal matrix with diagonal blocks $A$ and $B.$ The real and imaginary parts of any complex number $z$ are represented by $\Re(z)$ and $\Im(z)$, respectively.
-For any matrix $A\in{\mathbb{R}}^{n\times n},$ $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$ denotes the maximum and minimum eigenvalue of $A,$ respectively.
+For any symmetric matrix $A\in{\mathbb{R}}^{n\times n},$ $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$ denotes the maximum and minimum eigenvalue of $A,$ respectively.
 The notation $A>0$ signifies that $A$ is SPD, and $A>B$ identifies $A-B>0.$ Abbreviations used in the paper are as follows:
-SPP: saddle point problem, SPD: symmetric positive definite, BD: block diagonal, IBD: inexact block diagonal, SS: shift-splitting, RSS: relaxed shift-splitting, EGSS: extensive generalized shift-spliting, PESS: parameterized enhanced shift-splitting, GMRES: generalized minimum residual process.
+SPP: saddle point problem, SPD: symmetric positive definite, BD: block diagonal, IBD: inexact block diagonal, SS: shift-splitting, RSS: relaxed shift-splitting, EGSS: extensive generalized shift-spliting, PESS: parameterized enhanced shift-splitting, LPESS: local parameterized enhanced shift-splitting GMRES: generalized minimum residual process.
 2 The parameterized enhanced shift-splitting (PESS) iterative process and the preconditioner
-Let $s>0$ be a real number. Then, we write the matrix $\mathcal{A}$ in the form
-$\mathcal{A}=(\Sigma+s\mathcal{A})-(\Sigma-(1-s)\mathcal{A}):={\mathscr{P}}_{PESS}-\mathcal{Q}_{PESS},\,\mbox{where}$ (2.1)
-${\mathscr{P}}_{PESS}=\left[\begin{matrix}\Lambda_{1}+sA&sB^{T}&0\\ -sB&\Lambda_{2}&-sC^{T}\\ 0&sC&\Lambda_{3}\end{matrix}\right],\mathcal{Q}_{PESS}:=\left[\begin{matrix}\Lambda_{1}-(1-s)A&-(1-s)B^{T}&0\\ (1-s)B&\Lambda_{2}&(1-s)C^{T}\\ 0&-(1-s)C&\Lambda_{3}\end{matrix}\right],$
+In this section, we proposed a parameterized enhanced shift-splitting (PESS) iterative process and the corresponding PESS preconditioner.
+Let $s>0$ be a real number.
+Then, we split the matrix $\mathcal{A}$ in the form
+$\mathcal{A}=(\Sigma+s\mathcal{A})-(\Sigma-(1-s)\mathcal{A})=:{\mathscr{P}}_{PESS}-\mathcal{Q}_{PESS},\,\mbox{where}$ (2.1)
+${\mathscr{P}}_{PESS}=\left[\begin{matrix}\Lambda_{1}+sA&sB^{T}&0\\ -sB&\Lambda_{2}&-sC^{T}\\ 0&sC&\Lambda_{3}\end{matrix}\right],\mathcal{Q}_{PESS}=\left[\begin{matrix}\Lambda_{1}-(1-s)A&-(1-s)B^{T}&0\\ (1-s)B&\Lambda_{2}&(1-s)C^{T}\\ 0&-(1-s)C&\Lambda_{3}\end{matrix}\right],$
 $\Sigma=\mathrm{diag}(\Lambda_{1},\Lambda_{2},\Lambda_{3}),$ and $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ are SPD matrices.
-The matrix ${\mathscr{P}}_{PESS}$ is nonsingular for $s>0.$ The special matrix splitting (2.1) leads us to the subsequent iterative process for solving the three-by-three block SPP given in (1.1): The PESS iterative process.
+The matrix ${\mathscr{P}}_{PESS}$ is nonsingular for $s>0.$ The special matrix splitting (2.1) leads us to the subsequent iterative process for solving the three-by-three block SPP given in (1.1):
+The PESS iterative process.
 Let $s$ be a positive constant, and let $\Lambda_{1}\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}$ $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices.
@@
 ${\mathscr{P}}_{PESS}w=r,$ (2.3)
-where $r=[r_{1}^{T},r_{2}^{T},r_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}$ and $w=[w_{1}^{T},w_{2}^{T},w_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}.$ Specifying $\hat{X}=\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C$ and $\hat{A}=\Lambda_{1}+sA+s^{2}B^{T}\hat{X}^{-1}B,$ then ${\mathscr{P}}_{PESS}$ can be written in the following way
-${\mathscr{P}}_{PESS}=\left[\begin{matrix}I&sB^{T}\hat{X}^{-1}&0\\ 0&I&-sC^{T}\Lambda_{3}^{-1}\\ 0&0&I\end{matrix}\right]\left[\begin{matrix}\hat{A}&0&0\\ 0&\hat{X}&0\\ 0&0&\Lambda_{3}\end{matrix}\right]\left[\begin{matrix}I&0&0\\ -s\hat{X}^{-1}B&I&0\\ 0&s\Lambda_{3}^{-1}C&I\end{matrix}\right].$ (2.13)
+where $r=[r_{1}^{T},r_{2}^{T},r_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}$ and $w=[w_{1}^{T},w_{2}^{T},w_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}.$ Specifying $\widehat{X}=\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C$ and $\widehat{A}=\Lambda_{1}+sA+s^{2}B^{T}\widehat{X}^{-1}B,$ then ${\mathscr{P}}_{PESS}$ can be written in the following way
+${\mathscr{P}}_{PESS}=\left[\begin{matrix}I&sB^{T}\widehat{X}^{-1}&0\\ 0&I&-sC^{T}\Lambda_{3}^{-1}\\ 0&0&I\end{matrix}\right]\left[\begin{matrix}\widehat{A}&0&0\\ 0&\widehat{X}&0\\ 0&0&\Lambda_{3}\end{matrix}\right]\left[\begin{matrix}I&0&0\\ -s\widehat{X}^{-1}B&I&0\\ 0&s\Lambda_{3}^{-1}C&I\end{matrix}\right].$ (2.4)
 The decomposition (2.4) leads us to the following Algorithm 1 to determine the solution of the system (2.3).
 Algorithm 1 Computation of $w$ from ${\mathscr{P}}_{PESS}w=r$
-1:Solve $(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)v_{1}=r_{2}+sC^{T}\Lambda_{3}^{-1}r_{3}$ to find $v_{1};$
+1:Solve $\widehat{X}v_{1}=r_{2}+sC^{T}\Lambda_{3}^{-1}r_{3}$ to find $v_{1};$
 2:Compute $v=r_{1}-sB^{T}v_{1};$
-3:Solve $(\Lambda_{1}+sA+s^{2}B^{T}(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)^{-1}B)w_{1}=v$ to find $w_{1}$
-4:Solve $(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)v_{2}=sBw_{1}$ to find $v_{2}$;
+3:Solve $\widehat{A}w_{1}=v$ to find $w_{1}$
+4:Solve $\widehat{X}v_{2}=sBw_{1}$ to find $v_{2}$;
 5:Compute $w_{2}=v_{1}+v_{2};$
 6:Solve $\Lambda_{3}w_{3}=r_{3}-sCw_{2}$ for $w_{3}.$
-The implementation of Algorithm 1 requires to solve two linear subsystems with coefficient matrices $\hat{X}=\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C$ and $\hat{A}=\Lambda_{1}+sA+s^{2}B^{T}(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)^{-1}B$.
-Since $\Lambda_{1},\Lambda_{2},\Lambda_{3},\hat{X},$ and $\hat{A}$ are SPD and $s>0$, we can apply the Cholesky factorization to solve these linear subsystems exactly or inexactly by the preconditioned conjugate gradient method.
-Similarly, we can solve step $6$ of the Algorithm 1.
+The implementation of Algorithm 1 requires to solve two linear subsystems with coefficient matrices $\widehat{X}$ and $\widehat{A}$.
+Since $\Lambda_{1},\Lambda_{2},\Lambda_{3},\widehat{X},$ and $\widehat{A}$ are SPD and $s>0$, we can apply the Cholesky factorization to solve these linear subsystems exactly or inexactly by the preconditioned conjugate gradient method.
 3 Convergence analysis of the PESS iterative process
@@
 Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices.
-Then $\Sigma^{-1}\mathcal{A}$ (normal-(((or, $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})$ is positive stable.
+Then $\Sigma^{-1}\mathcal{A}$ ((((or, $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})$ is positive stable.
 Proof.
@@
 Then by Lemma 3.1, $\mathcal{A}$ is positive stable implies that, the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ and consequently, $\Sigma^{-1}\mathcal{A}$ is positive stable.
-$\blacksquare$ By [22], the stationary iterative process (2.2) converges to the exact solution of the three-by-three block SPP (1.1) for any initial guess vector if and only if $|\vartheta(\mathcal{T})|<1,$ where $\mathcal{T}$ is the iteration matrix.
-Now, we establish an if and only if condition that precisely determines the convergence of the PESS iterative process.
+$\blacksquare$
+By [22], the stationary iterative process (2.2) converges to the exact solution of the three-by-three block SPP (1.1) for any initial guess vector if and only if $|\vartheta(\mathcal{T})|<1,$ where $\mathcal{T}$ is the iteration matrix.
+Now, we establish an if and only if a condition that precisely determines the convergence of the PESS iterative process.
 Theorem 3.1.
@@
 $(1-(1-s)\Re(\mu))^{2}+(1-s)^{2}\Im(\mu)^{2}<(1+s\Re(\mu))^{2}+s^{2}\Im(\mu)^{2}.$
-This implies $(2s-1)|\mu|^{2}+2\Re(\mu)^{2}>0.$ Since, $\vartheta(\mathcal{T})=\vartheta{\widetilde{(}\mathcal{T}}),$ the proof is conclusive.
-$\blacksquare$ Using the condition in Theorem 3.1, next, we present two sufficient conditions that ensure the convergence of the PESS iterative process.
-The first one is presented as follows:
+This implies $|\theta|<1$ if and only if $(2s-1)|\mu|^{2}+2\Re(\mu)>0.$ Since, $\vartheta(\mathcal{T})=\vartheta{\widetilde{(}\mathcal{T}}),$ the proof is conclusive. $\blacksquare$
+Using the condition in Theorem 3.1, next, we present two sufficient conditions that ensure the convergence of the PESS iterative process. The first one is presented as follows:
 Corollary 3.1.
@@
 $\Re(\mu)=\frac{{\bf p}^{*}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}){\bf p}}{2{\bf p}^{*}{\bf p}}\geq\frac{1}{2}\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}).$
-Again, we have $|\mu|<\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}),$ then we get the following
-$\frac{1}{2}-\frac{\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{2\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}>\frac{1}{2}-\frac{\Re(\mu)}{|\mu|^{2}}.$ (3.9)
-If we apply $s>0$, then we get the desired result from (3.9).
-$\blacksquare$ The sufficient condition on Corollary 3.2 is difficult to find for large $\mathcal{A}$ due to the involvement of computation of $\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})$ and $\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}.$ Thus, we consider $s\geq\frac{1}{2}$ for practical implementation.
+Again, we have $|\mu|\leq\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}),$ then we get the following
+$\frac{1}{2}-\frac{\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{2\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}\geq\frac{1}{2}-\frac{\Re(\mu)}{|\mu|^{2}}.$ (3.6)
+If we apply $s>0$, then we get the desired result from (3.6). $\blacksquare$
+The sufficient condition on Corollary 3.2 is difficult to find for large $\mathcal{A}$ due to the involvement of computation of $\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})$ and $\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}.$ Thus, we consider $s\geq\frac{1}{2}$ for practical implementation.
 4 Spectral distribution of the PESS preconditioned matrix
@@
 Then, we get $\lambda=1-\theta.$ Now, from Corollary 3.1, for $s\geq\frac{1}{2},$ it holds that $|\theta|<1.$ Hence, it follows that, for $s\geq\frac{1}{2},$ $|\lambda-1|<1,$ which completes the proof.
-$\blacksquare$ Let $(\lambda,{\bf p}=[u^{T},v^{T},w^{T}]^{T})$ be an eigenpair of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then, we have $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p},$ which can be written as
+$\blacksquare$
+Let $(\lambda,{\bf p}=[u^{T},v^{T},w^{T}]^{T})$ be an eigenpair of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then, we have $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p},$ which can be written as
 $Au+B^{T}v=\lambda(\Lambda_{1}+sA)u+s\lambda B^{T}v,$ (4.2a)
@@
 Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0.$ Let $(\lambda,{\bf p}=[u^{T},v^{T},w^{T}]^{T})$ be an eigenpair of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then, the following holds
-  1. 1.
+  1. $(1)$
     $u\neq 0.$
-  2. 2.
+  2. $(2)$
     When $\lambda$ is real, $\lambda>0.$
-  3. 3.
+  3. $(3)$
     $\Re(\mu)>0,$ where $\mu=\lambda/(1-s\lambda).$
 Proof.
-  1. 1.
-    Let $u=0$.
+$(1)$ Let $u=0$.
 Then by (4.2a), we get $(1-s\lambda)B^{T}v=0,$ which shows that $v=0,$ as $B$ has full row rank and $\lambda\neq\frac{1}{s}.$ Furthermore, when combined with (4.2c) leads to $w=0,$ as $\Lambda_{3}$ is SPD.
@@
 Hence, $u\neq 0.$
-  2. 2.
-    Let ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ be an eigenvector corresponding to the real eigenvalue $\lambda.$ Then, we have ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}{\bf p}=\lambda{\bf p},$ or, $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p}.$ Consequently, $\lambda$ can be written as
+$(2)$ Let ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ be an eigenvector corresponding to the real eigenvalue $\lambda.$ Then, we have ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}{\bf p}=\lambda{\bf p},$ or, $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p}.$ Consequently, $\lambda$ can be written as
 $\lambda$ $=$ $\frac{{\bf p}^{T}\mathcal{A}{\bf p}+{\bf p}^{T}\mathcal{A}^{T}\bf p}{2({\bf p}^{T}{\mathscr{P}}_{PESS}{\bf p}+{\bf p}^{T}{\mathscr{P}}_{PESS}^{T}\bf p)}$
@@
 The last inequality follows from the assumptions that $A,\Lambda_{1},\Lambda_{2}$, and $\Lambda_{3}$ are SPD matrices.
-  3. 3.
-    The system of linear equations between (4.2a)-(4.2c) can be reformulated as
+$(3)$ The system of linear equations (4.2a)-(4.2c) can be reformulated as
 $\left\{\begin{array}[]{lcl}Au+B^{T}v=(\lambda/(1-s\lambda))\Lambda_{1}u,&\\ -Bu-C^{T}w=(\lambda/(1-s\lambda))\Lambda_{2}v,&\\ Cv=(\lambda/(1-s\lambda))\Lambda_{3}w.\end{array}\right.$ (4.3)
@@
 From Proposition 4.1, it follows that when $\lambda$ is real, then $\mu>0.$ Furthermore, the values of $\lambda$ lies in $(0,1/s)$ for all $s>0$.
-In the following theorem, we provide sharper bounds for real eigenvalues of the PESS preconditioned matrix.
+In the following theorem, we provide sharper bounds for real eigenvalues of the PESS preconditioned matrix. Before that, we introduce the following notation:
+$\xi_{\max}:=\lambda_{\max}(\Lambda_{1}^{-1}A),\leavevmode\nobreak\ \xi_{\min}:=\lambda_{\min}(\Lambda_{1}^{-1}A),\leavevmode\nobreak\ \eta_{\max}:=\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T}),$ (4.4)
+$\leavevmode\nobreak\ \eta_{\min}:=\lambda_{\min}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \theta_{\max}:=\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C).$ (4.5)
 Theorem 4.2.
 Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0.$ Suppose $\lambda$ is a real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then
-$\lambda\in\left(0,\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}\right].$ (4.4)
+$\lambda\in\left(0,\frac{\xi_{\max}}{1+s\xi_{\max}}\right].$ (4.6)
 Proof.
@@
 Since $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$ is similar to $\Lambda_{1}^{-1}A$ and the function $f:{\mathbb{R}}\rightarrow{\mathbb{R}}$ defined by $f(p):=p/(1+sp),$ where $s>0,$ is monotonically increasing in $p>0$, we obtain the following bound for $\lambda:$
-$\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\min}(\Lambda_{1}^{-1}A)}\leq\lambda\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}.$ (4.7)
+$\frac{\xi_{\min}}{1+s\xi_{\min}}\leq\lambda\leq\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.9)
 […]
@@
 Combining (4.15) and (4.16), we get the following bounds for $\lambda$
-$0<\lambda^{\pm}<\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}.$ (4.15)
+$0<\lambda^{\pm}<\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.17)
 Next, consider the case $Cv\neq 0.$ Then from (4.2c), we get $w=\frac{1}{\mu}\Lambda_{3}^{-1}Cv.$ Substituting the value of $w$ in (4.2b), we obtain
@@
 Now, we assume that
-$\lambda>\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}.$ (4.17)
-Since by Remark 4.2, we get $01-s\lambda>$ 0 and the above assertion yields $\mu>\lambda_{\max}(\Lambda_{1}^{-1}A).$ Hence, the matrix $\mu\Lambda_{1}-A$ is nonsingular and (4.2a) gives
+$\lambda>\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.19)
+Since by Remark 4.2, we get $01-s\lambda>$ 0 and the above assertion yields $\mu>\xi_{\max}.$ Hence, the matrix $\mu\Lambda_{1}-A$ is nonsingular and (4.2a) gives
 $u=\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v.$ (4.20)
@@
 Consequently, the matrix $B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}$ is also a SPD matrix and this implies $v^{T}B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v>0,$ for all $v\not=0.$ This leads to a contradiction to (4.22) as $v^{T}C^{T}\Lambda_{3}^{-1}Cv\geq 0.$ Hence, the assumption (4.19) is not true and we get
-$\lambda\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}.$ (4.21)
-Again, from (1) of Proposition 4.1, we have $\lambda>0.$ Combining the above together with (4.7) and (4.15), we obtained the desired bounds in (4.4) for $\lambda.$ Hence, the proof is concluded.
-$\blacksquare$ Since the preconditioners ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{EGSS}$ are special cases of the PESS preconditioner, next, we obtain refined bounds for real eigenvalues of the SS and EGSS preconditioned matrices from Theorem 4.2.
+$\lambda\leq\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.23)
+Again, from ($(1)$) of Proposition 4.1, we have $\lambda>0.$ Combining the above together with (4.9) and (4.17), we obtained the desired bounds in (4.6) for $\lambda.$ Hence, the proof is concluded. ∎
+Since the preconditioners ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{EGSS}$ are special cases of the PESS preconditioner, next, we obtain refined bounds for real eigenvalues of the SS and EGSS preconditioned matrices from Theorem 4.2.
 Corollary 4.1.
 Let ${\mathscr{P}}_{SS}$ be defined as in (1.10) with $\alpha>0$ and let $\lambda$ be an real eigenvalue of the SS preconditioned matrix ${\mathscr{P}}_{SS}^{-1}\mathcal{A}.$ Then
-$\lambda\in\left(0,\frac{2\lambda_{\max}(A)}{\alpha+\lambda_{\max}(A)}\right].$ (4.22)
+$\lambda\in\left(0,\frac{2\kappa_{\max}}{\alpha+\kappa_{\max}}\right],$ (4.24)
+where $\kappa_{\max}=\lambda_{\max}(A).$
 Proof.
 Since ${\mathscr{P}}_{SS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, the bounds in (4.24) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha I$ and $s=\frac{1}{2}$ in Theorem 4.2.
-$\blacksquare$ In the next result, we discuss bounds for the real eigenvalues of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$.
+$\blacksquare$
+In the next result, we discuss bounds for the real eigenvalues of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$.
 Corollary 4.2.
 Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.11) with $\alpha>0$ and let $\lambda$ be an real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}.$ Then,
-$\lambda\in\left(0,\frac{2\lambda_{\max}(P^{-1}A)}{\alpha+\lambda_{\max}(P^{-1}A)}\right].$ (4.23)
+$\lambda\in\left(0,\frac{2\tilde{\kappa}_{\max}}{\alpha+\tilde{\kappa}_{\max}}\right],$ (4.25)
+where $\tilde{\kappa}_{\max}=\lambda_{\max}(P^{-1}A).$
 Proof.
 Since ${\mathscr{P}}_{EGSS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, and the bounds are in (4.25) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha P$ and $s=\frac{1}{2}$ in Theorem 4.2.
-$\blacksquare$ The following result shows the bounds when $\lambda$ is a non-real eigenvalue.
+$\blacksquare$
+The following result shows the bounds when $\lambda$ is a non-real eigenvalue.
 Theorem 4.3.
 Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices.
-Let $s>0$ and $\lambda$ be a non-real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0.$ Suppose ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the eigenvector corresponding to $\lambda.$ Then the following holds $(1).$   If $Cv=0,$ then $\lambda$ satisfies
-$\frac{\lambda_{\min}(\Lambda^{-1}A)}{2+s\lambda_{\min}(\Lambda_{1}^{-1}A)}\leq|\lambda|\leq\sqrt{\frac{\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}{1+s\lambda_{\min}(\Lambda_{1}^{-1}A)+s^{2}\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}}.$
-$(2).$   If $Cv\neq 0,$ then real and imaginary part of $\lambda/(1-s\lambda)$ satisfies
-$\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)\,\lambda_{\min}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}{2\left(\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)+\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)\right)}\leq\Re\left(\frac{\lambda}{1-s\lambda}\right)\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2},$
-$\left|\Im\left(\frac{\lambda}{1-s\lambda}\right)\right|\leq\sqrt{\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}.$
+Let $s>0$ and $\lambda$ be a non-real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0.$ Suppose ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the eigenvector corresponding to $\lambda.$ Then the following holds
+  1. $(1)$
+    If $Cv=0,$ then $\lambda$ satisfies
+    $\frac{\xi_{\min}}{2+s\xi_{\min}}\leq|\lambda|\leq\sqrt{\frac{\eta_{\max}}{1+s\xi_{\min}+s^{2}\eta_{\max}}}.$
+  2. $(2)$
+    If $Cv\neq 0,$ then real and imaginary part of $\lambda/(1-s\lambda)$ satisfies
+    $\frac{\xi_{\min}\,\eta_{\min}}{2\left(\xi_{\max}^{2}+\eta_{\max}+\theta_{\max}\right)}\leq\Re\left(\frac{\lambda}{1-s\lambda}\right)\leq\frac{\xi_{\max}}{2}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \left|\Im\left(\frac{\lambda}{1-s\lambda}\right)\right|\leq\sqrt{\eta_{\max}+\theta_{\max}},$
+where $\xi_{\max},\xi_{\min},\eta_{\max},\eta_{\min}$ and $\theta_{\max}$ is defined as in (4.4) and (4.5).
 Proof.
-$(1).$   Let $\lambda$ be an eigenvalue of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0$ and the corresponding eigenvector is ${\bf p}=[u^{T},v^{T},w^{T}]^{T}.$ Then, the system of linear equations in (4.2a)- (4.2c) holds.
-We assert that, $v\neq 0,$ otherwise from (4.2a), we get $Au=\lambda(\Lambda_{1}+sA)u,$ which further implies $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}\bar{u}=\lambda\bar{u},$ where $\bar{u}=(\Lambda_{1}+sA)^{\frac{1}{2}}u.$ Since, $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}$ is a SPD matrix and by Lemma 4.1, $u\neq 0,$ and hence $\bar{u}\neq 0,$ which implies $\lambda$ is real, thus a contradiction to $\Im(\lambda)\neq 0.$ Initially, we consider the case $Cv=0.$ The aforementioned discussion suggests that $Au\neq\lambda(\Lambda_{1}+sA)u$ for any $u\neq 0.$ Consequently, $(\mu\Lambda_{1}-A)$ is nonsingular.
+$(1)$   Let $\lambda$ be an eigenvalue of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0$ and the corresponding eigenvector is ${\bf p}=[u^{T},v^{T},w^{T}]^{T}.$ Then, the system of linear equations in (4.2a)- (4.2c) holds.
+We assert that, $v\neq 0,$ otherwise from (4.2a), we get $Au=\lambda(\Lambda_{1}+sA)u,$ which further implies $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}\bar{u}=\lambda\bar{u},$ where $\bar{u}=(\Lambda_{1}+sA)^{\frac{1}{2}}u.$ Since, $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}$ is a SPD matrix and by Lemma 4.1, $u\neq 0,$ and hence $\bar{u}\neq 0,$ which implies $\lambda$ is real, thus a contradiction to $\Im(\lambda)\neq 0.$
+Initially, we consider the case $Cv=0.$ The aforementioned discussion suggests that $Au\neq\lambda(\Lambda_{1}+sA)u$ for any $u\neq 0.$ Consequently, $(\mu\Lambda_{1}-A)$ is nonsingular.
 Again, since $Cv=0,$ from (4.2c), and similar to (4.10)-(4.12), we obtain quadratic equation (4.13) in $\lambda$, which has complex solutions
@@
 $\Psi_{1}(p,q,s)=\frac{q}{s^{2}q+sp+1},$
-where $p,q,s>0.$ Then $\Psi_{1}$ is monotonically increasing in the argument $q>0$ while decreasing in the argument $p>0.$ Due to the fact that $q\leq\lambda_{\max}(\Lambda_{1}^{-1}B^{T}\Lambda_{2}^{-1}B)=\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T}),$ we get the following bounds
-$|\lambda^{\pm i}|^{2}$ $=$ $\Psi_{1}(p,q,s)\leq\Psi_{1}(\lambda_{\min}(\Lambda_{1}^{-1}A),\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T}),s)$ (4.26)
-$=$ $\frac{\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}{1+s\lambda_{\min}(\Lambda_{1}^{-1}A)+s^{2}\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}.$
+where $p,q,s>0.$ Then $\Psi_{1}$ is monotonically increasing in the argument $q>0$ while decreasing in the argument $p>0.$ Due to the fact that $q\leq\lambda_{\max}(\Lambda_{1}^{-1}B^{T}\Lambda_{2}^{-1}B)=\eta_{\max},$ we get the following bounds
+$|\lambda^{\pm i}|^{2}=\Psi_{1}(p,q,s)$$\leq$ $\Psi_{1}(\xi_{\min},\eta_{\max},s)=\frac{\eta_{\max}}{1+s\xi_{\min}+s^{2}\eta_{\max}}.$ (4.28)
 Again, considering the condition $q>(p^{2}/4),$ (4.27) adheres to the inequality $(p/(2+sp))\leq|\lambda^{\pm i}|.$ Notably, the function $\Psi_{2}:{\mathbb{R}}\longrightarrow{\mathbb{R}}$ defined by $\Psi_{2}(p)=p/(2+sp),$ where $p,s>0,$ is monotonically increasing in the argument $p>0,$ we obtain the following bound
-$|\lambda^{\pm i}|=\Psi_{2}(p)\geq\Psi_{2}(\lambda_{\min}(\Lambda_{1}^{-1}A))=\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)}{2+s\lambda_{\min}(\Lambda_{1}^{-1}A)}.$ (4.27)
-Hence, combining (4.26) and (4.27), proof for the part $(1).$ follows.
-$(2).$   Next, consider the case $Cv\neq 0,$ then (4.2c) yields $w=\frac{1}{\mu}\Lambda_{3}^{-1}Cv,$ where $\mu=\lambda/(1-s\lambda).$ Substituting this in (4.2b), we get
+$|\lambda^{\pm i}|=\Psi_{2}(p)\geq\Psi_{2}(\xi_{\min})=\frac{\xi_{\min}}{2+s\xi_{\min}}.$ (4.29)
+Hence, combining (4.28) and (4.29), proof for the part $(1)$ follows.
+$(2)$   Next, consider the case $Cv\neq 0,$ then (4.2c) yields $w=\frac{1}{\mu}\Lambda_{3}^{-1}Cv,$ where $\mu=\lambda/(1-s\lambda).$ Substituting this in (4.2b), we get
 $Bu+\frac{1}{\mu}C^{T}\Lambda_{3}^{-1}Cv=-\mu\Lambda_{2}v.$ (4.30)
@@
 Therefore, we get
-$|\Im(\mu)|\leq\sqrt{\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}.$ (4.35)
-Furthermore, notice that $\Theta_{1}\geq\left(\Re(\mu)-\lambda_{\max}(\Lambda_{1}^{-1}A)\right)\Theta_{2},$ then from (4.33), we deduce
-$-\Re(\mu)\geq\left(\Re(\mu)-\lambda_{\max}(\Lambda_{1}^{-1}A)\right)\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\Theta_{2}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v}+\frac{\Re(\mu)}{|\mu|^{2}}\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}.$ (4.36)
+$|\Im(\mu)|\leq\sqrt{\eta_{\max}+\theta_{\max}}.$ (4.37)
+Furthermore, notice that $\Theta_{1}\geq\left(\Re(\mu)-\xi_{\max}\right)\Theta_{2},$ then from (4.35), we deduce
+$-\Re(\mu)\geq\left(\Re(\mu)-\xi_{\max}\right)\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\Theta_{2}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v}+\frac{\Re(\mu)}{|\mu|^{2}}\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}.$ (4.38)
 Using (4.36) to (4.38), we get
-$2\Re(\mu)\leq\lambda_{\max}(\Lambda_{1}^{-1}A)-\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{|\mu|^{2}}\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}\leq\lambda_{\max}(\Lambda_{1}^{-1}A).$ (4.37)
+$2\Re(\mu)\leq\xi_{\max}-\frac{\xi_{\max}}{|\mu|^{2}}\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}\leq\xi_{\max}.$ (4.39)
 Hence, (4.39) yields the following bound
-$\Re(\mu)\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2}.$ (4.38)
+$\Re(\mu)\leq\frac{\xi_{\max}}{2}.$ (4.40)
 On the other side, from (4.35) and (4.36), we deduce that
-$2\Re(\mu)=\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V(\Re(\mu)\Theta_{2}-\Theta_{1})V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v}\geq\lambda_{\min}(\Lambda_{1}^{-1}A)\,\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\mathcal{X}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v},$ (4.39)
+$2\Re(\mu)=\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V(\Re(\mu)\Theta_{2}-\Theta_{1})V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v}\geq\xi_{\min}\,\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\mathcal{X}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v},$ (4.41)
 where ${\displaystyle{\mathcal{X}=\mathrm{diag}\left(\frac{1}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right).}}$ Note that
@@
 Now, using Proposition 4.1 and (4.40), we get
-$-\theta_{i}<\Re(\mu)-\theta_{i}\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2}-\theta_{i},$
-which further yields $(\Re(\mu)-\theta_{i})^{2}\leq\max\left\{\left(\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2}-\theta_{i}\right)^{2},\theta_{i}^{2}\right\}.$ Hence, we get
-$\displaystyle{\max_{i}}(\Re(\mu)-\theta_{i})^{2}$ $=$ $\displaystyle{\max_{i}}\{(\Re(\mu)-\lambda_{\min}(\Lambda_{1}^{-1}A))^{2},(\Re(\mu)-\lambda_{\max}(\Lambda_{1}^{-1}A))^{2}\}$ (4.40)
-$\leq$ $\displaystyle{\max_{i}}\left\{\left(\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2}-\lambda_{\min}(\Lambda_{1}^{-1}A)\right)^{2},\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)\right\}$
-$=$ $\lambda_{\max}^{2}(\Lambda_{1}^{-1}A).$
+$-\theta_{i}<\Re(\mu)-\theta_{i}\leq\frac{\xi_{\max}}{2}-\theta_{i},$
+which further yields $(\Re(\mu)-\theta_{i})^{2}\leq\max\left\{\left(\frac{\xi_{\max}}{2}-\theta_{i}\right)^{2},\theta_{i}^{2}\right\}.$ Hence, we get
+$\displaystyle{\max_{i}}(\Re(\mu)-\theta_{i})^{2}$ $=$ $\displaystyle{\max_{i}}\{(\Re(\mu)-\xi_{\min})^{2},(\Re(\mu)-\xi_{\max})^{2}\}$ (4.42)
+$\leq$ $\displaystyle{\max_{i}}\left\{\left(\frac{\xi_{\max}}{2}-\xi_{\min}\right)^{2},\xi^{2}_{\max}\right\}=\xi^{2}_{\max}.$
 Therefore, by (4.42) and from the bound in (4.37), we obtain
-$\mathcal{X}\geq\frac{I}{\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)+\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}$ (4.41)
+$\mathcal{X}\geq\frac{I}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}$ (4.43)
 Combining (4.41) and (4.43) leads to the following bounds
-$2\Re(\mu)$$\geq$ $\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)}{\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)+\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}\frac{v^{*}B\Lambda_{1}^{-1}B^{T}v}{v^{*}\Lambda_{2}v}$ (4.42)
-$\geq$ $\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)\lambda_{\min}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}{\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)+\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}.$
-Hence, the proof of part $(2).$ follows by merging the inequalities of (4.35) and (4.42).
-$\blacksquare$ Utilizing the established bounds in Theorem 4.3, we derive the subsequent estimations for non-real eigenvalues of SS and EGSS preconditioned matrices.
+$2\Re(\mu)$$\geq$ $\frac{\xi_{\min}}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}\frac{v^{*}B\Lambda_{1}^{-1}B^{T}v}{v^{*}\Lambda_{2}v}\geq\frac{\xi_{\min}\eta_{\min}}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}.$ (4.44)
+Hence, the proof of part $(2)$ follows by merging the inequalities of (4.37) and (4.44). $\blacksquare$
+Remark 4.3.
+From the bounds in $(2)$ of Theorem 4.3, we obtain
+${\frac{1}{|\tau|^{2}}\left(\frac{\xi_{\min}\eta_{\min}}{2(\xi_{\max}^{2}+\eta_{\max}+\theta_{\max})}+\frac{1}{s}\right)\leq\frac{1}{s}-\Re(\lambda)\leq\frac{1}{|\tau|^{2}}(\frac{\xi_{\max}}{2}+\frac{1}{s})}$
+and
+$|\Im(\lambda)|\leq\frac{1}{s\tau\bar{\tau}}\sqrt{\eta_{\max}+\theta_{\max}},$
+where $\tau=s\mu+1.$ Thus, the real and imaginary parts of the eigenvalues of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ cluster better as $s$ grows.
+Therefore, the PESS preconditioner can accelerate the rate of convergence of the Krylov subspace process, like $GMRES$.
+Utilizing the established bounds in Theorem 4.3, we derive the subsequent estimations for non-real eigenvalues of SS and EGSS preconditioned matrices.
 Corollary 4.3.
 Let ${\mathscr{P}}_{SS}$ be defined as in (1.10) with $\alpha>0$ and let $\lambda$ be a non-real eigenvalue of the SS preconditioned matrix ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector.
-Then, $(1).$   If $Cv=0,$ then $\lambda$ satisfies
-$\frac{2\lambda_{\min}(A)}{2\alpha+\lambda_{\min}(A)}\leq|\lambda|\leq\sqrt{\frac{4\lambda_{\max}(BB^{T})}{\alpha^{2}+\alpha\lambda_{\min}(A)+\lambda_{\max}(BB^{T})}}.$
-$(2).$    If $Cv\neq 0,$ the real and imaginary part of $\lambda/(2-\lambda)$ satisfies
-[…]
-$\left|\Im\left(\frac{\lambda}{2-\lambda}\right)\right|\leq\frac{\sqrt{\lambda_{\max}(BB^{T})+\lambda_{\max}(C^{T}C)}}{\alpha}.$
+Then,
+$(1)$   If $Cv=0,$ $\lambda$ satisfies
+$\frac{2\kappa_{\min}}{2\alpha+\kappa_{\min}}\leq|\lambda|\leq\sqrt{\frac{4\tau_{\max}}{\alpha^{2}+\alpha\kappa_{\min}+\tau_{\max}}}.$
+$(2)$    If $Cv\neq 0,$ the real and imaginary part of $\lambda/(2-\lambda)$ satisfies
+$\frac{\kappa_{\min}\,\tau_{\min}}{2\alpha\left(\kappa^{2}_{\max}+\tau_{\max}+\beta_{\max}\right)}\leq\Re\left(\frac{\lambda}{2-\lambda}\right)\leq\frac{\kappa_{\max}}{2\alpha}\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \left|\Im\left(\frac{\lambda}{2-\lambda}\right)\right|\leq\frac{\sqrt{\tau_{\max}+\beta_{\max}}}{\alpha},$
+where $\kappa_{\min}=\lambda_{\min}(A),$ $\tau_{\max}=\lambda_{\max}(BB^{T}),$ $\tau_{\min}=\lambda_{\min}(BB^{T})$ and $\beta_{\max}=\lambda_{\max}(C^{T}C).$
 Proof.
@@
 Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.11), and let $P,Q$ and $W$ are SPD matrices with $\alpha,\beta,\gamma>0.$ Let $\lambda$ be a non-real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector.
-Then $(1).$ If $Cv=0,$ then $\lambda$ satisfies
-$\frac{2\lambda_{\min}(P^{-1}A)}{2\alpha+\lambda_{\min}(P^{-1}A)}\leq|\lambda|\leq\sqrt{\frac{4\lambda_{\max}(Q^{-1}BP^{-1}B^{T})}{\alpha\beta+\beta\lambda_{\min}(P^{-1}A)+\lambda_{\max}(Q^{-1}BP^{-1}B^{T})}}.$
-$(2).$   If $Cv\neq 0,$ then the real and imaginary part of $\lambda/(2-\lambda)$ satisfies
-$\frac{\lambda_{\min}(P^{-1}A)\,\lambda_{\min}(Q^{-1}BP^{-1}B^{T})}{2\left(\beta\lambda_{\max}^{2}(P^{-1}A)+\alpha\lambda_{\max}(Q^{-1}BP^{-1}B^{T})+\frac{\alpha^{2}}{\gamma}\lambda_{\max}(Q^{-1}C^{T}W^{-1}C)\right)}$
-[…]
-$\left|\Im\left(\frac{\lambda}{2-\lambda}\right)\right|\leq\sqrt{\frac{1}{\alpha\beta}\lambda_{\max}(Q^{-1}BP^{-1}B^{T})+\frac{1}{\beta\gamma}\lambda_{\max}(\Lambda_{2}^{-1}C^{T}W^{-1}C)}.$
+Then $(1)$ If $Cv=0,$ $\lambda$ satisfies
+$\frac{2\tilde{\kappa}_{\min}}{2\alpha+\tilde{\kappa}_{\min}}\leq|\lambda|\leq\sqrt{\frac{4\tilde{\tau}_{\max}}{\alpha\beta+\beta\tilde{\kappa}_{\min}+\tilde{\tau}_{\max}}}.$
+$(2)$   If $Cv\neq 0,$ the real and imaginary part of $\lambda/(2-\lambda)$ satisfies
+$\frac{\tilde{\kappa}_{\min}\,\tilde{\tau}_{\min}}{2\left(\beta\tilde{\kappa}^{2}_{\max}+\alpha\tilde{\tau}_{\max}+(\alpha^{2}/\gamma)\tilde{\beta}_{\max}\right)}\leq\Re\left(\frac{\lambda}{2-\lambda}\right)\leq\frac{\tilde{\kappa}_{\max}}{2\alpha}\leavevmode\nobreak\ \mbox{and}$
+$\left|\Im\left(\frac{\lambda}{2-\lambda}\right)\right|\leq\sqrt{\frac{1}{\alpha\beta}\tilde{\tau}_{\max}+\frac{1}{\beta\gamma}\tilde{\beta}_{\max}},$
+where $\tilde{\kappa}_{\min}=\lambda_{\min}(P^{-1}A),$ $\tilde{\tau}_{\max}=\lambda_{\max}(Q^{-1}BP^{-1}B^{T}),$ $\tilde{\tau}_{\min}=\lambda_{\min}(Q^{-1}BP^{-1}B^{T})$ and $\tilde{\beta}_{\max}=\lambda_{\max}(C^{T}C).$
 Proof.
@@
 $\blacksquare$
+5 Local PESS (LPESS) preconditioner
+To enhance the efficiency of the PESS preconditioner, in this section, we propose a relaxed version of the PESS preconditioner by incorporating the concept of RSS preconditioner [9].
+By omitting the term $\Lambda_{1}$ from the $(1,1)$-block of ${\mathscr{P}}_{PESS}$, we present the local PESS (LPESS) preconditioner, denoted as ${\mathscr{P}}_{LPESS}$, defined as follows:
+${\mathscr{P}}_{LPESS}:=\left[\begin{matrix}A&sB^{T}&0\\ -sB&\Lambda_{2}&-sC^{T}\\ 0&sC&\Lambda_{3}\end{matrix}\right].$ (5.1)
+The implementation of the LPESS preconditioner is similar to the Algorithm 1.
+However, there is one modification in step $3,$ i.e., we need to solve a linear subsystem $(sA+s^{2}B^{T}\widehat{X}^{-1}B)w_{1}=v$ instead of $\widehat{A}w_{1}=v.$
+To illustrate the efficiency of the LPESS preconditioner, we study the spectral distribution of the preconditioned matrix ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}.$ To achieve this, we consider the following decomposition of the matrices $\mathcal{A}$ and ${\mathscr{P}}_{LPESS}$ as follows:
+$\mathcal{A}=\mathfrak{L}\mathfrak{D}\mathfrak{U}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ {\mathscr{P}}_{LPESS}=\mathfrak{L}\widetilde{\mathfrak{D}}\mathfrak{U},$ (5.2)
+where
+$\mathfrak{L}=\left[\begin{matrix}I&0&0\\ -BA^{-1}&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{U}=\left[\begin{matrix}I&A^{-1}B^{T}&0\\ 0&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{D}=\left[\begin{matrix}A&0&0\\ 0&Q&-C^{T}\\ 0&C&0\end{matrix}\right],\widetilde{\mathfrak{D}}=\left[\begin{matrix}sA&0&0\\ 0&\Lambda_{2}+sQ&-C^{T}\\ 0&C&\Lambda_{3}\end{matrix}\right],$
+and $Q=BA^{-1}B^{T}.$ Using the decomposition in (5.2), we have
+${\mathscr{P}}^{-1}_{LPESS}\mathcal{A}=\mathfrak{U}^{-1}\left[\begin{matrix}s^{-1}I&0\\ 0&M^{-1}K\end{matrix}\right]\mathfrak{U},$ (5.3)
+where $M=\left[\begin{matrix}\Lambda_{2}+sQ&-sC^{T}\\ sC&\Lambda_{3}\end{matrix}\right]$ and $K=\left[\begin{matrix}Q&-C^{T}\\ C&0\end{matrix}\right].$ Since, ${\mathscr{P}}^{-1}_{LPESS}\mathcal{A}$ is similar to $\left[\begin{matrix}s^{-1}I&0\\ 0&M^{-1}K\end{matrix}\right],$ ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has eigenvalue $\frac{1}{s}$ with multiplicity $n$ and the remaining eigenvalues satisfies the generalized eigenvalue problem $Kp=\lambda Mp.$
+Theorem 5.1.
+Let $A\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}$ and $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices.
+Assume that $s>0,$ then LPESS preconditioned matrix ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has $n$ eigenvalue equal to $\frac{1}{s}.$ The remaining $m+p$ eigenvalues satisfies the following:
+  1. $(1)$
+    real eigenvalues located in the interval
+    $\Big{[}\min\Big{\{}\frac{\vartheta_{\min}}{1+s\vartheta_{\min}},\frac{\theta_{\min}}{\vartheta_{\max}+s\theta_{\min}}\Big{\}},\frac{\vartheta_{\max}}{1+s\vartheta_{\max}}\Big{]}.$
+  2. $(2)$
+    If $\lambda$ is any non-real eigenvalue (i.e., $\Im(\lambda)\neq 0$), then
+    $\frac{\vartheta_{\min}}{2+s\vartheta_{\min}}\leq|\lambda|\leq\sqrt{\frac{\theta_{\max}}{1+s\vartheta_{\min}+s^{2}\theta_{\max}}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \frac{1}{s(1+s\sqrt{\theta_{\max}})}\leq|\lambda-\frac{1}{s}|\leq\frac{2}{s(2+s\vartheta_{\min})},$
+where $\vartheta_{\max}:=\lambda_{\max}(\Lambda_{2}^{-1}Q),$ $\vartheta_{\min}:=\lambda_{\min}(\Lambda_{2}^{-1}Q)$ and $\theta_{\min}:=\lambda_{\min}(\Lambda_{3}^{-1}C\Lambda_{1}^{-1}C^{T}).$
+Proof.
+Observe that,
+$M^{-1}K=\mathcal{F}^{-1}\widetilde{M}^{-1}\widetilde{K}\mathcal{F},$ (5.4)
+where $\mathcal{F}=\left[\begin{matrix}I&0\\ 0&-I\end{matrix}\right],$ $\widetilde{M}=\left[\begin{matrix}\Lambda_{2}+sQ&C^{T}\\ -C&\Lambda_{3}\end{matrix}\right]$ and $\widetilde{K}=\left[\begin{matrix}Q&C^{T}\\ -C&0\end{matrix}\right].$ Hence, $M^{-1}K$ is similar to $\widetilde{M}^{-1}\widetilde{K}.$ Now, $\widetilde{M}^{-1}\widetilde{K}$ is in the form of the preconditioned matrix discussed in [28].
+Therefore, by applying the Corollary 4.1 and Theorem 4.3 of [28], we obtain the desired bounds in $(1)$ and $(2).$blacksquare
 6 Numerical experiments
-In this section, we conduct a few numerical experiments to showcase the superiority and efficiency of the proposed PESS preconditioner over the existing preconditioners to enhance the convergence speed of the Krylov subspace iterative process to solve three-by-three block SPPs.
-Our study involves a comparative analysis among GMRES process and the PGMRES process employing the proposed PESS preconditioner ${\mathscr{P}}_{PESS}$ and the existing preconditioners ${\mathscr{P}}_{BD}$, ${\mathscr{P}}_{IBD}$, ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS}$ and ${\mathscr{P}}_{EGSS}.$ The numerical results are reported from the aspect of iteration counts (abbreviated as “IT”) and elapsed CPU times in seconds (abbreviated as “CPU”).
-Each subsystem featured in Algorithm 1 is precisely solved by applying the Cholesky factorization of the coefficient matrices.
+In this section, we conduct a few numerical experiments to showcase the superiority and efficiency of the proposed PESS and LPESS preconditioners over the existing preconditioners to enhance the convergence speed of the Krylov subspace iterative process to solve three-by-three block SPPs.
+Our study involves a comparative analysis among GMRES process and the PGMRES process employing the proposed preconditioners ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS},$ and the existing preconditioners ${\mathscr{P}}_{BD}$, ${\mathscr{P}}_{IBD}$, ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS}$ and ${\mathscr{P}}_{EGSS}$ (baseline preconditioners).
+The numerical results are reported from the aspect of iteration counts (abbreviated as “IT”) and elapsed CPU times in seconds (abbreviated as “CPU”).
+Each subsystem involving $\widehat{A}$ and $\widehat{X}$ featured in Algorithm 1 is precisely solved by applying the Cholesky factorization of the coefficient matrices.
 For all iterative process, the initial guess vector is ${\bf u}_{0}=0\in{\mathbb{R}}^{n+m+p}$, and the termination criterion is
@@
 The vector ${\bf d}\in{\mathbb{R}}^{n+m+p}$ is chosen so that the exact solution of the system (1.1) is ${\bf u}_{*}=[1,1,\ldots,1]^{T}\in{\mathbb{R}}^{n+m+p}.$ All numerical tests are run in MATLAB (version R2023a) on a Windows 11 operating system with Intel(R) Core(TM) i7-10700 CPU, 2.90GHz, 16 GB memory.
+Before proceeding, we discuss the optimal choice for the parameters of PESS preconditioner.
+We can rewrite the PESS preconditioner as ${\mathscr{P}}_{PESS}=s\widetilde{{\mathscr{P}}}_{PESS},$ where $\widetilde{{\mathscr{P}}}_{PESS}=\left[\begin{matrix}\frac{1}{s}\Lambda_{1}+A&B^{T}&0\\ -B&\frac{1}{s}\Lambda_{2}&-C^{T}\\ 0&C&\frac{1}{s}\Lambda_{3}\end{matrix}\right].$ Since, the pre-factor $s$ has not much effect on the performance of PESS preconditioner, investigating the optimal parameters of ${\mathscr{P}}_{PESS}$ and $\widetilde{{\mathscr{P}}}_{PESS}$ are equivalent.
+As we know, a preconditioner performs efficiently when it is as close as possible to the coefficient matrix $\mathcal{A},$ observe that $\widetilde{{\mathscr{P}}}_{PESS}-\mathcal{A}={\displaystyle{\frac{1}{s}\Sigma}},$ which tends to a zero matrix as $s$ tends to positive infinity for fixed $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}.$ Thus, PESS encounters enhance efficiency for large value of $s.$ However, $s$ can not be too large as the coefficient matrix $\Lambda_{1}+sA+s^{2}B^{T}(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)^{-1}B$ of the linear subsystem in step $3$ of Algorithm 1 becomes very-ill conditioned.
+Similar investigations also hold for LPESS preconditioner.
+Nevertheless, in Figure 7, we show the adaptability of the PESS and LPESS preconditioners by varying the parameter $s.$
+Example 1.
+We consider the three-by-three block SPP (1.1) taken from [18], where
+$A=\left[\begin{matrix}I\otimes G+G\otimes I&0\\ 0&I\otimes G+G\otimes I\end{matrix}\right]\in{\mathbb{R}}^{2l^{2}\times 2l^{2}},\,\,B=\left[\begin{matrix}I\otimes F&F\otimes I\end{matrix}\right]\in{\mathbb{R}}^{l^{2}\times 2l^{2}},$ and $C=E\otimes F\in{\mathbb{R}}^{l^{2}\times l^{2}},$ where $G=\frac{1}{(l+1)^{2}}\,\mathrm{tridiag}(-1,2,-1)\in{\mathbb{R}}^{l\times l},\quad F=\frac{1}{l+1}\,\mathrm{tridiag}(0,1,-1)\in{\mathbb{R}}^{l\times l}$ , and $E=\mathrm{diag}(1,l+1,\ldots,l^{2}-l+1)\in{\mathbb{R}}^{l\times l}.$ Here, $E\otimes F$ denotes the Kronecker product of two matrices $E$ and $F,$ and $\mathrm{tridiag}(a,b,c)\in{\mathbb{R}}^{l\times l}$ denotes the $l\times l$ tridiagonal matrix with diagonal, subdiagonal, and superdiagonal entries are $a,b,$ and $c,$ respectively.
+For this problem, the size of the matrix $\mathcal{A}$ is $4l^{2}.$
 Table 2: Numerical results of GMRES, BD, IBD, SS, RSS, EGSS and PESS PGMRES process for Example 1.
@@
 Case I
-SS IT $5$ $4$ $4$ $4$ $4$
-CPU $0.2221$ $1.1664$ $6.6960$ $32.3304$ $106.8526$
-RES $084.7675e-$ 08 $078.7496e-$ 07 $077.1592e-$ 07 $076.2088e-$ 07 $075.5601e-$ 07
-RSS IT $5$ $4$ $4$ $4$ $4$
-CPU $0.2099$ $1.1297$ $6.5668$ $32.6876$ $109.8077$
-RES $083.2028e-$ 08 $079.5518e-$ 07 $077.9278e-$ 07 $076.9203e-$ 07 $076.2180e-$ 07
+SS IT $4$ $4$ $4$ $4$ $4$
+CPU $0.2415$ $1.3375$ $6.7059$ $33.6902$ $111.3118$
+RES $087.7528e-$ 08 $085.5120E-$ 08 $084.5134E-$ 08 $083.9157E-$ 08 $083.5073E-$ 08
+RSS IT $4$ $4$ $4$ $4$ $4$
+CPU $0.2776$ $1.0467$ $6.5207$ $32.4428$ $110.4587$
+RES $088.0898e-$ 08 $086.0033e-$ 08 $084.9839e-$ 08 $084.3510e-$ 08 $083.9097e-$ 08
 EGSS IT $4$ $4$ $4$ $4$ $4$
-CPU $0.2592$ $1.3542$ $7.9751$ $34.3676$ $125.4022$
-RES $081.8724e-$ 08 $081.7563e-$ 08 $081.7090e-$ 08 $081.6765e-$ 08 $081.6497e-$ 08
-${}^{\dagger}$ IT 2 2 2 2 2
-CPU $0.2078$ 0.8102 4.6363 21.1566 75.4631
+CPU $0.3061$ $1.2771$ $6.9552$ $32.8960$ $119.7763$
+RES $105.9583e-$ 10 $104.4128e-$ 10 $103.6626e-$ 10 $103.2009e-$ 10 $102.8807e-$ 10
+PESS IT 2 2 2 2 2
+CPU $0.2285$ $\bf 0.9854$ $\bf 4.8795$ $\bf 23.3687$ $\bf 79.9355$
 $s=12$ RES $073.1630e-$ 07 $072.3239e-$ 07 $071.9486e-$ 07 $071.7260e-$ 07 $071.5754e-$ 07
-${}^{\dagger}$ IT 2 2 2 2 2
-CPU $0.2289$ 0.8097 4.6984 21.1552 74.09290
-$s=13$ RES $072.6946e-$ 07 $071.9800e-$ 07 $071.6603e-$ 07 $071.4708e-$ 07 $071.3426e-$ 07
+LPESS IT 2 2 2 2 2
+CPU $0.2834$ $\bf 0.7989$ $\bf 4.5599$ $\bf 21.1664$ $\bf 72.8376$
+$s=12$ RES $073.1180e-$ 07 $072.2463e-$ 07 $071.8481e-$ 07 $071.6071e-$ 07 $071.4411e-$ 07
 Case II
-SS IT $10$ $9$ $9$ $9$ $9$
-CPU $0.2879$ $2.0186$ $12.7565$ $63.8265$ $211.2261$
-RES $071.4343e-$ 07 $077.5194e-$ 07 $076.0005e-$ 07 $075.2050e-$ 07 $074.6709e-$ 07
-RSS IT $9$ $9$ $9$ $9$ $9$
-CPU $0.1990$ $1.8589$ $12.7943$ $62.9284$ $211.9332$
-RES $077.8940e-$ 07 $075.4297e-$ 07 $074.5044e-$ 07 $073.8357e-$ 07 $073.3306e-$ 07
-EGSS IT $9$ $9$ $9$ $9$ $9$
-CPU $0.2130$ $1.4008$ $8.7647$ $41.7527$ $156.1426$
-RES $071.1981e-$ 07 $071.7824e-$ 07 $071.5671e-$ 07 $071.4056e-$ 07 $071.2878e-$ 07
-${}^{\dagger}$ IT $3$ $3$ $3$ $3$ $3$
-CPU $0.2344$ $0.9795$ 5.8979 28.3526 98.0878
-$s=12$ RES $081.0992e-$ 08 $082.4637e-$ 08 $084.8787e-$ 08 $087.2624e-$ 08 $088.4769e-$ 08
-${}^{\dagger}$ IT $3$ $3$ $3$ $3$ $3$
-CPU $0.1929$ $0.9965$ $5.9099$ $28.0443$ $101.0273$
-$s=13$ RES $098.6456e-$ 09 $081.9393e-$ 08 $083.8703e-$ 08 $085.8810e-$ 08 $087.0253e-$ 08
+SS IT $7$ $7$ $7$ $7$ $7$
+CPU 0.2780 $1.6189$ $10.6707$ $54.4493$ $184.6025$
+RES $076.7967e-$ 07 $074.9738e-$ 07 $074.1383e-$ 07 $073.6189e-$ 07 $073.2556e-$ 07
+RSS IT $7$ $7$ $7$ $7$ $7$
+CPU $0.3248$ $1.4644$ $10.4985$ $52.4559$ $177.1895$
+RES $075.2397e-$ 07 $073.6832e-$ 07 $072.7532e-$ 07 $072.1869e-$ 07 $071.8240e-$ 07
+EGSS IT $5$ $5$ $4$ $4$ $4$
+CPU $0.2588$ $1.3792$ $6.7546$ $34.8358$ $118.2298$
+RES $086.8491e-$ 08 $071.5239e-$ 07 $079.7099e-$ 07 $076.3466e-$ 07 $074.8159e-$ 07
+PESS IT $3$ $3$ $3$ $3$ $3$
+CPU $0.2971$ $1.1597$ $6.6682$ $31.2670$ $103.8616$
+$s=12$ RES $087.4100e-$ 08 $087.4642e-$ 08 $087.3327e-$ 08 $086.9803e-$ 08 $086.1920e-$ 08
+LPESS IT $3$ $3$ $3$ $3$ $3$
+CPU $0.2226$ $0.9905$ $6.1339$ $30.0975$ $98.6443$
+$s=12$ RES $091.1683e-$ 09 $092.3215e-$ 09 $093.1540e-$ 09 $093.3907e-$ 09 $093.0271e-$ 09
 […]
 […]
-Example 1.
-We consider the three-by-three block SPP (1.1) taken from [18], where
-$A=\left[\begin{matrix}I\otimes G+G\otimes I&0\\ 0&I\otimes G+G\otimes I\end{matrix}\right]\in{\mathbb{R}}^{2l^{2}\times 2l^{2}},\,\,B=\left[\begin{matrix}I\otimes F&F\otimes I\end{matrix}\right]\in{\mathbb{R}}^{l^{2}\times 2l^{2}},$ and $C=E\otimes F\in{\mathbb{R}}^{l^{2}\times l^{2}},$ where $G=\frac{1}{(l+1)^{2}}\,\mathrm{tridiag}(-1,2,-1)\in{\mathbb{R}}^{l\times l},\quad F=\frac{1}{l+1}\,\mathrm{tridiag}(0,1,-1)\in{\mathbb{R}}^{l\times l}$ , and $E=\mathrm{diag}(1,l+1,\ldots,l^{2}-l+1)\in{\mathbb{R}}^{l\times l}.$ Here, $E\otimes F$ denotes the Kronecker product of two matrices $E$ and $F,$ and $\mathrm{tridiag}(a,b,c)\in{\mathbb{R}}^{l\times l}$ denotes the $l\times l$ tridiagonal matrix with diagonal, subdiagonal, and superdiagonal entries are $a,b,$ and $c,$ respectively.
-For this problem, the size of the matrix $\mathcal{A}$ is $4l^{2}.$ We consider the parameter choices for the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS},{\mathscr{P}}_{EGSS}$ and ${\mathscr{P}}_{PESS}$ in two cases.
-  • In Case I: $\alpha=0.2$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=0.2,\beta=2,\gamma=0.002$ and $P=I,Q=I,W=I$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=\frac{1}{2}\alpha I,\Lambda_{2}=\frac{1}{2}\beta I,\Lambda_{3}=\frac{1}{2}\gamma I$ for ${\mathscr{P}}_{PESS}.$
-  • In Case II: $\alpha=2$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=2,\beta=2,\gamma=0.002$ and $P=I,Q=I,W=CC^{T}$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=\frac{1}{2}\alpha I,\Lambda_{2}=\frac{1}{2}\beta I,\Lambda_{3}=\frac{1}{2}\gamma CC^{T}$ for ${\mathscr{P}}_{PESS}.$
-The optimal values (experimentally) for $s$ in the range $[10,20]$ are determined to be $12$ and $13$ for minimal CPU times.
-Following [18], selection of $\hat{A}$ and $\hat{S}$ (SPD approximations of $A$ and $S,$ respectively) for IBD preconditioner are done as follows:
-$\hat{A}=LL^{T},\quad\hat{S}=\mathrm{diag}(B\hat{A}^{-1}B^{T}),$
+We consider the parameter choices for the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS},{\mathscr{P}}_{EGSS},{\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}$ in two cases.
+  • In Case I: $\alpha=0.1$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=0.1,\beta=1,\gamma=0.001$ and $P=I,Q=I,W=I$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=I,\Lambda_{2}=I,\Lambda_{3}=0.001I$ for ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}.$
+[…]
+The parameter selection in Case II is made as in [31].
+The optimal value (experimentally) for $s$ in the range $[10,20]$ is determined to be $12$ for minimal CPU times.
+Following [18], selection of $\widehat{A}$ and $\widehat{S}$ (SPD approximations of $A$ and $S,$ respectively) for IBD preconditioner are done as follows:
+$\widehat{A}=LL^{T},\quad\widehat{S}=\mathrm{diag}(B\widehat{A}^{-1}B^{T}),$
 where $L$ is the incomplete Cholesky factor of $A$ produced by the Matlab function:
@@
 (e) $l=80$
-Figure 1: Convergence curves for IT versus RES of PGMRES process employing BD, IBD, SS, RSS, EGSS, and PESS $(s=13)$ preconditioners in Case II for Example 1.
+Figure 1: Convergence curves for IT versus RES of PGMRES process employing BD, IBD, SS, RSS, EGSS, PESS and LPESS $(s=12)$ preconditioners in Case II for Example 1.
 (a) $\mathcal{A}$
@@
 (f) ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$
-(g) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=12$
-Figure 2: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=16$ for Example 1.
+(g) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=13$
+(h) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=13$
+Figure 2: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with $l=16$ for Example 1.
 Figure 3: Parameter $s$ vs condition number of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ for Example 1.
-Figure 1 illustrates convergence curves pertaining to preconditioners BD, IBD, SS, RSS, EGSS, and PESS ($s=13$) for Case II with $l=16,32,48,64$ and $80$ These curves depict the relationship between the relative residue at each iteration step (RES) and IT counts.
-To further illustrate the superiority of the PESS preconditioner, spectral distributions of $\mathcal{A}$ and the preconditioned matrices ${\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{RSS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for the Case II with $l=16$ and $s=12$ are displayed in Figure 2.
+Figure 1 illustrates convergence curves pertaining to preconditioners BD, IBD, SS, RSS, EGSS, PESS and LPESS ($s=12$) for Case II with $l=16,32,48,64$ and $80$ These curves depict the relationship between the relative residue at each iteration step (RES) and IT counts.
+To further illustrate the superiority of the PESS preconditioner, spectral distributions of $\mathcal{A}$ and the preconditioned matrices ${\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{RSS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for the Case II with $l=16$ and $s=13$ are displayed in Figure 2.
 To investigate the robustness of the proposed PESS preconditioner, we measure the condition number of the ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ which is for any nonsingular matrix $A$ defined by $\kappa(A):=\|A^{-1}\|_{2}\|A\|_{2}.$ In Figure 3, the influence of the parameter $s$ on the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ is depicted.
@@
   • From Table 2, it is observed that the GMRES has a very slow convergence speed.
-In both Cases I and II, our proposed PESS preconditioner outperforms all other compared existing preconditioners from the aspects of IT and CPU times.
-For example, in Case II with $l=80,$ our proposed PESS preconditioner is $71\%$, $30\%$, $54\%$, $54\%$, and $37\%$ more efficient than the existing BD, IBD, SS, RSS, and EGSS preconditioner, respectively.
+In both Cases I and II, our proposed PESS and LPESS preconditioners outperform all other compared existing preconditioners from the aspects of IT and CPU times.
+For example, in Case I with $l=80,$ our proposed PESS preconditioner is $76\%$, $43\%$, $28\%$, $27\%$, and $33\%$ more efficient than the existing BD, IBD, SS, RSS, and EGSS preconditioner, respectively.
+Moreover, LPESS preconditioners performs approximately $78\%,$ $48\%,$ $35\%,$ $34\%,$ and $39\%$ faster than BD, IBD, SS, RSS, and EGSS preconditioner, respectively.
 Similar patterns are noticed for $l=16,32,48$ and $64$.
-  • In Figure 1, it is evident that the PESS preconditioner has a faster convergence speed than the other baseline preconditioners when applied to the PGMRES process for all $l=16,32,48,64$ and $80$
-  • According to Figure 2, the spectrum of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ has better clustering properties than the baseline preconditioned matrices, consequently improves the computational performance of our proposed PESS PGMRES process.
-Moreover, the real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ are contained in the interval $(0,0.08330]$ and for ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ they are in $(0,1.9983],$ which are consistent with bounds of Theorem 4.2, Corollaries 3.1, and 3.2, respectively.
+  • In Figure 1, it is evident that the PESS and LPESS preconditioners have a faster convergence speed than the other baseline preconditioners when applied to the PGMRES process for all $l=16,32,48,64$ and $80$
+  • According to Figure 2, the spectrum of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LESS}^{-1}\mathcal{A}$ have better clustering properties than the baseline preconditioned matrices, consequently improves the computational efficiency of our proposed PESS and LPESS PGMRES processes.
+Moreover, the real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ are contained in the interval $(0,0.0714]$ and for ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ are in $(0,1.9991]$ and $(0,1],$ respectively, which are consistent with bounds of Theorem 4.2, Corollaries 3.1, and 3.2, respectively.
 For non-real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ obtained bounds are consistent with Theorem 4.3.
   • For $l=32,$ the computed condition numbers of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},$ and ${\mathscr{P}}_{IBD}^{-1}\mathcal{A}$ are $5.4289e+04,$ $9.5567e+09,$ and $9.5124e+09,$ respectively, which are very large.
-Whereas Figure 3 illustrate a decreasing trend in the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ (for instance, when $s=50$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=3.422$) with increasing values of $s.$ This observation highlights that ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ emerges as well-conditioned, consequently the solution obtained by the proposed PESS PGMRES process is more reliable and robust.
-The above discussions affirm that the proposed PESS preconditioner is efficient, robust, and better well-conditioned with respect to the baseline preconditioners.
+Whereas Figure 3 illustrate a decreasing trend in the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ (for instance, when $s=50$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=3.4221$) with increasing values of $s.$ This observation highlights that ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ emerges as well-conditioned, consequently the solution obtained by the proposed PESS PGMRES process is more reliable and robust.
+The above discussions affirm that the proposed PESS and LPESS preconditioner is efficient, robust, and better well-conditioned with respect to the baseline preconditioners.
 Figure 4: Relationship of norm error of solution with increasing noise percentage, employing proposed ${\mathscr{P}}_{PESS}$ for Case II with $s=12$ for $l=16,32,48,64$ and $80$ of Example 1.
@@
 In the following, we consider the perturbed counterpart $(\mathcal{A}+\Delta\mathcal{A})\tilde{\textbf{u}}=\textbf{d}$ of the three-by-three block SPP (1.1), where the perturbation matrix $\Delta\mathcal{A}$ has same structure as of $\mathcal{A}.$ We construct the perturbation matrix $\Delta\mathcal{A}$ by adding noise to the block matrices $B$ and $C$ of $\mathcal{A}$ as follows:
-$\Delta B=\varepsilon*N_{P}*std(B).*randn(m,n)\quad\text{and}\quad\Delta C=\varepsilon*N_{P}*std(C).*randn(p,m),$
-where $\varepsilon=10^{-4}$ and $N_{P}$ is the noise percentage and $std(B)$ and $std(C)$ are the standard deviation of $B$ and $C,$ respectively.
+$\Delta B=10^{-4}*N_{P}*std(B).*randn(m,n)\quad\text{and}\quad\Delta C=10^{-4}*N_{P}*std(C).*randn(p,m),$
+where $N_{P}$ is the noise percentage and $std(B)$ and $std(C)$ are the standard deviation of $B$ and $C,$ respectively.
 Here, $randn(m,n)$ is the is an $m$-by-$n$ normally distributed random matrix, generated by the MATLAB command $randn$.
@@
 We noticed that with growing $N_{P},$ the norm error of the solution remains consistently less than $10^{-8},$ which demonstrates the robustness of the proposed PESS preconditioner and the solution u is not sensitive to small perturbation on $\mathcal{A}$.
-Table 3: Numerical results of GMRES, BD, IBD, SS, RSS, EGSS, and PESS PGMRES processes for Example 2.
+Example 2.
+[…]
+$-\Delta{\bf u}+\nabla p=0,\quad\text{in}\quad\Xi,$ (6.1)
+$\quad\quad\quad\nabla\cdot{\bf u}=0,\quad\text{in}\quad\Xi.$
+[…]
+Table 3: Numerical results of GMRES, BD, IBD, SS, RSS, EGSS, PESS and LPESS PGMRES processes for Example 2.
 Process h $1/8$ $1/16$ $1/32$ $1/64$
 size($\mathcal{A}$) $288$ $1088$ $4224$ $16640$
-GMRES IT $160$ $555$ $2047$bf{--}- -
-CPU $\bf{0.0279}$ $0.2233$ $55.1948$ $2097.3235$
+GMRES IT $158$ $548$ $2048$bf{--}- -
+CPU $0.2325$ $0.6960$ $124.7595$ $4545.5072$
 RES $078.0113e-$ 07 $078.8023e-$ 07 $079.0887e-$ 07 $064.0481e-$ 06
 IBD IT $34$ $35$ $25$ $23$
-CPU $\bf{0.0066}$ $\bf 0.0697$ $\bf 1.4026$ $53.9580$
-RES 8.8956e-07 6.2172e-07 9.5786e-07 7.9687e-07
+CPU $\bf 0.0147$ $\bf 0.1166$ $1.9070$ $32.6213$
+RES $079.4111e-$ 07 $076.9932e-$ 07 $079.2544e-$ 07 $079.3340e-$ 07
 Case I
-SS IT $7$ $8$ $11$ $9$
-CPU $0.1387$ $0.2258$ $2.4943$ $70.2884$
-RES $071.4964e-$ 07 $074.2135e-$ 07 $072.2106e-$ 07 $078.7734e-$ 07
-RSS IT $\bf 4$ $5$ $5$ $6$
-CPU $0.1487$ $0.2089$ $1.4538$ $49.1211$
-RES $079.0256e-$ 07 $083.7846e-$ 08 $071.9761e-$ 07 $081.0091e-$ 08
-EGSS IT $9$ $10$ $13$ $18$
-CPU $0.1437$ $0.2381$ $2.7262$ $133.7547$
-RES $081.8639e-$ 08 $073.9243e-$ 07 $071.6563e-$ 07 $075.1266e-$ 07
-${}^{\dagger}$ IT 4 4 5 5
-CPU $0.1455$ 0.2062 $1.5472$ 44.7190
-$s=25$ RES $071.4060e-$ 07 $075.5647e-$ 07 $088.6496e-$ 08 $074.3922e-$ 07
+SS IT $5$ $6$ $9$ $11$
+CPU $0.3047$ $0.4422$ $2.4279$ $51.5825$
+RES $074.5523e-$ 07 $077.9331e-$ 07 $072.5307e-$ 07 $074.5327e-$ 07
+RSS IT $\bf 4$ $\bf 4$ $5$ $5$
+CPU $0.2044$ $0.3436$ $1.6047$ $27.6128$
+RES $083.4853e-$ 08 $079.2155e-$ 07 $081.7469e-$ 08 $072.7261e-$ 07
+EGSS IT $7$ $9$ $11$ $15$
+CPU $0.2389$ $0.4051$ $4.4353$ $52.7804$
+RES $071.6861e-$ 07 $084.0517e-$ 08 $071.3223e-$ 07 $071.774e-$ 07
+PESS IT $4$ $4$ $5$ $5$
+CPU $0.2068$ $0.3480$ $1.7633$ $25.2920$
+$s=30$ RES $082.3499e-$ 08 $074.2761e-$ 07 $082.1247e-$ 08 $072.0987e-$ 07
+LPESS IT $\bf 3$ $\bf{3}$ 4 3
+CPU $\bf 0.1990$ $\bf 0.2578$ $\bf 1.4778$ $\bf 17.9730$
+$s=30$ RES $071.6606e-$ 07 $078.3803e-$ 07 $104.9289e-$ 10 $075.2658e-$ 07
 Case II
-SS IT $41$ $56$ $77$ $66$
-CPU $0.1717$ $0.5602$ $14.0111$ $466.4250$
-RES $078.2375e-$ 07 $078.9915e-$ 07 $079.1640e-$ 07 $079.7171e-$ 07
-RSS IT $17$ $18$ $15$ $15$
-CPU $0.1564$ $0.2734$ $2.9773$ $109.9137$
-RES $071.6292e-$ 07 $071.4261e-$ 07 $075.8149e-$ 07 $076.0840e-$ 07
+SS IT $29$ $38$ $60$ $52$
+CPU $0.2999$ $1.1144$ $13.4580$ $213.2381$
+RES $077.3560e-$ 07 $079.5790e-$ 07 $079.2515e-$ 07 $079.6425e-$ 07
+RSS IT $12$ $13$ $14$ $12$
+CPU $0.3797$ $0.6019$ $3.4491$ $50.3476$
+RES $073.2456e-$ 07 $075.3086e-$ 07 $076.6600e-$ 07 $074.2834e-$ 07
 EGSS IT $9$ $9$ $9$ $7$
-CPU $0.1403$ $0.2274$ $2.0706$ $54.4502$
-RES $084.1786e-$ 08 $091.4701e-$ 09 $092.4144e-$ 09 $072.6089e-$ 07
-${}^{\dagger}$ IT $5$ 5 5 5
-CPU $0.1278$ $0.2112$ 1.5175 44.0049
-$s=26$ RES $082.0196e-$ 08 $081.8653e-$ 08 $081.5968e-$ 08 $085.7013e-$ 08
-[…]
-[…]
-Example 2.
-[…]
-[…]
+CPU $0.2459$ $0.4180$ $2.4320$ $27.9359$
+RES $097.9001e-$ 09 $098.6623e-$ 09 $097.4644e-$ 09 $071.2457e-$ 07
+PESS IT $5$ $5$ $5$ $5$
+CPU $0.2091$ $0.4184$ $1.8118$ $25.8048$
+$s=26$ RES $082.3801e-$ 08 $081.1965e-$ 08 $098.4923e-$ 09 $085.9356e-$ 08
+LPESS IT $4$ $5$ $\bf 5$ $\bf 5$
+CPU $0.2070$ $0.3662$ $\bf 1.5330$ $\bf 23.5997$
+$s=26$ RES $072.4250e-$ 07 $128.6182e-$ 12 $115.3930e-$ 11 $108.1253e-$ 10
 […]
@@
 (d) $\textbf{h}=1/64$
-Figure 5: Convergence curves for IT versus RES of the PGMRES processes by employing IBD, SS, RSS, EGSS, and PESS preconditioners in Case II for Example 2.
+Figure 5: Convergence curves for IT versus RES of the PGMRES processes by employing IBD, SS, RSS, EGSS, PESS and LPESS preconditioners in Case II for Example 2.
 […]
-  • In Case I: $\alpha=0.02$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=0.02,\beta=0.2,\gamma=0.002$ and $P=I,Q=I,W=I$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=\frac{1}{2}\alpha I,\Lambda_{2}=\frac{1}{2}\beta I,\Lambda_{3}=\frac{1}{2}\gamma I$ for ${\mathscr{P}}_{PESS}.$
-  • In Case II: $\alpha=2$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=2,\beta=2,\gamma=0.002$ and $P=A,Q=I,W=CC^{T}$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=\frac{1}{2}\alpha A,\Lambda_{2}=\frac{1}{2}\beta I,\Lambda_{3}=\frac{1}{2}\gamma CC^{T}.$
-For the IBD preconditioner, the matrices $\hat{A}$ and $\hat{S}$ are constructed as in Example 1.
-In the interval $[10,30]$, the empirical optimal choice for $s$ is found to be $25$ for Case I and $26$ for Case II.
 Table 3 shows that the GMRES process has a very slow convergence speed and also does not converge for $\textbf{h}=1/64$ within $7000$ iterations.
-The proposed PESS preconditioner requires almost five times fewer iterations compared to the IBD preconditioner for convergence.
-Moreover, in both cases, the PESS preconditioned process outperforms the SS, RSS, and EGSS preconditioners in terms of IT and CPU times.
-Notably, IT for the PESS preconditioner, in both cases, remains constant, whereas for SS, RSS, and EGSS, IT increases as the size of $\mathcal{A}$ increases.
-Furthermore, in Case II with $l=1/64$, our proposed PESS preconditioner is approximately $18\%$, $91\%$, $60\%$, and $19\%$ more efficient than the existing IBD, SS, RSS, and EGSS preconditioner, respectively.
+The proposed preconditioners require almost five times fewer iterations compared to the IBD preconditioner for convergence.
+Moreover, in both cases, the PESS and LPESS preconditioners outperform the SS, RSS, and EGSS preconditioners in terms of IT and CPU times.
+Notably, IT for the PESS preconditioner, in both cases, remains constant, whereas for SS and EGSS, IT increases as the size of $\mathcal{A}$ increases.
+Furthermore, in Case II with $l=1/64$, our proposed LPESS preconditioner is approximately $28\%$, $90\%$, $53\%$, and $16\%$ more efficient than the existing IBD, SS, RSS, and EGSS preconditioner, respectively.
 Similar patterns are observed for other values of $l.$
-The convergence curves in Figure 5 demonstrate the rapid convergence of the proposed PESS preconditioner compared to the baseline IBD, SS, RSS, EGSS, and PESS preconditioners in terms of RES versus IT counts.
 (a) $\mathcal{A}$
@@
 (f) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=26$
-Figure 6: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with ${\bf h}=1/8.$
-(a) CPU versus parameters $s$ and $\alpha$
-(b) IT versus parameters $s$ and $\alpha$
-Figure 7: Characteristic curves for CPU (left) and IT (right) of the proposed PESS iteration process by varying $s$ in the interval $[1,30]$ with step size $1$ and $\alpha$ in the interval $[0.05,1]$ with step size $0.05$ with $\textbf{h}=1/32$ in Case I for Example 2.
-Moreover, Figure 6 illustrates the spectral distributions of the original matrix $\mathcal{A},$ and the preconditioned matrices ${\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{RSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with ${\bf{h}}=1/8.$ According to Figure 6, eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ demonstrate a superior clustering compared to the other preconditioned matrices, which leads to a favorable convergence speed for the proposed PESS PGMRES process.
-Furthermore, the system (1.1) exhibits ill-conditioning nature with $\kappa(\mathcal{A})=5.07e+05.$ While for Case I and $\mathbf{h}=1/16$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=1.4024,$ indicating that the proposed PESS preconditioned system is well-conditioned, ensuring an efficient and robust solution.
-In addition, to demonstrate the relationship of the parameters and the convergence speed of the PESS preconditioner, we plot the graph of CPU times and IT counts by varying the parameters $s$ and $\alpha,$ in Figure 7.
-For $\textbf{h}=1/32$, $s$ varying in the interval $[1,30]$ with step size one and $\alpha$ from the interval $[0.05,1]$ with step size $0.05$ Other parameters remain consistent with Case I.
-We draw the following observations from Figure 7.
-  • For the given value of $s$, the proposed PESS preconditioner is insensitive to the change of the parameter $\alpha.$
-  • With an increase in $s$, IT and CPU times show a substantial decrease irrespective of the value of $\alpha.$
+(g) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=26$
+Figure 6: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf h}=1/8.$
+The convergence curves in Figure 5 demonstrate the rapid convergence of the proposed PESS and LPESS PGMRES processes compared to the baseline IBD, SS, RSS, and EGSS ones in terms of RES versus IT counts.
+Moreover, Figure 6 illustrates the spectral distributions of the original matrix $\mathcal{A},$ and the preconditioned matrices ${\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{RSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf{h}}=1/8.$ According to Figure 6, eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ demonstrate a superior clustering compared to the other preconditioned matrices, which leads to a favorable convergence speed for the proposed PESS and LPESS PGMRES processes.
+(a) PESS PGMRES process
+(b) LPESS PGMRES process
+Figure 7: Characteristic curves for IT of the proposed PESS (left) and LPESS (right) PGMRES processes by varying $s$ in the interval $[1,100]$ with step size $1$ with $\textbf{h}=1/16$ in Case I for Example 2.
+Furthermore, the system (1.1) exhibits ill-conditioning nature with $\kappa(\mathcal{A})=5.0701e+05.$ While for Case I with $\mathbf{h}=1/16$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=1.3353$ and $\kappa({\mathscr{P}}_{LPESS}^{-1}\mathcal{A})=1.2056,$ indicating that the proposed PESS and LPESS preconditioned systems are well-conditioned, ensuring an efficient and robust solution.
+In addition, to demonstrate the relationship of the parameter $s$ and the convergence speed of the PESS and LPESS preconditioners, we plot graphs of IT counts by varying the parameters $s$ in the interval $[1,100]$ with step size one in Figure 7.
+We consider $\textbf{h}=1/16$ and other choices for $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ as in Case I.
+Figure 7 shows that, with the increasing value of $s,$ decreasing trend in the $IT$ counts for both the PESS and LPESS preconditioner.
+Moreover, for $s>22,$ using LPESS preconditioner, we can solve the three-by-three $SPP$ mentioned only in $3$ iterations.
 7 Conclusion
-In this paper, we proposed the PESS iterative process and corresponding PESS preconditioner ${\mathscr{P}}_{PESS}$ to solve the three-by-three block SPP (1.1).
+In this paper, we proposed the PESS iterative process and corresponding PESS preconditioner ${\mathscr{P}}_{PESS}$ and its relaxed variant LPESS preconditioner to solve the three-by-three block SPP (1.1).
 For the convergence of the proposed PESS iterative process, necessary and sufficient criteria are derived.
-Moreover, we estimated the spectral bounds of the proposed PESS preconditioned matrices.
+Moreover, we estimated the spectral bounds of the proposed PESS and LPESS preconditioned matrices.
 This empowers us to derive spectral bounds for SS and EGSS preconditioned matrices, which have not yet been investigated in the literature to the best of our knowledge.
-Numerous experimental analyses are performed to demonstrate the effectiveness of our proposed PESS preconditioner.
-The key observations are as follows: $(i)$ the proposed PESS preconditioner is found to outperform the existing baseline preconditioners in terms of IT and CPU times.
-$(ii)$ The proposed PESS preconditioner significantly reduces the condition number of $\mathcal{A}$, consequently showing the its proficiency in solving three-by-three block SPPs.
-$(iii)$ The PESS preconditioned matrix has a more clustered spectral distribution than the baseline preconditioned matrices.
+Numerous experimental analyses are performed to demonstrate the effectiveness of our proposed PESS and LPESS preconditioners.
+The key observations are as follows: $(i)$ the proposed PESS and LPESS preconditioners are found to outperform the existing baseline preconditioners in terms of IT and CPU times.
+$(ii)$ The proposed preconditioners significantly reduces the condition number of $\mathcal{A}$, consequently showing the its proficiency in solving three-by-three block SPPs.
+$(iii)$ The PESS and LPESS preconditioned matrices have a more clustered spectral distribution than the baseline preconditioned matrices.
 $(iv)$ Sensitivity analysis conducted by introducing different percentages of noise on the system (1.1) showcases the robustness of the proposed PESS preconditioner.
-Acknowledgements
-The second author (Pinki Khatun) sincerely acknowledges the Council of Scientific &\&& Industrial Research (CSIR) in New Delhi, India, for financial assistance in the form of a fellowship (File no.
-09/1022(0098)/2020-EMR-I).
+Further investigation to determine the optimal parameter $s$ can be done in future research endeavors.
 Declaration of competing interest
 The authors disclose that they have no conflicting interests.
-Data availability
-No data was used for the research described in the article.
+Acknowledgments
+Pinki Khatun sincerely acknowledges the Council of Scientific &\&& Industrial Research (CSIR) in New Delhi, India, for financial assistance in the form of a fellowship (File no.
+09/1022(0098)/2020-EMR-I).

### v3

@@
 Additionally, we introduce a local PESS (LPESS) preconditioner by relaxing the PESS preconditioner.
-Necessary and sufficient criteria are established for the convergence of the proposed PESS iterative process for any random initial guess.
+Necessary and sufficient criteria are established for the convergence of the proposed PESS iterative process for any initial guess.
 Furthermore, we meticulously investigate the spectral bounds of the PESS and LPESS preconditioned matrices.
 Moreover, empirical investigations have been performed for the sensitivity analysis of the proposed PESS preconditioner, which unveils its robustness.
-Numerical experiments are carried out to demonstrate the enhanced efficiency and robustness of the proposed PESS and LPESS preconditioners compared to the existing block diagonal and shift-splitting preconditioners.
-Keywords. Krylov subspace iterative process $\cdot$ Preconditioner $\cdot$ Shift-splitting $\cdot$ Saddle point problem $\cdot$ Robust $\cdot$ GMRES
-AMS subject classification. 65F08 $\cdot$ 65F10 $\cdot$ 65F50
+Numerical experiments are carried out to demonstrate the enhanced efficiency and robustness of the proposed PESS and LPESS preconditioners compared to the existing state-of-the-art preconditioners.
+Keywords. Krylov subspace iterative process, Preconditioner, Shift-splitting, Saddle point problem, Robust, GMRES
+AMS subject classification. 65F08, 65F10, 65F50
 1 Introduction
@@
 $\mathcal{A}{\bf u}={\bf d},$ (1.1)
-where $\mathcal{A}=\left[\begin{matrix}A&B^{T}&0\\ -B&0&-C^{T}\\ 0&C&0\end{matrix}\right],\,{\bf u}=\left[\begin{matrix}x\\ y\\ z\end{matrix}\right],\,{\bf d}=\left[\begin{matrix}f\\ g\\ h\end{matrix}\right],$ and $A\in{\mathbb{R}}^{n\times n}$ is a symmetric positive definite (SPD) matrix, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are the full row rank matrices.
-Here, $B^{T}$ stands for the transpose of $B,$ and $f\in{\mathbb{R}}^{n},\,g\in{\mathbb{R}}^{m}$ and $h\in{\mathbb{R}}^{p}$ are known vectors.
-Under the stated assumptions on block matrices A, B, and C, $\mathcal{A}$ is nonsingular, which ensures the existence of a unique solution for the system (1.1).
+where $\mathcal{A}=\left[\begin{matrix}A&B^{T}&0\\ -B&0&-C^{T}\\ 0&C&0\end{matrix}\right],\,{\bf u}=\left[\begin{matrix}x\\ y\\ z\end{matrix}\right],\,{\bf d}=\left[\begin{matrix}f\\ g\\ h\end{matrix}\right],$ $A\in{\mathbb{R}}^{n\times n}$ is a symmetric positive definite (SPD) matrix, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are the full row rank matrices.
+Here, $B^{T}$ and $C^{T}$ stand for the transpose of $B$ and $C$, respectively; $f\in{\mathbb{R}}^{n},\,g\in{\mathbb{R}}^{m}$ and $h\in{\mathbb{R}}^{p}$ are known vectors.
+Under the stated assumptions on block matrices A, B and C, $\mathcal{A}$ is nonsingular, which ensures the existence of a unique solution for the system (1.1).
 Due to the growing popularity of three-by-three block SPPs in a variety of applications, it is essential to determine the efficient and robust numerical solution of three-by-three block SPPs (1.1).
@@
 Moreover, the saddle point matrix $\mathcal{A}$ can be very sensitive, which leads to a significant slowdown in the solution algorithm.
-Therefore, it is important to develop novel, efficient, and robust preconditioners for the fast convergence of the Krylov subspace iterative process, which can handle the sensitivity of the problem (1.1) when the matrix $\mathcal{A}$ is ill-conditioned.
+Therefore, it is important to develop novel, efficient and robust preconditioners for the fast convergence of the Krylov subspace iterative process, which can handle the sensitivity of the problem (1.1) when the matrix $\mathcal{A}$ is ill-conditioned.
 Consider the following partitioning of $\mathcal{A}:$
@@
 Then the linear system (1.1) can be identified as the standard two-by-two block SPP [10].
-In the last decades, colossal attention has been given to the development of numerical solution methods for standard two-by-two block SPP, such as Uzawa methods [7, 13, 2], Krylov subspace iterative process [17, 26], null space methods [15, 25].
+In the last decades, colossal attention has been given to the development of numerical solution methods for standard two-by-two block SPP, such as Uzawa methods [11, 12], Hermitian and skew-Hermitian splitting-type (HSS-type) methods [13, 14], successive overrelaxation-type (SOR-type) methods [15, 16] and null space methods [17, 18].
+For a more recent comprehensive survey see [19].
 However, the properties of the submatrices in (1.8) are different from the standard two-by-two block SPP.
-Notice that the leading blocks are skew-symmetric and positive semidefinite in the first partitioning.
-The second partitioning highlights that the $(1,2)$ block is rank deficient, and the $(2,2)$ block is skew-symmetric.
-In contrast, in the standard SPP, $(1,2)$ block is full row rank, and the $(2,2)$ block is generally symmetric positive semidefinite.
-Therefore, the existing literature for solving the two-by-two SPPs (see [5]) may not be applied to solve (1.1) directly, and it is essential to develop new preconditioners for three-by-three block SPP that exploit the specific structure of the saddle point matrix $\mathcal{A}$, which is sensitive in nature.
+Notice that the leading block is not SPD in the first partitioning.
+The second partitioning highlights that the $(1,2)$ block is rank deficient and the $(2,2)$ block is skew-symmetric.
+In contrast, in the standard two-by-two block SPP, $(1,2)$ block is full row rank, and the $(2,2)$ block is generally symmetric positive semidefinite.
+Therefore, the existing literature for solving the two-by-two block SPPs (see [10]) may not be applied to solve (1.1) directly and it is essential to develop new preconditioners for three-by-three block SPPs that exploit the specific structure of the saddle point matrix $\mathcal{A}$, which is sensitive in nature.
 Recently, various effective preconditioners have been developed in the literature for solving three-by-three block SPPs.
-For instance, Huang and Ma [18] has studied the exact block diagonal (BD) and inexact BD (IBD) preconditioners ${\mathscr{P}}_{BD}$ and ${\mathscr{P}}_{IBD}$ in the following forms:
+For instance, Huang and Ma [8] have studied the exact block diagonal (BD) and inexact BD (IBD) preconditioners ${\mathscr{P}}_{BD}$ and ${\mathscr{P}}_{IBD}$ in the following forms:
 ${\mathscr{P}}_{BD}=\left[\begin{matrix}A&0&0\\ 0&S&0\\ 0&0&CS^{-1}C^{T}\end{matrix}\right]\quad\mbox{and}\quad{\mathscr{P}}_{IBD}=\left[\begin{matrix}\widehat{A}&0&0\\ 0&\widehat{S}&0\\ 0&0&C\widehat{S}^{-1}C^{T}\end{matrix}\right],$ (1.9)
 where $S=BA^{-1}B^{T},$ $\widehat{A}$ and $\widehat{S}$ are SPD approximations of $A$ and $S,$ respectively.
-Although the spectrum of the preconditioned matrices ${\mathscr{P}}^{-1}_{BD}\mathcal{A}$ and ${\mathscr{P}}^{-1}_{IBD}\mathcal{A}$ have good clustering properties, these preconditioners have certain shortcomings, such as they are time-consuming, expensive, require excessive number of iterations, and condition numbers are very large.
+Although the spectrum of the preconditioned matrices ${\mathscr{P}}^{-1}_{BD}\mathcal{A}$ and ${\mathscr{P}}^{-1}_{IBD}\mathcal{A}$ have good clustering properties, these preconditioners have certain shortcomings, such as they are time-consuming, expensive, require excessive number of iterations and condition numbers are very large.
 For more details, see the paper [8].
-To overcome these drawbacks of BD preconditioners, researchers have started exploring shift-splitting (SS) preconditioners, which was initially developed for a non-Hermitian system of linear equations by Bai et al.
+Inspired by the HSS iteration method [13], Salkuyeh et al.
+[20] split the coefficient matrix $\mathcal{A}$ and presented the alternating positive semidefinite splitting (APSS) iteration method.
+They proved the iteration method is unconditional convergence and proposed the corresponding APSS preconditioner to solve the system (1.1).
+Moreover, to improve the effectiveness of the APSS preconditioner, many relaxed and modified versions of APSS have been designed; see [21, 22, 23].
+For instance, by relaxing the $(1,1)$ block and introducing a new parameter $\beta>0$ in the APSS preconditioner, Chen and Ren [21] proposed the following modified APSS (MAPSS) preconditioner:
+${\mathscr{P}}_{MAPSS}=\left[\begin{matrix}A&B^{T}&-\frac{1}{\alpha}B^{T}C^{T}\\ -B&\alpha I&-C^{T}\\ 0&C&\beta I\end{matrix}\right],$ (1.10)
+where $\alpha>0$ and $I$ stands for the identity matrix of the appropriate dimension.
+Motivated by the work in [24] and by incorporating the ideas of shifting and the BD preconditioner, the authors in [25] proposed two preconditioners along with their inexact variants.
+One of these preconditioners is given by:
+${\mathscr{P}}_{SL}=\left[\begin{matrix}A&B^{T}&0\\ -B&C^{T}C&0\\ 0&C&I\end{matrix}\right].$ (1.11)
+Three block preconditioners are developed by Xie and Li [26] for the system (1.1) with the equivalent symmetric coefficient matrix.
+It is also demonstrated in [26] that the preconditioned matrices possess no more than three distinct eigenvalues.
+Huang [27] proposed a variant of the Uzawa iterative method for the theree-by-three block SPP by introducing two variable parameters.
+Additionally, Huang et al.
+[28] generalized the well-known Uzawa method to solve the linear system (1.1) and propose the inexact Uzawa method.
+In addition to the preconditioners discussed above, recent literature [29, 30, 31] have introduced several other preconditioners to solve the three-by-three block SPP (1.1).
+The shift-splitting (SS) preconditioners were initially developed for a non-Hermitian system of linear equations by Bai et al.
 [32] and later for two-by-two block SPPs [33, 34, 35].
-Recently, Cao [9] enhanced this idea and introduced the following SS preconditioner ${\mathscr{P}}_{SS}$ and relaxed SS (RSS) preconditioner ${\mathscr{P}}_{RSS}$ for the three-by-three block SPP (1.1)
+Recently, Cao [36] enhanced this idea and introduced the following SS preconditioner ${\mathscr{P}}_{SS}$ and relaxed SS (RSS) preconditioner ${\mathscr{P}}_{RSS}$ for the three-by-three block SPP (1.1):
 ${\mathscr{P}}_{SS}=\frac{1}{2}\left[\begin{matrix}\alpha I+A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right]\,\,\mbox{and}\,\,{\mathscr{P}}_{RSS}=\frac{1}{2}\left[\begin{matrix}A&B^{T}&0\\ -B&\alpha I&-C^{T}\\ 0&C&\alpha I\end{matrix}\right],$ (1.12)
 where $\alpha>0$ and verified unconditionally convergence of the associated SS iterative process.
-Here, $I$ stands for the identity matrix of the appropriate dimension.
 Wang and Zhang [37] generalized the SS preconditioner by introducing a new parameter $\beta>0$ in the $(3,3)$ block.
+By merging the lopsided and shift technique, a lopsided SS preconditioner is presented by Zhang et al.
+[38].
 Further, Yin et al.
-[31] developed an extensive generalized SS (EGSS) preconditioner
-${\mathscr{P}}_{EGSS}=\frac{1}{2}\left[\begin{matrix}\alpha P+A&B^{T}&0\\ -B&\beta Q&-C^{T}\\ 0&C&\gamma W\end{matrix}\right],$ (1.11)
+[39] developed the following extensive generalized SS (EGSS) preconditioner:
+${\mathscr{P}}_{EGSS}=\frac{1}{2}\left[\begin{matrix}\alpha P+A&B^{T}&0\\ -B&\beta Q&-C^{T}\\ 0&C&\gamma W\end{matrix}\right]$ (1.13)
 for the three-by-three block SPP (1.1), where $\alpha,\beta,\gamma>0$ and $P,Q,W$ are SPD matrices and investigated its convergent properties.
-There are a number of other preconditioners and iterative processes to solve the three-by-three block SPP (1.1) has been studied recently in the literature [19, 1, 24, 29, 4].
+By relaxing the $(1,1)$ block and eliminating the prefactor $1/2,$ Liang and Zhu [40] have proposed relaxed and preconditioned generalized shift-splitting (RPGSS) preconditioner
+${\mathscr{P}}_{RPGSS}=\left[\begin{matrix}A&B^{T}&0\\ -B&\beta Q&-C^{T}\\ 0&C&\gamma W\end{matrix}\right],$ (1.14)
+where $Q\in{\mathbb{R}}^{m\times m}$ and $W\in{\mathbb{R}}^{p\times p}$ are SPD and $\beta,\gamma>0$.
 Despite exhibiting favorable performance of SS, RSS and EGSS preconditioners, they do not outperform BD and IBD preconditioners.
 Therefore, to improve the convergence speed and efficiency of the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS}$ and ${\mathscr{P}}_{EGSS},$ this paper presents a parameterized enhanced shift-splitting (PESS) preconditioner by introducing a parameter $s>0$ for there-by-three block SPPs.
-As per the direct correlation between the rate of convergence in Krylov subspace iterative processes and the spectral properties of the preconditioned matrix, we perform in-depth spectral distribution of PESS preconditioned matrix.
-We summarize the following main contributions of this paper:
+As per the direct correlation between the rate of convergence in Krylov subspace iterative processes and the spectral properties of the preconditioned matrix, we perform in-depth spectral distribution of the PESS preconditioned matrix.
+We summarize the main contributions of this paper as follows:
   • We propose the PESS iterative process along with its associated PESS preconditioner and its relaxed version known as the local PESS (LPESS) preconditioner to solve the three-by-three block SPP (1.1).
@@
 These investigations also encompass the unconditional convergence of other exiting SS preconditioners in the literature [36, 39].
-  • We have conducted a comprehensive analysis of the spectral distribution for the PESS preconditioned matrix.
-Our framework allows us to derive the spectral distribution for the SS and EGSS preconditioned matrices.
+  • We have conducted a comprehensive analysis of the spectral distribution for the PESS and LPESS preconditioned matrices.
+Our framework allows us to derive the spectral distribution for the existing SS and EGSS preconditioned matrices.
   • We empirically show that our proposed preconditioner significantly reduces the condition number of the ill-conditioned saddle point matrix $\mathcal{A}$.
 Thereby establishing an efficiently solvable, well-conditioned system.
-  • Numerical experiments show that the proposed PESS and LPESS preconditioners outperform the existing BD and SS preconditioners.
+  • Numerical experiments show that the proposed PESS and LPESS preconditioners outperform all the compared baseline preconditioners.
 Moreover, we perform sensitivity analysis by perturbing the ill-conditioned matrix $\mathcal{A}$ in (1.1), demonstrating the robustness of the proposed preconditioner.
@@
 In Section 5, we present the LPESS preconditioner.
+In Section 6, we discuss strategies for selecting the appropriate parameters for the proposed preconditioners.
 In Section 7, we conduct numerical experiments to illustrate the computational efficiency and robustness of the proposed preconditioner.
@@
 The following notations are adopted throughout the paper.
-For any matrix $A\in{\mathbb{R}}^{n\times n},$ $A^{T},\sigma(A)$ and $\vartheta(A)$ denotes transpose, spectrum and spectral radius of $A,$ respectively.
+For any matrix $A\in{\mathbb{R}}^{n\times n},$ $A^{T},\sigma(A)$ and $\vartheta(A)$ denote transpose, spectrum and spectral radius of $A,$ respectively.
 For any $x\in\mathbb{C}^{n},$ $x^{*}$ denotes its conjugate transpose.
 The symbol $\mathrm{diag}(A,B)$ represents a block diagonal matrix with diagonal blocks $A$ and $B.$ The real and imaginary parts of any complex number $z$ are represented by $\Re(z)$ and $\Im(z)$, respectively.
-For any symmetric matrix $A\in{\mathbb{R}}^{n\times n},$ $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$ denotes the maximum and minimum eigenvalue of $A,$ respectively.
-The notation $A>0$ signifies that $A$ is SPD, and $A>B$ identifies $A-B>0.$ Abbreviations used in the paper are as follows:
-SPP: saddle point problem, SPD: symmetric positive definite, BD: block diagonal, IBD: inexact block diagonal, SS: shift-splitting, RSS: relaxed shift-splitting, EGSS: extensive generalized shift-spliting, PESS: parameterized enhanced shift-splitting, LPESS: local parameterized enhanced shift-splitting GMRES: generalized minimum residual process.
+For any matrix $A\in{\mathbb{R}}^{n\times n}$ having real eigenvalues, $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$ denote the maximum and minimum eigenvalue of $A,$ respectively.
+The notation $A>0$ signifies that $A$ is SPD and $A>B$ identifies $A-B>0.$
+Abbreviations used in the paper are as follows:
+SPP: saddle point problem, SPD: symmetric positive definite, BD: block diagonal, IBD: inexact block diagonal, MAPSS: modified alternating positive semidefinite splitting, SS: shift-splitting, RSS: relaxed shift-splitting, EGSS: extensive generalized shift-splitting, RPGSS: relaxed and preconditioned generalized shift-splitting, PESS: parameterized enhanced shift-splitting, LPESS: local parameterized enhanced shift-splitting, GMRES: generalized minimum residual process.
 2 The parameterized enhanced shift-splitting (PESS) iterative process and the preconditioner
@@
 ${\mathscr{P}}_{PESS}=\left[\begin{matrix}\Lambda_{1}+sA&sB^{T}&0\\ -sB&\Lambda_{2}&-sC^{T}\\ 0&sC&\Lambda_{3}\end{matrix}\right],\mathcal{Q}_{PESS}=\left[\begin{matrix}\Lambda_{1}-(1-s)A&-(1-s)B^{T}&0\\ (1-s)B&\Lambda_{2}&(1-s)C^{T}\\ 0&-(1-s)C&\Lambda_{3}\end{matrix}\right],$
-$\Sigma=\mathrm{diag}(\Lambda_{1},\Lambda_{2},\Lambda_{3}),$ and $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ are SPD matrices.
-The matrix ${\mathscr{P}}_{PESS}$ is nonsingular for $s>0.$ The special matrix splitting (2.1) leads us to the subsequent iterative process for solving the three-by-three block SPP given in (1.1):
+$\Sigma=\mathrm{diag}(\Lambda_{1},\Lambda_{2},\Lambda_{3})$ and $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ are SPD matrices.
+The matrix ${\mathscr{P}}_{PESS}$ is nonsingular for $s>0.$ The special matrix splitting (2.1) leads us to the subsequent iterative process for solving the three-by-three block SPP given in (1.1).
 The PESS iterative process.
-Let $s$ be a positive constant, and let $\Lambda_{1}\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}$ $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices.
+Let $s$ be a positive constant and let $\Lambda_{1}\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}{,}$ $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices.
 For any initial guess ${\bf u}_{0}\in{\mathbb{R}}^{n+m+p}$ and $k=0,1,2\ldots,$ until a specified termination criterion is fulfilled, we compute
 ${\bf u}_{k+1}=\mathcal{T}{\bf u}_{k}+{\bf c},$ (2.2)
-where ${\bf u}_{k}=[x_{k}^{T},y_{k}^{T},z_{k}^{T}]^{T},{\bf c}={\mathscr{P}}_{PESS}^{-1}{\bf d},$ and $\mathcal{T}={\mathscr{P}}_{PESS}^{-1}\mathcal{Q}_{PESS}$ is called the iteration matrix for the PESS iterative process.
+where ${\bf u}_{k}=[x_{k}^{T},y_{k}^{T},z_{k}^{T}]^{T},{\bf c}={\mathscr{P}}_{PESS}^{-1}{\bf d}$ and $\mathcal{T}={\mathscr{P}}_{PESS}^{-1}\mathcal{Q}_{PESS}$ is called the iteration matrix for the PESS iterative process.
 Moreover, the matrix splitting (2.2) introduces a new preconditioner ${\mathscr{P}}_{PESS}$, identified as PESS preconditioner.
-This preconditioner is a generalization of the previous work in the literature, see for example [9, 27, 31] for specific choices of $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ and $s,$ which are listed in Table 1.
+This preconditioner generalizes the previous work in the literature; for example, refer to [36, 37, 39] for specific choices of $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ and $s,$ which are listed in Table 1.
 Table 1: ${\mathscr{P}}_{PESS}$ as a generalization of the above SS preconditioners for different choices of $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ and $s.$
@@
 matrices and $\alpha,\beta,\gamma>0$
-In the implementation of the PESS iterative process or the PESS preconditioner to enhance the rate of convergence of the Krylov subspace iterative process like GMRES, at each iterative step, we solve the following system of linear equations
+In the implementation of the PESS iterative process or the PESS preconditioner to enhance the rate of convergence of the Krylov subspace iterative process like GMRES, at each iterative step, we solve the following system of linear equations:
 ${\mathscr{P}}_{PESS}w=r,$ (2.3)
-where $r=[r_{1}^{T},r_{2}^{T},r_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}$ and $w=[w_{1}^{T},w_{2}^{T},w_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}.$ Specifying $\widehat{X}=\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C$ and $\widehat{A}=\Lambda_{1}+sA+s^{2}B^{T}\widehat{X}^{-1}B,$ then ${\mathscr{P}}_{PESS}$ can be written in the following way
-${\mathscr{P}}_{PESS}=\left[\begin{matrix}I&sB^{T}\widehat{X}^{-1}&0\\ 0&I&-sC^{T}\Lambda_{3}^{-1}\\ 0&0&I\end{matrix}\right]\left[\begin{matrix}\widehat{A}&0&0\\ 0&\widehat{X}&0\\ 0&0&\Lambda_{3}\end{matrix}\right]\left[\begin{matrix}I&0&0\\ -s\widehat{X}^{-1}B&I&0\\ 0&s\Lambda_{3}^{-1}C&I\end{matrix}\right].$ (2.4)
+where $r=[r_{1}^{T},r_{2}^{T},r_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}$ and $w=[w_{1}^{T},w_{2}^{T},w_{3}^{T}]^{T}\in{\mathbb{R}}^{n+m+p}.$ Specifying $\widehat{X}=\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C$ and $\widetilde{A}=\Lambda_{1}+sA+s^{2}B^{T}\widehat{X}^{-1}B,$ then ${\mathscr{P}}_{PESS}$ can be written in the following way:
+${\mathscr{P}}_{PESS}=\left[\begin{matrix}I&sB^{T}\widehat{X}^{-1}&0\\ 0&I&-sC^{T}\Lambda_{3}^{-1}\\ 0&0&I\end{matrix}\right]\left[\begin{matrix}\widetilde{A}&0&0\\ 0&\widehat{X}&0\\ 0&0&\Lambda_{3}\end{matrix}\right]\left[\begin{matrix}I&0&0\\ -s\widehat{X}^{-1}B&I&0\\ 0&s\Lambda_{3}^{-1}C&I\end{matrix}\right].$ (2.4)
 The decomposition (2.4) leads us to the following Algorithm 1 to determine the solution of the system (2.3).
@@
 2:Compute $v=r_{1}-sB^{T}v_{1};$
-3:Solve $\widehat{A}w_{1}=v$ to find $w_{1}$
+3:Solve $\widetilde{A}w_{1}=v$ to find $w_{1};$
 4:Solve $\widehat{X}v_{2}=sBw_{1}$ to find $v_{2}$;
@@
 6:Solve $\Lambda_{3}w_{3}=r_{3}-sCw_{2}$ for $w_{3}.$
-The implementation of Algorithm 1 requires to solve two linear subsystems with coefficient matrices $\widehat{X}$ and $\widehat{A}$.
-Since $\Lambda_{1},\Lambda_{2},\Lambda_{3},\widehat{X},$ and $\widehat{A}$ are SPD and $s>0$, we can apply the Cholesky factorization to solve these linear subsystems exactly or inexactly by the preconditioned conjugate gradient method.
+The implementation of Algorithm 1 requires to solve two linear subsystems with coefficient matrices $\widehat{X}$ and $\widetilde{A}$.
+Since $\Lambda_{1},\Lambda_{2},\Lambda_{3},\widehat{X}$ and $\widetilde{A}$ are SPD and $s>0$, we can apply the Cholesky factorization to solve these linear subsystems exactly or inexactly by the preconditioned conjugate gradient method.
 3 Convergence analysis of the PESS iterative process
@@
 Lemma 3.1.
-[9] Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices.
+[36] Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices.
 Then, the saddle point matrix $\mathcal{A}$ is positive stable.
 Lemma 3.2.
-Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices.
-Then $\Sigma^{-1}\mathcal{A}$ ((((or, $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})$ is positive stable.
+Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices.
+Then $\Sigma^{-1}\mathcal{A}$ ((((or $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})$ is positive stable.
 Proof.
@@
 $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}=\left[\begin{matrix}\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}&\Lambda_{1}^{-\frac{1}{2}}B^{T}\Lambda_{2}^{-\frac{1}{2}}&0\\ -\Lambda_{2}^{-\frac{1}{2}}B\Lambda_{1}^{-\frac{1}{2}}&0&\Lambda_{2}^{-\frac{1}{2}}C^{T}\Lambda_{3}^{-\frac{1}{2}}\\ 0&\Lambda_{3}^{-\frac{1}{2}}C\Lambda_{2}^{-\frac{1}{2}}&0\end{matrix}\right].$ (3.1)
-Since the matrices $\Lambda_{1},\Lambda_{2},$ and $\Lambda_{3}$ are SPD, then $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$ is SPD, and further $\Lambda_{2}^{-\frac{1}{2}}B\Lambda_{1}^{-\frac{1}{2}}$ and $\Lambda_{3}^{-\frac{1}{2}}C\Lambda_{2}^{-\frac{1}{2}}$ are of full row rank.
+Given that the matrices $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ are SPD, then $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$ is SPD.
+Furthermore, the matrices $\Lambda_{2}^{-\frac{1}{2}}B\Lambda_{1}^{-\frac{1}{2}}$ and $\Lambda_{3}^{-\frac{1}{2}}C\Lambda_{2}^{-\frac{1}{2}}$ are of full row rank.
 Consequently, the block structure of $\mathcal{A}$ and the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ are identical.
-Then by Lemma 3.1, $\mathcal{A}$ is positive stable implies that, the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ and consequently, $\Sigma^{-1}\mathcal{A}$ is positive stable.
+Then by Lemma 3.1, $\mathcal{A}$ is positive stable implies that the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ and consequently, $\Sigma^{-1}\mathcal{A}$ is positive stable.
 $\blacksquare$
 By [41], the stationary iterative process (2.2) converges to the exact solution of the three-by-three block SPP (1.1) for any initial guess vector if and only if $|\vartheta(\mathcal{T})|<1,$ where $\mathcal{T}$ is the iteration matrix.
-Now, we establish an if and only if a condition that precisely determines the convergence of the PESS iterative process.
+Now, we establish an if and only if condition that precisely determines the convergence of the PESS iterative process.
 Theorem 3.1.
-Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0$.
+Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices and $s>0$.
 Then, the PESS iterative process (2.2) converges to the unique solution of the three-by-three block SPP (1.1) if and only if
@@
 $\theta=\frac{1-(1-s)\mu}{1+s\mu}.$
-Since $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ is nonsingular then $\mu\neq 0,$ and we have
+Since $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ is nonsingular then $\mu\neq 0$ and we have
 $|\theta|=\frac{|1-(1-s)\mu|}{|1+s\mu|}=\sqrt{\frac{(1-(1-s)\Re(\mu))^{2}+(1-s)^{2}\Im(\mu)^{2}}{(1+s\Re(\mu))^{2}+s^{2}\Im(\mu)^{2}}}.$ (3.5)
@@
 This implies $|\theta|<1$ if and only if $(2s-1)|\mu|^{2}+2\Re(\mu)>0.$ Since, $\vartheta(\mathcal{T})=\vartheta{\widetilde{(}\mathcal{T}}),$ the proof is conclusive. $\blacksquare$
-Using the condition in Theorem 3.1, next, we present two sufficient conditions that ensure the convergence of the PESS iterative process. The first one is presented as follows:
+Using the condition in Theorem 3.1, next, we present two sufficient conditions that ensure the convergence of the PESS iterative process. The first one is presented as follows.
 Corollary 3.1.
-Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and if $s\geq\frac{1}{2},$ then $\vartheta(\mathcal{T})<1,$ i.e., the PESS iterative process always converges to the unique solution of the three-by-three block SPP (1.1).
+Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices and if $s\geq\frac{1}{2},$ then $\vartheta(\mathcal{T})<1,$ i.e., the PESS iterative process always converges to the unique solution of the three-by-three block SPP (1.1).
 Proof.
@@
 Remark 3.1.
-By applying Corollary 3.1, the unconditional convergence of the existing preconditioners, namely ${\mathscr{P}}_{SS}$, ${\mathscr{P}}_{GSS},$ and ${\mathscr{P}}_{EGSS}$ can be obtained from ${\mathscr{P}}_{PESS}$ by sustituting $s=\frac{1}{2}$.
+By applying Corollary 3.1, the unconditional convergence of the existing preconditioners, namely ${\mathscr{P}}_{SS}$, ${\mathscr{P}}_{GSS}$ and ${\mathscr{P}}_{EGSS}$ can be obtained from ${\mathscr{P}}_{PESS}$ by sustituting $s=1/2$.
 Next, we present a stronger sufficient condition to Corollary 3.1 for the convergence of the PESS iterative process.
 Corollary 3.2.
-Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, if
-$s>\max\left\{\frac{1}{2}\left(1-\frac{\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}\right),0\right\},$
-then the PESS iterative process is convergent for any random initial guess.
+Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, if
+$s>\max\left\{\frac{1}{2}\left(1-\frac{\lambda_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}\right),0\right\},$
+then the PESS iterative process is convergent for any initial guess.
 Proof.
 Note that, as $\Re(\mu)>0,$ the condition in (3.2) holds if and only if $\frac{1}{2}-\frac{\Re(\mu)}{|\mu|^{2}}<s.$ Let ${\bf p}$ be an eigenvector corresponding to $\mu,$ then we have
-$\Re(\mu)=\frac{{\bf p}^{*}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}){\bf p}}{2{\bf p}^{*}{\bf p}}\geq\frac{1}{2}\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}).$
-Again, we have $|\mu|\leq\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}),$ then we get the following
-$\frac{1}{2}-\frac{\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{2\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}\geq\frac{1}{2}-\frac{\Re(\mu)}{|\mu|^{2}}.$ (3.6)
+$\Re(\mu)=\frac{{\bf p}^{*}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}){\bf p}}{2{\bf p}^{*}{\bf p}}\geq\frac{1}{2}\lambda_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}).$
+Since $|\mu|\leq\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}),$ we obtain the following:
+$\frac{1}{2}-\frac{\lambda_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{2\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}\geq\frac{1}{2}-\frac{\Re(\mu)}{|\mu|^{2}}.$ (3.6)
 If we apply $s>0$, then we get the desired result from (3.6). $\blacksquare$
-The sufficient condition on Corollary 3.2 is difficult to find for large $\mathcal{A}$ due to the involvement of computation of $\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})$ and $\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}.$ Thus, we consider $s\geq\frac{1}{2}$ for practical implementation.
+The sufficient condition on Corollary 3.2 is difficult to find for large $\mathcal{A}$ due to the involvement of computation of $\lambda_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})$ and $\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}.$ Thus, we consider $s\geq\frac{1}{2}$ for practical implementation.
 4 Spectral distribution of the PESS preconditioned matrix
@@
 Theorem 4.1.
-Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are full row rank matrices.
+Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are full row rank matrices.
 Suppose that $\lambda$ is an eigenvalue of the PESS preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$.
@@
 Proof.
-Assume that $\lambda$ and $\theta$ are the eigenvalue of the PESS preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and the iteration matrix $\mathcal{T}$, respectively.
+Assume that $\lambda$ and $\theta$ are the eigenvalues of the PESS preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and the iteration matrix $\mathcal{T}$, respectively.
 On the other hand, from the matrix splitting (2.1), we have
 ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}={\mathscr{P}}_{PESS}^{-1}({\mathscr{P}}_{PESS}-\mathcal{Q}_{PESS})=I-\mathcal{T}.$
-Then, we get $\lambda=1-\theta.$ Now, from Corollary 3.1, for $s\geq\frac{1}{2},$ it holds that $|\theta|<1.$ Hence, it follows that, for $s\geq\frac{1}{2},$ $|\lambda-1|<1,$ which completes the proof.
+Then, we get $\lambda=1-\theta.$ Now, from Corollary 3.1, it holds that $|\theta|<1$ for $s\geq\frac{1}{2}.$ Hence, it follows that, $|\lambda-1|<1$ for $s\geq\frac{1}{2},$ which completes the proof.
 $\blacksquare$
@@
 Remark 4.1.
-Notice that from (4.2a)-(4.2c), we get $\lambda\neq\frac{1}{s},$ otherwise , ${\bf p}=[u^{T},v^{T},w^{T}]^{T}={0}\in{\mathbb{R}}^{n+m+p},$ which is impossible as ${\bf p}$ is an eigenvector.
+Notice that from (4.2a)-(4.2c), we get $\lambda\neq 1/s,$ otherwise , ${\bf p}=[u^{T},v^{T},w^{T}]^{T}={0}\in{\mathbb{R}}^{n+m+p},$ which is impossible as ${\bf p}$ is an eigenvector.
 Proposition 4.1.
-Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0.$ Let $(\lambda,{\bf p}=[u^{T},v^{T},w^{T}]^{T})$ be an eigenpair of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then, the following holds
+Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices and $s>0.$ Let $(\lambda,{\bf p}=[u^{T},v^{T},w^{T}]^{T})$ be an eigenpair of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then, the following holds:
   1. $(1)$
@@
 $(1)$ Let $u=0$.
-Then by (4.2a), we get $(1-s\lambda)B^{T}v=0,$ which shows that $v=0,$ as $B$ has full row rank and $\lambda\neq\frac{1}{s}.$ Furthermore, when combined with (4.2c) leads to $w=0,$ as $\Lambda_{3}$ is SPD.
+Then by (4.2a), we get $(1-s\lambda)B^{T}v=0,$ which shows that $v=0,$ as $B$ has full row rank and $\lambda\neq 1/s.$ Furthermore, when combined with (4.2c) leads to $w=0,$ as $\Lambda_{3}$ is SPD.
 Combining the above facts, it follows that ${\bf p}=[u^{T},v^{T},w^{T}]^{T}=0,$ which is impossible as ${\bf p}$ is an eigenvector.
 Hence, $u\neq 0.$
-$(2)$ Let ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ be an eigenvector corresponding to the real eigenvalue $\lambda.$ Then, we have ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}{\bf p}=\lambda{\bf p},$ or, $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p}.$ Consequently, $\lambda$ can be written as
+$(2)$ Let ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ be an eigenvector corresponding to the real eigenvalue $\lambda.$ Then, we have ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}{\bf p}=\lambda{\bf p},$ or $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}\,{\bf p}.$ Consequently, $\lambda$ can be written as
 $\lambda$ $=$ $\frac{{\bf p}^{T}\mathcal{A}{\bf p}+{\bf p}^{T}\mathcal{A}^{T}\bf p}{2({\bf p}^{T}{\mathscr{P}}_{PESS}{\bf p}+{\bf p}^{T}{\mathscr{P}}_{PESS}^{T}\bf p)}$
 $=$ $\frac{u^{T}Au}{2(su^{T}Au+u^{T}\Lambda_{1}u+v^{T}\Lambda_{2}v+w^{T}\Lambda_{3}w)}>0.$
-The last inequality follows from the assumptions that $A,\Lambda_{1},\Lambda_{2}$, and $\Lambda_{3}$ are SPD matrices.
+The last inequality follows from the assumptions that $A,\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ are SPD matrices.
 $(3)$ The system of linear equations (4.2a)-(4.2c) can be reformulated as
 $\left\{\begin{array}[]{lcl}Au+B^{T}v=(\lambda/(1-s\lambda))\Lambda_{1}u,&\\ -Bu-C^{T}w=(\lambda/(1-s\lambda))\Lambda_{2}v,&\\ Cv=(\lambda/(1-s\lambda))\Lambda_{3}w.\end{array}\right.$ (4.3)
-Now, the systems of linear equations in (4.3) can also be written as $\mathcal{A}{\bf p}=\mu\Sigma{\bf p},$ or, $\Sigma^{-1}\mathcal{A}{\bf p}=\mu{\bf p},$ where $\mu=\lambda/(1-s\lambda).$ By Lemma (3.2), we have $\Sigma^{-1}\mathcal{A}$ is positive stable, which implies $\Re(\mu)>0,$ hence the proof is completed.
+The system of linear equations in (4.3) can also be expressed as $\mathcal{A}{\bf p}=\mu\Sigma{\bf p},$ or $\Sigma^{-1}\mathcal{A}{\bf p}=\mu{\bf p},$ where $\mu=\lambda/(1-s\lambda).$ By Lemma (3.2), we have $\Sigma^{-1}\mathcal{A}$ is positive stable, which implies $\Re(\mu)>0,$ hence the proof is completed.
 $\blacksquare$
@@
 Theorem 4.2.
-Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0.$ Suppose $\lambda$ is a real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then
+Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices and $s>0.$ Suppose $\lambda$ is a real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then
 $\lambda\in\left(0,\frac{\xi_{\max}}{1+s\xi_{\max}}\right].$ (4.6)
@@
 First, we assume that $v=0.$ Subsequently, (4.2a) and (4.2c) are simplified to
-$Au=\lambda(\Lambda_{1}+sA)u,\quad\lambda\Lambda_{3}w=0,$ (4.7)
-respectively. The second equation of (4.7) gives $w=0,$ as $\Lambda_{3}$ is SPD. Multiplying $u^{T}$ to the first equation of (4.7) from the left, we get
+$Au=\lambda(\Lambda_{1}+sA)u\leavevmode\nobreak\ \leavevmode\nobreak\ {\text{and}}\leavevmode\nobreak\ \leavevmode\nobreak\ \lambda\Lambda_{3}w=0,$ (4.7)
+respectively. The second equation in (4.7) gives $w=0,$ as $\Lambda_{3}$ is SPD. Premultiplying by $u^{T}$ to the first equation of (4.7) , we get
 $\lambda=\frac{u^{T}Au}{u^{T}\Lambda_{1}u+su^{T}Au}=\frac{p}{1+sp},$
-where $p={\displaystyle{\frac{u^{T}Au}{u^{T}\Lambda_{1}u}}}>0.$ Now, for any nonzero vector $u\in{\mathbb{R}}^{n},$ we can have
+where $p={\displaystyle{\frac{u^{T}Au}{u^{T}\Lambda_{1}u}}}>0.$ Now, for any nonzero vector $u\in{\mathbb{R}}^{n},$ we have
 $\lambda_{\min}(\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}})\leq\frac{u^{T}Au}{u^{T}\Lambda_{1}u}=\frac{(\Lambda_{1}^{\frac{1}{2}}u)^{T}\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}\Lambda_{1}^{\frac{1}{2}}u}{(\Lambda_{1}^{\frac{1}{2}}u)^{T}\Lambda_{1}^{\frac{1}{2}}u}\leq\lambda_{\max}(\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}).$ (4.8)
@@
 $\lambda u^{T}Au=\lambda^{2}u^{T}\Lambda_{1}u+s\lambda^{2}u^{T}Au+(1-s\lambda)^{2}u^{T}B^{T}\Lambda_{2}^{-1}Bu,$ (4.12)
-which is equivalent to
+which is further equivalent to
 $\lambda^{2}-\lambda\frac{2sq+p}{s^{2}q+sp+1}+\frac{q}{s^{2}q+sp+1}=0,$ (4.13)
-where $q={\displaystyle{\frac{u^{T}B^{T}\Lambda_{2}^{-1}Bu}{u^{T}\Lambda_{1}u}}}.$ By solving (4.13), we obtain the following real solutions for $\lambda$
-$\lambda^{\pm}=\frac{2sq+p\pm\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)}$ (4.14)
+where $q={\displaystyle{\frac{u^{T}B^{T}\Lambda_{2}^{-1}Bu}{u^{T}\Lambda_{1}u}}}.$ By solving (4.13), we obtain the following real solutions for $\lambda$:
+$\lambda^{\pm}=\frac{2sq+p\pm\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)},$ (4.14)
 where $p^{2}-4q\geq 0.$ Consider the functions $\Phi_{1},\Phi_{2}:{\mathbb{R}}\times{\mathbb{R}}\times{\mathbb{R}}\longrightarrow{\mathbb{R}}$ defined by
@@
 respectively, where $p,s>0,$ $q\geq 0$ and $p^{2}>4q$.
-Then $\Phi_{1}$ and $\Phi_{2}$ is monotonically increasing in the argument $p>0,$ and $q\geq 0,$ and decreasing in the argument $q\geq 0$ and $p>0,$ respectively.
-Now, using (4.8), (4.14) and monotonicity of the functions $\Phi_{1}$ and $\Phi_{2},$ we get the following bounds
+Then $\Phi_{1}$ and $\Phi_{2}$ are strictly monotonically increasing in the argument $p>0$ and $q\geq 0,$ and decreasing in the argument $q\geq 0$ and $p>0,$ respectively.
+Now, using (4.8), (4.14) and monotonicity of the functions $\Phi_{1}$ and $\Phi_{2},$ we get the following bounds:
 $\lambda^{+}=\Phi_{1}(p,q,s)<\Phi_{1}(\lambda_{\max}(\Lambda^{-1}_{1}A),0,s),$ (4.15)
 $\Phi_{2}(\lambda_{\max}(\Lambda^{-1}_{1}A),0,s)<\lambda^{-}=\Phi_{2}(p,q,s).$ (4.16)
-Combining (4.15) and (4.16), we get the following bounds for $\lambda$
+Combining (4.15) and (4.16), we get the following bounds for $\lambda$:
 $0<\lambda^{\pm}<\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.17)
@@
 $\lambda>\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.19)
-Since by Remark 4.2, we get $01-s\lambda>$ 0 and the above assertion yields $\mu>\xi_{\max}.$ Hence, the matrix $\mu\Lambda_{1}-A$ is nonsingular and (4.2a) gives
+By Remark 4.2, we get $01-s\lambda>$ 0 and the above assertion yields $\mu>\xi_{\max}.$ Hence, the matrix $\mu\Lambda_{1}-A$ is nonsingular and (4.2a) gives
 $u=\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v.$ (4.20)
@@
 $(1-s\lambda)B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v+\frac{(1-s\lambda)}{\mu}C^{T}\Lambda_{3}^{-1}Cv=-\lambda\Lambda_{2}v.$ (4.21)
-Note that, $v\neq 0.$ Premultiplying by $v^{T}$ on the both sides of (4.21) leads to the following identity
+Note that, $v\neq 0.$ Thus, premultiplying by $v^{T}$ on the both sides of (4.21) leads to the following identity:
 $(1-s\lambda)v^{T}B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v+\lambda v^{T}\Lambda_{2}v=-\frac{(1-s\lambda)}{\mu}v^{T}C^{T}\Lambda_{3}^{-1}Cv.$ (4.22)
-On the other hand, $\mu\Lambda_{1}-A>\Lambda_{1}(\mu-\ \lambda_{\max}(\Lambda^{-1}A))I>0,$ and hence a SPD matrix.
-Consequently, the matrix $B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}$ is also a SPD matrix and this implies $v^{T}B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v>0,$ for all $v\not=0.$ This leads to a contradiction to (4.22) as $v^{T}C^{T}\Lambda_{3}^{-1}Cv\geq 0.$ Hence, the assumption (4.19) is not true and we get
+On the other hand, $\mu\Lambda_{1}-A>\Lambda_{1}(\mu-\ \lambda_{\max}(\Lambda^{-1}A))I>0$ and hence is a SPD matrix.
+Consequently, the matrix $B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}$ is also a SPD matrix and this implies $v^{T}B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v>0$ for all $v\not=0.$ This leads to a contradiction to (4.22) as $v^{T}C^{T}\Lambda_{3}^{-1}Cv\geq 0.$ Hence, the assumption (4.19) is not true and we get
 $\lambda\leq\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.23)
-Again, from ($(1)$) of Proposition 4.1, we have $\lambda>0.$ Combining the above together with (4.9) and (4.17), we obtained the desired bounds in (4.6) for $\lambda.$ Hence, the proof is concluded. ∎
+Again, we have $\lambda>0$ from $(1)$ of Proposition 4.1. Combining the above together with (4.9) and (4.17), we obtain the desired bounds in (4.6) for $\lambda.$ Hence, the proof is concluded. ∎
 Since the preconditioners ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{EGSS}$ are special cases of the PESS preconditioner, next, we obtain refined bounds for real eigenvalues of the SS and EGSS preconditioned matrices from Theorem 4.2.
@@
 Proof.
-Since ${\mathscr{P}}_{SS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, the bounds in (4.24) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha I$ and $s=\frac{1}{2}$ in Theorem 4.2.
+Since ${\mathscr{P}}_{SS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=1/2$ as discussed in Table 1, the bounds in (4.24) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha I$ and $s=1/2$ in Theorem 4.2.
 $\blacksquare$
@@
 Corollary 4.2.
-Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.11) with $\alpha>0$ and let $\lambda$ be an real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}.$ Then,
+Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.13) with $\alpha>0$ and let $\lambda$ be an real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}.$ Then
 $\lambda\in\left(0,\frac{2\tilde{\kappa}_{\max}}{\alpha+\tilde{\kappa}_{\max}}\right],$ (4.25)
@@
 Proof.
-Since ${\mathscr{P}}_{EGSS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, and the bounds are in (4.25) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha P$ and $s=\frac{1}{2}$ in Theorem 4.2.
+Since ${\mathscr{P}}_{EGSS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=1/2$ as discussed in Table 1 and the bounds in (4.25) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha P$ and $s=1/2$ in Theorem 4.2.
 $\blacksquare$
@@
 Theorem 4.3.
-Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices.
-Let $s>0$ and $\lambda$ be a non-real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0.$ Suppose ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the eigenvector corresponding to $\lambda.$ Then the following holds
+Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices.
+Let $s>0$ and $\lambda$ be a non-real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0.$ Suppose ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the eigenvector corresponding to $\lambda.$ Then the following holds:
   1. $(1)$
@@
     $\frac{\xi_{\min}\,\eta_{\min}}{2\left(\xi_{\max}^{2}+\eta_{\max}+\theta_{\max}\right)}\leq\Re\left(\frac{\lambda}{1-s\lambda}\right)\leq\frac{\xi_{\max}}{2}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \left|\Im\left(\frac{\lambda}{1-s\lambda}\right)\right|\leq\sqrt{\eta_{\max}+\theta_{\max}},$
-where $\xi_{\max},\xi_{\min},\eta_{\max},\eta_{\min}$ and $\theta_{\max}$ is defined as in (4.4) and (4.5).
+where $\xi_{\max},\xi_{\min},\eta_{\max},\eta_{\min}$ and $\theta_{\max}$ are defined as in (4.4) and (4.5).
 Proof.
 $(1)$   Let $\lambda$ be an eigenvalue of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0$ and the corresponding eigenvector is ${\bf p}=[u^{T},v^{T},w^{T}]^{T}.$ Then, the system of linear equations in (4.2a)-(4.2c) holds.
-We assert that, $v\neq 0,$ otherwise from (4.2a), we get $Au=\lambda(\Lambda_{1}+sA)u,$ which further implies $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}\bar{u}=\lambda\bar{u},$ where $\bar{u}=(\Lambda_{1}+sA)^{\frac{1}{2}}u.$ Since, $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}$ is a SPD matrix and by Lemma 4.1, $u\neq 0,$ and hence $\bar{u}\neq 0,$ which implies $\lambda$ is real, thus a contradiction to $\Im(\lambda)\neq 0.$
+We assert that, $v\neq 0,$ otherwise from (4.2a), we get $Au=\lambda(\Lambda_{1}+sA)u,$ which further implies $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}\bar{u}=\lambda\bar{u},$ where $\bar{u}=(\Lambda_{1}+sA)^{\frac{1}{2}}u.$ Since $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}$ is a SPD matrix and by Lemma 4.1 $u\neq 0$ and thus we have $\bar{u}\neq 0.$ This implies $\lambda$ is real, leading to a contradiction to $\Im(\lambda)\neq 0.$
 Initially, we consider the case $Cv=0.$ The aforementioned discussion suggests that $Au\neq\lambda(\Lambda_{1}+sA)u$ for any $u\neq 0.$ Consequently, $(\mu\Lambda_{1}-A)$ is nonsingular.
-Again, since $Cv=0,$ from (4.2c), and similar to (4.10)-(4.12), we obtain quadratic equation (4.13) in $\lambda$, which has complex solutions
-$\lambda^{\pm i}=\frac{2sq+p\pm i\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)}$ (4.26)
+Again, since $Cv=0,$ from (4.2c) and following a similar process as in (4.10)-(4.12), we obtain quadratic equation (4.13) in $\lambda$, which has complex solutions:
+$\lambda^{\pm i}=\frac{2sq+p\pm i\sqrt{4q-p^{2}}}{2(s^{2}q+sp+1)}$ (4.26)
 for $4q>p^{2}.$ From (4.26), we get
@@
 $\Psi_{1}(p,q,s)=\frac{q}{s^{2}q+sp+1},$
-where $p,q,s>0.$ Then $\Psi_{1}$ is monotonically increasing in the argument $q>0$ while decreasing in the argument $p>0.$ Due to the fact that $q\leq\lambda_{\max}(\Lambda_{1}^{-1}B^{T}\Lambda_{2}^{-1}B)=\eta_{\max},$ we get the following bounds
+where $p,q,s>0.$ Then $\Psi_{1}$ is monotonically increasing in the argument $q>0$ while decreasing in the argument $p>0.$ Due to the fact that $q\leq\lambda_{\max}(\Lambda_{1}^{-1}B^{T}\Lambda_{2}^{-1}B)=\eta_{\max},$ we get the following bounds:
 $|\lambda^{\pm i}|^{2}=\Psi_{1}(p,q,s)$$\leq$ $\Psi_{1}(\xi_{\min},\eta_{\max},s)=\frac{\eta_{\max}}{1+s\xi_{\min}+s^{2}\eta_{\max}}.$ (4.28)
-Again, considering the condition $q>(p^{2}/4),$ (4.27) adheres to the inequality $(p/(2+sp))\leq|\lambda^{\pm i}|.$ Notably, the function $\Psi_{2}:{\mathbb{R}}\longrightarrow{\mathbb{R}}$ defined by $\Psi_{2}(p)=p/(2+sp),$ where $p,s>0,$ is monotonically increasing in the argument $p>0,$ we obtain the following bound
+Again, considering the condition $q>(p^{2}/4),$ (4.27) adheres to the inequality $(p/(2+sp))\leq|\lambda^{\pm i}|.$ Notably, the function $\Psi_{2}:{\mathbb{R}}\longrightarrow{\mathbb{R}}$ defined by $\Psi_{2}(p)=p/(2+sp),$ where $p,s>0,$ is monotonically increasing in the argument $p>0,$ we obtain the following bound:
 $|\lambda^{\pm i}|=\Psi_{2}(p)\geq\Psi_{2}(\xi_{\min})=\frac{\xi_{\min}}{2+s\xi_{\min}}.$ (4.29)
@@
 $Bu+\frac{1}{\mu}C^{T}\Lambda_{3}^{-1}Cv=-\mu\Lambda_{2}v.$ (4.30)
-On the other side, nonsingularity of $(\mu\Lambda_{1}-A)$ and (4.2a) enables us the following identity
+On the other side, nonsingularity of $(\mu\Lambda_{1}-A)$ and (4.2a) enables us the following identity:
 $u=(\mu\Lambda_{1}-A)^{-1}B^{T}v.$ (4.31)
@@
 $v^{*}B(\mu\Lambda_{1}-A)^{-1}B^{T}v+\frac{1}{\mu}v^{*}C^{T}\Lambda_{3}^{-1}Cv=-\mu v^{*}\Lambda_{2}v.$ (4.33)
-Consider the following eigenvalue decomposition of $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$
+Consider the following eigenvalue decomposition of $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$:
 $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}=V\mathcal{D}V^{T},$
-where $\mathcal{D}=\mathrm{diag}(\theta_{i})\in{\mathbb{R}}^{n\times n}$ and $\theta_{i}>0,\,i=1,2,\ldots,n$ are in $\sigma(\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}})$ and $V\in{\mathbb{R}}^{n\times n}$ is an orthonormal matrix.
+where $\mathcal{D}=\mathrm{diag}(\theta_{i})\in{\mathbb{R}}^{n\times n}$ and $\theta_{i}>0,\,i=1,2,\ldots,n,$ are in $\sigma(\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}})$ and $V\in{\mathbb{R}}^{n\times n}$ is an orthonormal matrix.
 Then, (4.33) can be written as
@@
 Since $(\mu I-\mathcal{D})^{-1}=\Theta_{1}-{\bf{i\mkern 1.0mu}}\Im(\mu)\Theta_{2},$ where ${\bf{i\mkern 1.0mu}}$ denotes the imaginary unit and
-$\Theta_{1}={\displaystyle{\mathrm{diag}\left(\frac{\Re(\mu)-\theta_{i}}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right),\Theta_{2}=\mathrm{diag}\left(\frac{1}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right)}},$
-from (4.34), and the fact $\Im(\mu)\neq 0,$ we get
+$\Theta_{1}={\displaystyle{\mathrm{diag}\left(\frac{\Re(\mu)-\theta_{i}}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right),\leavevmode\nobreak\ \Theta_{2}=\mathrm{diag}\left(\frac{1}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right)}},$
+from (4.34) and the fact $\Im(\mu)\neq 0,$ we get
 $v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\Theta_{1}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v+\frac{\Re(\mu)}{|\mu|^{2}}v^{*}C^{T}\Lambda_{3}^{-1}Cv=-\Re(\mu)v^{*}\Lambda_{2}v,$ (4.35)
@@
 $2\Re(\mu)\leq\xi_{\max}-\frac{\xi_{\max}}{|\mu|^{2}}\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}\leq\xi_{\max}.$ (4.39)
-Hence, (4.39) yields the following bound
+Hence, (4.39) yields the following bound:
 $\Re(\mu)\leq\frac{\xi_{\max}}{2}.$ (4.40)
@@
 Therefore, by (4.42) and from the bound in (4.37), we obtain
-$\mathcal{X}\geq\frac{I}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}$ (4.43)
-Combining (4.41) and (4.43) leads to the following bounds
+$\mathcal{X}\geq\frac{I}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}.$ (4.43)
+Combining (4.41) and (4.43) leads to the following bounds:
 $2\Re(\mu)$$\geq$ $\frac{\xi_{\min}}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}\frac{v^{*}B\Lambda_{1}^{-1}B^{T}v}{v^{*}\Lambda_{2}v}\geq\frac{\xi_{\min}\eta_{\min}}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}.$ (4.44)
@@
 From the bounds in $(2)$ of Theorem 4.3, we obtain
-${\frac{1}{|\tau|^{2}}\left(\frac{\xi_{\min}\eta_{\min}}{2(\xi_{\max}^{2}+\eta_{\max}+\theta_{\max})}+\frac{1}{s}\right)\leq\frac{1}{s}-\Re(\lambda)\leq\frac{1}{|\tau|^{2}}(\frac{\xi_{\max}}{2}+\frac{1}{s})}$
+${\frac{1}{|\tau|^{2}}\left(\frac{\xi_{\min}\eta_{\min}}{2(\xi_{\max}^{2}+\eta_{\max}+\theta_{\max})}+\frac{1}{s}\right)\leq\frac{1}{s}-\Re(\lambda)\leq\frac{1}{|\tau|^{2}}\left(\frac{\xi_{\max}}{2}+\frac{1}{s}\right)}$
 and
@@
 Corollary 4.3.
-Let ${\mathscr{P}}_{SS}$ be defined as in (1.10) with $\alpha>0$ and let $\lambda$ be a non-real eigenvalue of the SS preconditioned matrix ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector.
+Let ${\mathscr{P}}_{SS}$ be defined as in (1.12) with $\alpha>0$ and let $\lambda$ be a non-real eigenvalue of the SS preconditioned matrix ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector.
 Then,
@@
 Proof.
-Since ${\mathscr{P}}_{SS}$ is a special case of ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, then desired bounds will be obtained by setting $\Lambda_{1}=\Lambda_{2}=\Lambda_{3}=\frac{1}{2}\alpha I$ and $s=\frac{1}{2}$ into Theorem 4.3.
+Since ${\mathscr{P}}_{SS}$ is a special case of ${\mathscr{P}}_{PESS}$ for $s=1/2$ as discussed in Table 1, then desired bounds will be obtained by setting $\Lambda_{1}=\Lambda_{2}=\Lambda_{3}=\frac{1}{2}\alpha I$ and $s=1/2$ into Theorem 4.3.
 $\blacksquare$
 Corollary 4.4.
-Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.11), and let $P,Q$ and $W$ are SPD matrices with $\alpha,\beta,\gamma>0.$ Let $\lambda$ be a non-real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector.
+Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.13) and let $P,Q$ and $W$ are SPD matrices with $\alpha,\beta,\gamma>0.$ Let $\lambda$ be a non-real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector.
 Then $(1)$ If $Cv=0,$ $\lambda$ satisfies
@@
 Proof.
-Since ${\mathscr{P}}_{EGSS}$ is a special case of ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, then the desired bounds will be obtained by setting $\Lambda_{1}=\frac{1}{2}\alpha P,$ $\Lambda_{2}=\frac{1}{2}\beta Q$, $\Lambda_{3}=\frac{1}{2}\gamma W$ and $s=\frac{1}{2}$ in Theorem 4.3.
+Since ${\mathscr{P}}_{EGSS}$ is a special case of ${\mathscr{P}}_{PESS}$ for $s=1/2$ as discussed in Table 1, then the desired bounds will be obtained by setting $\Lambda_{1}=\frac{1}{2}\alpha P,$ $\Lambda_{2}=\frac{1}{2}\beta Q$, $\Lambda_{3}=\frac{1}{2}\gamma W$ and $s=1/2$ in Theorem 4.3.
 $\blacksquare$
@@
 where
-$\mathfrak{L}=\left[\begin{matrix}I&0&0\\ -BA^{-1}&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{U}=\left[\begin{matrix}I&A^{-1}B^{T}&0\\ 0&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{D}=\left[\begin{matrix}A&0&0\\ 0&Q&-C^{T}\\ 0&C&0\end{matrix}\right],\widetilde{\mathfrak{D}}=\left[\begin{matrix}sA&0&0\\ 0&\Lambda_{2}+sQ&-C^{T}\\ 0&C&\Lambda_{3}\end{matrix}\right],$
+$\mathfrak{L}=\left[\begin{matrix}I&0&0\\ -BA^{-1}&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{U}=\left[\begin{matrix}I&A^{-1}B^{T}&0\\ 0&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{D}=\left[\begin{matrix}A&0&0\\ 0&Q&-C^{T}\\ 0&C&0\end{matrix}\right],\widetilde{\mathfrak{D}}=\left[\begin{matrix}sA&0&0\\ 0&\Lambda_{2}+sQ&-sC^{T}\\ 0&sC&\Lambda_{3}\end{matrix}\right]$
 and $Q=BA^{-1}B^{T}.$ Using the decomposition in (5.2), we have
 ${\mathscr{P}}^{-1}_{LPESS}\mathcal{A}=\mathfrak{U}^{-1}\left[\begin{matrix}s^{-1}I&0\\ 0&M^{-1}K\end{matrix}\right]\mathfrak{U},$ (5.3)
-where $M=\left[\begin{matrix}\Lambda_{2}+sQ&-sC^{T}\\ sC&\Lambda_{3}\end{matrix}\right]$ and $K=\left[\begin{matrix}Q&-C^{T}\\ C&0\end{matrix}\right].$ Since, ${\mathscr{P}}^{-1}_{LPESS}\mathcal{A}$ is similar to $\left[\begin{matrix}s^{-1}I&0\\ 0&M^{-1}K\end{matrix}\right],$ ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has eigenvalue $\frac{1}{s}$ with multiplicity $n$ and the remaining eigenvalues satisfies the generalized eigenvalue problem $Kp=\lambda Mp.$
+where $M=\left[\begin{matrix}\Lambda_{2}+sQ&-sC^{T}\\ sC&\Lambda_{3}\end{matrix}\right]$ and $K=\left[\begin{matrix}Q&-C^{T}\\ C&0\end{matrix}\right].$ Since, ${\mathscr{P}}^{-1}_{LPESS}\mathcal{A}$ is similar to $\left[\begin{matrix}s^{-1}I&0\\ 0&M^{-1}K\end{matrix}\right],$ ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has eigenvalue $1/s$ with multiplicity $n$ and the remaining eigenvalues satisfies the generalized eigenvalue problem $Kp=\lambda Mp.$
 Theorem 5.1.
-Let $A\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}$ and $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices.
-Assume that $s>0,$ then LPESS preconditioned matrix ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has $n$ eigenvalue equal to $\frac{1}{s}.$ The remaining $m+p$ eigenvalues satisfies the following:
+Let $A\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}$ and $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices.
+Assume that $s>0,$ then LPESS preconditioned matrix ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has $n$ repeated eigenvalues equal to $1/s.$ The remaining $m+p$ eigenvalues satisfies the following:
   1. $(1)$
     real eigenvalues located in the interval
-    $\Big{[}\min\Big{\{}\frac{\vartheta_{\min}}{1+s\vartheta_{\min}},\frac{\theta_{\min}}{\vartheta_{\max}+s\theta_{\min}}\Big{\}},\frac{\vartheta_{\max}}{1+s\vartheta_{\max}}\Big{]}.$
+    $\Big{[}\min\Big{\{}\frac{\vartheta_{\min}}{1+s\vartheta_{\min}},\frac{\tilde{\theta}_{\min}}{\vartheta_{\max}+s\tilde{\theta}_{\min}}\Big{\}},\frac{\vartheta_{\max}}{1+s\vartheta_{\max}}\Big{]}.$
   2. $(2)$
     If $\lambda$ is any non-real eigenvalue (i.e., $\Im(\lambda)\neq 0$), then
-    $\frac{\vartheta_{\min}}{2+s\vartheta_{\min}}\leq|\lambda|\leq\sqrt{\frac{\theta_{\max}}{1+s\vartheta_{\min}+s^{2}\theta_{\max}}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \frac{1}{s(1+s\sqrt{\theta_{\max}})}\leq|\lambda-\frac{1}{s}|\leq\frac{2}{s(2+s\vartheta_{\min})},$
-where $\vartheta_{\max}:=\lambda_{\max}(\Lambda_{2}^{-1}Q),$ $\vartheta_{\min}:=\lambda_{\min}(\Lambda_{2}^{-1}Q)$ and $\theta_{\min}:=\lambda_{\min}(\Lambda_{3}^{-1}C\Lambda_{1}^{-1}C^{T}).$
+    $\frac{\vartheta_{\min}}{2+s\vartheta_{\min}}\leq|\lambda|\leq\sqrt{\frac{\tilde{\theta}_{\max}}{1+s\vartheta_{\min}+s^{2}\tilde{\theta}_{\max}}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \frac{1}{s(1+s\sqrt{\tilde{\theta}_{\max}})}\leq|\lambda-\frac{1}{s}|\leq\frac{2}{s(2+s\vartheta_{\min})},$
+where $\vartheta_{\max}:=\lambda_{\max}(\Lambda_{2}^{-1}Q),$ $\vartheta_{\min}:=\lambda_{\min}(\Lambda_{2}^{-1}Q)$ $\tilde{\theta}_{\max}:=\lambda_{\max}(\Lambda_{3}^{-1}C\Lambda_{1}^{-2}C^{T})$ and $\tilde{\theta}_{\min}:=\lambda_{\min}(\Lambda_{3}^{-1}C\Lambda_{1}^{-2}C^{T}).$
 Proof.
@@
 $M^{-1}K=\mathcal{F}^{-1}\widetilde{M}^{-1}\widetilde{K}\mathcal{F},$ (5.4)
-where $\mathcal{F}=\left[\begin{matrix}I&0\\ 0&-I\end{matrix}\right],$ $\widetilde{M}=\left[\begin{matrix}\Lambda_{2}+sQ&C^{T}\\ -C&\Lambda_{3}\end{matrix}\right]$ and $\widetilde{K}=\left[\begin{matrix}Q&C^{T}\\ -C&0\end{matrix}\right].$ Hence, $M^{-1}K$ is similar to $\widetilde{M}^{-1}\widetilde{K}.$ Now, $\widetilde{M}^{-1}\widetilde{K}$ is in the form of the preconditioned matrix discussed in [28].
+where $\mathcal{F}=\left[\begin{matrix}I&0\\ 0&-I\end{matrix}\right],$ $\widetilde{M}=\left[\begin{matrix}\Lambda_{2}+sQ&sC^{T}\\ -sC&\Lambda_{3}\end{matrix}\right]$ and $\widetilde{K}=\left[\begin{matrix}Q&C^{T}\\ -C&0\end{matrix}\right].$ Hence, $M^{-1}K$ is similar to $\widetilde{M}^{-1}\widetilde{K}.$ Now, $\widetilde{M}^{-1}\widetilde{K}$ is in the form of the preconditioned matrix discussed in [42].
 Therefore, by applying the Corollary 4.1 and Theorem 4.3 of [42], we obtain the desired bounds in $(1)$ and $(2).$blacksquare
+6 The strategy of parameter selection
+It is worth noting that the efficiency of the proposed PESS preconditioner depends on the selection of involved SPD matrices $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ and a positive real parameter $s.$ In this section, motivated by [42, 43, 44], we discuss a practical way to choose the parameter $s$ and the SPD matrices for the effectiveness of the proposed preconditioners.
+It is well acknowledged that the optimal parameter of PESS iteration process is obtained when $\vartheta(\mathcal{T})$ is minimized [43].
+For achieving this, similar to the approach in [44], we first define a function $\varphi(s)=\|\mathcal{Q}_{PESS}\|^{2}_{F}$ depending on the parameter $s$.
+Our aim is to minimize $\varphi(s).$ Then after some straightforward calculations, we obtain
+$\varphi(s)=\|\mathcal{Q}_{PESS}\|^{2}_{F}$ $=tr(\mathcal{Q}^{T}_{PESS}\mathcal{Q}_{PESS})$
+$=\|\Lambda_{1}\|^{2}_{F}+\|\Lambda_{2}\|^{2}_{F}+\|\Lambda_{3}\|^{2}_{F}+(s-1)^{2}\|A\|^{2}_{F}+2(s-1)tr(\Lambda_{1}A)$
+$\quad+2(s-1)^{2}\|B\|_{F}^{2})+2(s-1)^{2}\|C\|_{F}^{2}),$
+where $tr(\cdot)$ and $\|\cdot\|_{F}$ denote the trace and the Frobenious norm of a matrix, respectively.
+Now we choose the parameter $s$ and SPD matrices $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ to make $\varphi(s)$ as small as possible.
+Since $\|A\|^{2}_{F},\|B\|^{2}_{F},\|C\|^{2}_{F}$ and $tr(\Lambda_{1}A)$ are positive, we can select $s=1,$ then $\varphi(s)=\|\Lambda_{1}\|^{2}_{F}+\|\Lambda_{2}\|^{2}_{F}+\|\Lambda_{3}\|^{2}_{F}.$ Thus if we choose $\|\Lambda_{1}\|_{2},\|\Lambda_{2}\|_{2},\|\Lambda_{3}\|_{2}\,\rightarrow 0,$ we have $\varphi(s)\rightarrow 0$ and consequently, $\mathcal{Q}_{PESS}\rightarrow 0.$
+On the other hand, motivated by [45, 46], we discuss another strategy for choosing the parameter $s$.
+Notice that in Algorithm 1, we need to solve two linear system with coefficient matrices $\widehat{X}=\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C$ and $\widetilde{A}=\Lambda_{1}+sA+s^{2}B^{T}\widehat{X}B.$ Similar to [45, 46], we choose $s$ and $\|\Lambda_{2}\|_{2}$ as follows:
+$s=\sqrt{\frac{\|\Lambda_{2}\|_{2}}{\|C^{T}\Lambda_{3}^{-1}C\|_{2}}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \|\Lambda_{2}\|_{2}=\frac{\|B\|_{2}^{4}}{{4\|C^{T}\Lambda_{3}^{-1}C\|_{2}\|A\|_{2}^{2}}},$ (6.1)
+which balance the matrices $\Lambda_{2}$ and $C^{T}\Lambda_{3}^{-1}C$ in $\widehat{X}$ and the matrices $A$ and $C^{T}\Lambda_{3}^{-1}C$ in $\widetilde{A}$.
+In this case, we denote $s$ by $s_{est}$ and $\|\Lambda_{2}\|_{2}$ by $\beta_{est}.$ Numerical results are presented in Section 7 to demonstrate the effectiveness of ${\mathscr{P}}_{PESS}$ for the above choices of the parameters.
+In the sequel, we can rewrite the PESS preconditioner as ${\mathscr{P}}_{PESS}=s\widetilde{{\mathscr{P}}}_{PESS},$ where
+$\widetilde{{\mathscr{P}}}_{PESS}=\left[\begin{matrix}\frac{1}{s}\Lambda_{1}+A&B^{T}&0\\ -B&\frac{1}{s}\Lambda_{2}&-C^{T}\\ 0&C&\frac{1}{s}\Lambda_{3}\end{matrix}\right].$
+Since the prefactor $s$ has not much effect on the performance of PESS preconditioner, investigating the optimal parameters of ${\mathscr{P}}_{PESS}$ and $\widetilde{{\mathscr{P}}}_{PESS}$ are equivalent.
+A general criterion for a preconditioner to perform efficiently is that it should closely approximate the coefficient matrix $\mathcal{A}$ [10].
+Consequently, the difference $\widetilde{{\mathscr{P}}}_{PESS}-\mathcal{A}={\displaystyle{\frac{1}{s}\Sigma}}$ approaches zero matrix as $s$ tends to positive infinity for fixed $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}.$ Thus, the preconditioner PESS shows enhance efficiency for large value of $s.$ However, $s$ can not be too large as the coefficient matrix $\widetilde{A}$ of the linear subsystem in step $3$ of Algorithm 1 becomes very-ill conditioned.
+Similar investigations also hold for LPESS preconditioner.
+Nevertheless, in Figure 9, we show the adaptability of the PESS and LPESS preconditioners by varying the parameter $s.$
 7 Numerical experiments
 In this section, we conduct a few numerical experiments to showcase the superiority and efficiency of the proposed PESS and LPESS preconditioners over the existing preconditioners to enhance the convergence speed of the Krylov subspace iterative process to solve three-by-three block SPPs.
-Our study involves a comparative analysis among GMRES process and the PGMRES process employing the proposed preconditioners ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS},$ and the existing preconditioners ${\mathscr{P}}_{BD}$, ${\mathscr{P}}_{IBD}$, ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS}$ and ${\mathscr{P}}_{EGSS}$ (baseline preconditioners).
+Our study involves a comparative analysis among GMRES process and the PGMRES process employing the proposed preconditioners ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}$ and the existing baseline preconditioners ${\mathscr{P}}_{BD}$, ${\mathscr{P}}_{IBD}$, ${\mathscr{P}}_{MAPSS}$, ${\mathscr{P}}_{SL},$ ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS},$ ${\mathscr{P}}_{EGSS}$ and ${\mathscr{P}}_{RPGSS}$.
 The numerical results are reported from the aspect of iteration counts (abbreviated as “IT”) and elapsed CPU times in seconds (abbreviated as “CPU”).
-Each subsystem involving $\widehat{A}$ and $\widehat{X}$ featured in Algorithm 1 is precisely solved by applying the Cholesky factorization of the coefficient matrices.
-For all iterative process, the initial guess vector is ${\bf u}_{0}=0\in{\mathbb{R}}^{n+m+p}$, and the termination criterion is
+Each subsystem involving $\widehat{X}$ and $\widetilde{A}$ featured in Algorithm 1 are precisely solved by applying the Cholesky factorization of the coefficient matrices.
+For all iterative process, the initial guess vector is ${\bf u}_{0}=0\in{\mathbb{R}}^{n+m+p}$ and the termination criterion is
 ${\tt RES}:=\frac{\|\mathcal{A}{\bf u}^{k+1}-{\bf d}\|_{2}}{\|{\bf d}\|_{2}}<10^{-6}.$
 The vector ${\bf d}\in{\mathbb{R}}^{n+m+p}$ is chosen so that the exact solution of the system (1.1) is ${\bf u}_{*}=[1,1,\ldots,1]^{T}\in{\mathbb{R}}^{n+m+p}.$ All numerical tests are run in MATLAB (version R2023a) on a Windows 11 operating system with Intel(R) Core(TM) i7-10700 CPU, 2.90GHz, 16 GB memory.
-Before proceeding, we discuss the optimal choice for the parameters of PESS preconditioner.
-We can rewrite the PESS preconditioner as ${\mathscr{P}}_{PESS}=s\widetilde{{\mathscr{P}}}_{PESS},$ where $\widetilde{{\mathscr{P}}}_{PESS}=\left[\begin{matrix}\frac{1}{s}\Lambda_{1}+A&B^{T}&0\\ -B&\frac{1}{s}\Lambda_{2}&-C^{T}\\ 0&C&\frac{1}{s}\Lambda_{3}\end{matrix}\right].$ Since, the pre-factor $s$ has not much effect on the performance of PESS preconditioner, investigating the optimal parameters of ${\mathscr{P}}_{PESS}$ and $\widetilde{{\mathscr{P}}}_{PESS}$ are equivalent.
-As we know, a preconditioner performs efficiently when it is as close as possible to the coefficient matrix $\mathcal{A},$ observe that $\widetilde{{\mathscr{P}}}_{PESS}-\mathcal{A}={\displaystyle{\frac{1}{s}\Sigma}},$ which tends to a zero matrix as $s$ tends to positive infinity for fixed $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}.$ Thus, PESS encounters enhance efficiency for large value of $s.$ However, $s$ can not be too large as the coefficient matrix $\Lambda_{1}+sA+s^{2}B^{T}(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)^{-1}B$ of the linear subsystem in step $3$ of Algorithm 1 becomes very-ill conditioned.
-Similar investigations also hold for LPESS preconditioner.
-Nevertheless, in Figure 7, we show the adaptability of the PESS and LPESS preconditioners by varying the parameter $s.$
 Example 1.
-We consider the three-by-three block SPP (1.1) taken from [18], where
-$A=\left[\begin{matrix}I\otimes G+G\otimes I&0\\ 0&I\otimes G+G\otimes I\end{matrix}\right]\in{\mathbb{R}}^{2l^{2}\times 2l^{2}},\,\,B=\left[\begin{matrix}I\otimes F&F\otimes I\end{matrix}\right]\in{\mathbb{R}}^{l^{2}\times 2l^{2}},$ and $C=E\otimes F\in{\mathbb{R}}^{l^{2}\times l^{2}},$ where $G=\frac{1}{(l+1)^{2}}\,\mathrm{tridiag}(-1,2,-1)\in{\mathbb{R}}^{l\times l},\quad F=\frac{1}{l+1}\,\mathrm{tridiag}(0,1,-1)\in{\mathbb{R}}^{l\times l}$ , and $E=\mathrm{diag}(1,l+1,\ldots,l^{2}-l+1)\in{\mathbb{R}}^{l\times l}.$ Here, $E\otimes F$ denotes the Kronecker product of two matrices $E$ and $F,$ and $\mathrm{tridiag}(a,b,c)\in{\mathbb{R}}^{l\times l}$ denotes the $l\times l$ tridiagonal matrix with diagonal, subdiagonal, and superdiagonal entries are $a,b,$ and $c,$ respectively.
-For this problem, the size of the matrix $\mathcal{A}$ is $4l^{2}.$
-Table 2: Numerical results of GMRES, BD, IBD, SS, RSS, EGSS and PESS PGMRES process for Example 1.
-Process $l$ $16$ $32$ $48$ $64$ $80$
-size($\mathcal{A}$) $1024$ $4096$ $9216$ $16384$ $25600$
+Table 2: Numerical results of GMRES, BD, IBD, MAPSS, SL, SS, RSS, EGSS, RPGSS, PESS and LPESS PGMRES processes for Example 1.
+Process $l$ $16$ $32$ $48$ $64$ $80$ 128
+size($\mathcal{A}$) $1024$ $4096$ $9216$ $16384$ $25600$ $65536$
 […]
-CPU $0.6607$ $101.7659$ $989.0242$ $2116.1245$ $3755.9251$
-RES $078.2852e-$ 07 $079.9189e-$ 07 $079.8389e-$ 07 $031.0813e-$ 03 $031.9862e-$ 03
-BD IT $4$ $4$ $4$ $4$ $4$
-CPU 0.0611 $2.3135$ $17.0971$ $105.4514$ $337.4322$
-RES $131.2728e-$ 13 $131.7577e-$ 13 $136.6653e-$ 13 $126.0211e-$ 12 $122.6610e-$ 12
-IBD IT $22$ $22$ $21$ $21$ $21$
-CPU 0.0702 $1.1711$ $9.9850$ $44.2197$ $141.1108$
-RES $074.0221e-$ 07 $074.6059e-$ 07 $079.1711e-$ 07 $077.8518e-$ 07 $076.8357e-$ 07
+CPU $0.6607$ $101.7659$ $989.0242$ $2116.1245$ $3755.9251$ $8694.5762$
+RES $078.2852e-$ 07 $079.9189e-$ 07 $079.8389e-$ 07 $031.0813e-$ 03 $031.9862e-$ 03 $0.0037$
+BD IT $4$ $4$ $4$ $4$ $4$ $4$
+CPU 0.0611 $2.3135$ $17.0971$ $105.4514$ $337.4322$ $1355.6953$
+RES $131.2728e-$ 13 $131.7577e-$ 13 $136.6653e-$ 13 $126.0211e-$ 12 $122.6610e-$ 12 $076.4572e-$ 07
+IBD IT $22$ $22$ $21$ $21$ $21$ $27$
+CPU 0.0702 $1.1711$ $9.9850$ $44.2197$ $141.1108$ $3064.8487$
+RES $074.0221e-$ 07 $074.6059e-$ 07 $079.1711e-$ 07 $077.8518e-$ 07 $076.8357e-$ 07 $079.8667e-$ 07
+MAPSS IT $5$ $5$ $6$ $6$ $6$ $7$
+CPU $0.21299$ $1.2038$ $8.0537$ $36.5074$ $116.7495$ $514.1132$
+RES $074.6434e-$ 07 $079.0780e-$ 07 $073.8651e-$ 07 $073.8350e-$ 07 $074.3826e-$ 07 $072.4983e-$ 07
+SL IT $6$ $6$ $5$ $5$ $5$ $4$
+CPU $0.19624$ $1.0819$ $6.3741$ $30.6976$ $102.0981$ $481.7656$
+RES $082.5612e-$ 08 $087.4194e-$ 08 $089.4761e-$ 08 $093.7303e-$ 09 $091.7982e-$ 09 $086.1568e-$ 08
 Case I
-SS IT $4$ $4$ $4$ $4$ $4$
-CPU $0.2415$ $1.3375$ $6.7059$ $33.6902$ $111.3118$
-RES $087.7528e-$ 08 $085.5120E-$ 08 $084.5134E-$ 08 $083.9157E-$ 08 $083.5073E-$ 08
-RSS IT $4$ $4$ $4$ $4$ $4$
-CPU $0.2776$ $1.0467$ $6.5207$ $32.4428$ $110.4587$
-RES $088.0898e-$ 08 $086.0033e-$ 08 $084.9839e-$ 08 $084.3510e-$ 08 $083.9097e-$ 08
-EGSS IT $4$ $4$ $4$ $4$ $4$
-CPU $0.3061$ $1.2771$ $6.9552$ $32.8960$ $119.7763$
-RES $105.9583e-$ 10 $104.4128e-$ 10 $103.6626e-$ 10 $103.2009e-$ 10 $102.8807e-$ 10
-PESS IT 2 2 2 2 2
-CPU $0.2285$ $\bf 0.9854$ $\bf 4.8795$ $\bf 23.3687$ $\bf 79.9355$
-$s=12$ RES $073.1630e-$ 07 $072.3239e-$ 07 $071.9486e-$ 07 $071.7260e-$ 07 $071.5754e-$ 07
-LPESS IT 2 2 2 2 2
-CPU $0.2834$ $\bf 0.7989$ $\bf 4.5599$ $\bf 21.1664$ $\bf 72.8376$
-$s=12$ RES $073.1180e-$ 07 $072.2463e-$ 07 $071.8481e-$ 07 $071.6071e-$ 07 $071.4411e-$ 07
+SS IT $4$ $4$ $4$ $4$ $4$ $4$
+CPU $0.2415$ $1.3375$ $6.7059$ $33.6902$ $111.3118$ $439.1221$
+RES $087.7528e-$ 08 $085.5120e-$ 08 $084.5134e-$ 08 $083.9157e-$ 08 $083.5073e-$ 08 $082.7836e-$ 08
+RSS IT $4$ $4$ $4$ $4$ $4$ $4$
+CPU $0.2776$ $1.0467$ $6.5207$ $32.4428$ $110.4587$ $432.6755$
+RES $088.0898e-$ 08 $086.0033e-$ 08 $084.9839e-$ 08 $084.3510e-$ 08 $083.9097e-$ 08 $083.1111e-$ 08
+EGSS IT $4$ $4$ $4$ $4$ $4$ $4$
+CPU $0.3061$ $1.2771$ $6.9552$ $32.8960$ $119.7763$ $435.0198$
+RES $105.9583e-$ 10 $104.4128e-$ 10 $103.6626e-$ 10 $103.2009e-$ 10 $102.8807e-$ 10 $081.9145e-$ 08
+RPGSS IT $4$ $4$ $4$ $4$ $4$ $3$
+CPU $0.2249$ $1.1312$ $7.0322$ $35.0140$ $130.2593$ $371.4522$
+RES $105.9497e-$ 10 $104.4042e-$ 10 $103.6494e-$ 10 $103.1838e-$ 10 $102.8603e-$ 10 $079.9326e-$ 07
+PESS IT 2 2 2 2 2 $\bf{2}$
+CPU $0.2285$ $\bf 0.9854$ $\bf 4.8795$ $\bf 23.3687$ $\bf 79.9355$ $\bf 283.8561$
+$s=12$ RES $073.1630e-$ 07 $072.3239e-$ 07 $071.9486e-$ 07 $071.7260e-$ 07 $071.5754e-$ 07 $071.3135e-$ 07
+LPESS IT 2 2 2 2 2 $\bf 2$
+CPU $0.2834$ $\bf 0.7989$ $\bf 4.5599$ $\bf 21.1664$ $\bf 72.8376$ $\bf 268.3114$
+$s=12$ RES $073.1180e-$ 07 $072.2463e-$ 07 $071.8481e-$ 07 $071.6071e-$ 07 $071.4411e-$ 07 $071.1441e-$ 07
 Case II
-SS IT $7$ $7$ $7$ $7$ $7$
-CPU 0.2780 $1.6189$ $10.6707$ $54.4493$ $184.6025$
-RES $076.7967e-$ 07 $074.9738e-$ 07 $074.1383e-$ 07 $073.6189e-$ 07 $073.2556e-$ 07
-RSS IT $7$ $7$ $7$ $7$ $7$
-CPU $0.3248$ $1.4644$ $10.4985$ $52.4559$ $177.1895$
-RES $075.2397e-$ 07 $073.6832e-$ 07 $072.7532e-$ 07 $072.1869e-$ 07 $071.8240e-$ 07
-EGSS IT $5$ $5$ $4$ $4$ $4$
-CPU $0.2588$ $1.3792$ $6.7546$ $34.8358$ $118.2298$
-RES $086.8491e-$ 08 $071.5239e-$ 07 $079.7099e-$ 07 $076.3466e-$ 07 $074.8159e-$ 07
-PESS IT $3$ $3$ $3$ $3$ $3$
-CPU $0.2971$ $1.1597$ $6.6682$ $31.2670$ $103.8616$
-$s=12$ RES $087.4100e-$ 08 $087.4642e-$ 08 $087.3327e-$ 08 $086.9803e-$ 08 $086.1920e-$ 08
-LPESS IT $3$ $3$ $3$ $3$ $3$
-CPU $0.2226$ $0.9905$ $6.1339$ $30.0975$ $98.6443$
-$s=12$ RES $091.1683e-$ 09 $092.3215e-$ 09 $093.1540e-$ 09 $093.3907e-$ 09 $093.0271e-$ 09
+SS IT $7$ $7$ $7$ $7$ $7$ $7$
+CPU 0.2780 $1.6189$ $10.6707$ $54.4493$ $184.6025$ $703.1082$
+RES $076.7967e-$ 07 $074.9738e-$ 07 $074.1383e-$ 07 $073.6189e-$ 07 $073.2556e-$ 07 $072.5956e-$ 07
+RSS IT $7$ $7$ $7$ $7$ $7$ $7$
+CPU $0.3248$ $1.4644$ $10.4985$ $52.4559$ $177.1895$ $690.1611$
+RES $075.2397e-$ 07 $073.6832e-$ 07 $072.7532e-$ 07 $072.1869e-$ 07 $071.8240e-$ 07 $071.2720e-$ 07
+EGSS IT $5$ $5$ $4$ $4$ $4$ $4$
+CPU $0.2588$ $1.3792$ $6.7546$ $34.8358$ $118.2298$ $439.7907$
+RES $086.8491e-$ 08 $071.5239e-$ 07 $079.7099e-$ 07 $076.3466e-$ 07 $074.8159e-$ 07 $074.1376e-$ 07
+RPGSS IT $4$ $4$ $4$ $4$ $4$ $3$
+CPU $0.2445$ $1.2987$ $8.1982$ $28.5129$ $104.9127$ $397.4766$
+RES $085.2642e-$ 08 $071.1366e-$ 07 $087.1116e-$ 08 $078.2780e-$ 07 $075.6396e-$ 07 $072.5131e-$ 07
+PESS IT $3$ $3$ $3$ $3$ $3$ $3$
+CPU $0.2971$ $1.1597$ $6.6682$ $31.2670$ $103.8616$ $371.1214$
+$s=12$ RES $087.4100e-$ 08 $087.4642e-$ 08 $087.3327e-$ 08 $086.9803e-$ 08 $086.1920e-$ 08 $084.0967e-$ 08
+LPESS IT $3$ $3$ $3$ $3$ $3$ $3$
+CPU $0.2226$ $0.9905$ $6.1339$ $30.0975$ $98.6443$ $359.8141$
+$s=12$ RES $091.1683e-$ 09 $092.3215e-$ 09 $093.1540e-$ 09 $093.3907e-$ 09 $093.0271e-$ 09 $091.6141e-$ 09
 […]
 […]
-We consider the parameter choices for the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS},{\mathscr{P}}_{EGSS},{\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}$ in two cases.
-  • In Case I: $\alpha=0.1$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=0.1,\beta=1,\gamma=0.001$ and $P=I,Q=I,W=I$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=I,\Lambda_{2}=I,\Lambda_{3}=0.001I$ for ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}.$
-[…]
-The parameter selection in Case II is made as in [31].
-The optimal value (experimentally) for $s$ in the range $[10,20]$ is determined to be $12$ for minimal CPU times.
-Following [18], selection of $\widehat{A}$ and $\widehat{S}$ (SPD approximations of $A$ and $S,$ respectively) for IBD preconditioner are done as follows:
+Problem formulation: We consider the three-by-three block SPP (1.1) taken from [8] with
+$A=\left[\begin{matrix}I\otimes G+G\otimes I&0\\ 0&I\otimes G+G\otimes I\end{matrix}\right]\in{\mathbb{R}}^{2l^{2}\times 2l^{2}},\,\,B=\left[\begin{matrix}I\otimes F&F\otimes I\end{matrix}\right]\in{\mathbb{R}}^{l^{2}\times 2l^{2}},$ $C=E\otimes F\in{\mathbb{R}}^{l^{2}\times l^{2}},$ where $G=\frac{1}{(l+1)^{2}}\,\mathrm{tridiag}(-1,2,-1)\in{\mathbb{R}}^{l\times l},\quad F=\frac{1}{l+1}\,\mathrm{tridiag}(0,1,-1)\in{\mathbb{R}}^{l\times l}$ and $E=\mathrm{diag}(1,l+1,\ldots,l^{2}-l+1)\in{\mathbb{R}}^{l\times l}.$ Here, $E\otimes F$ denotes the Kronecker product of two matrices $E$ and $F$ and $\mathrm{tridiag}(a,b,c)\in{\mathbb{R}}^{l\times l}$ denotes the $l\times l$ tridiagonal matrix with diagonal entries $b$, subdiagonal entries $a$, and superdiagonal entries $c$.
+For this problem, the size of the matrix $\mathcal{A}$ is $4l^{2}.$
+Parameter selection: Following [8], selection of $\widehat{A}$ and $\widehat{S}$ (SPD approximations of $A$ and $S,$ respectively) for IBD preconditioner are done as follows:
 $\widehat{A}=LL^{T},\quad\widehat{S}=\mathrm{diag}(B\widehat{A}^{-1}B^{T}),$
@@
 […]
-The test problems generated for the values of $l=16,32,48,64,80,$ and their numerical results are reported in Table 2.
+For the preconditioner ${\mathscr{P}}_{\textit{MAPSS}},$ we take $\alpha=\sqrt[4]{\frac{tr(BB^{T}C^{T}C)}{m}}$ and $\beta=10^{-4}$ as in [21].
+We consider the parameter choices for the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS},{\mathscr{P}}_{EGSS},$ ${{\mathscr{P}}_{RPGSS}},$ ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}$ in the following two cases.
+  • In Case I: $\alpha=0.1$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=0.1,\beta=1,\gamma=0.001$ and $P=I,Q=I,W=I$ for ${\mathscr{P}}_{EGSS}$ and ${\mathscr{P}}_{RPGSS}$; and $\Lambda_{1}=I,\Lambda_{2}=I,\Lambda_{3}=0.001I$ for ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}.$
+  • In Case II: $\alpha=1$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=1,\beta=1,\gamma=0.001$ and $P=A,Q=I,W=CC^{T}$ for ${\mathscr{P}}_{EGSS}$ and ${\mathscr{P}}_{RPGSS}$; and $\Lambda_{1}=A,\Lambda_{2}=I,\Lambda_{3}=0.001CC^{T}$ for ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}.$
+The parameter selection in Case II is made as in [39].
+Results for experimentally found optimal parameter: The optimal value (experimentally) for $s$ in the range $[10,20]$ is determined to be $12$ for minimal CPU times.
+The test problems generated for the values of $l=16,32,48,64,80,{128},$ and their numerical results are reported in Table 2.
+Table 3: Numerical results of PESS-I, LPESS-I, PESS-II and LPESS-II PGMRES processes for Example 1.
+Process $l$ $16$ $32$ $48$ $64$ $80$ $128$
+size($\mathcal{A}$) $1024$ $4096$ $9216$ $16384$ $25600$ $65536$
+PESS-I IT $2$ $2$ $2$ $2$ $2$ $2$
+($s=1,$ $\Lambda_{1}=0.01I,$ CPU $0.2517$ $0.8373$ $4.6626$ $21.8133$ $77.9420$ $308.1418$
+$\Lambda_{2}=0.1I,$ $\Lambda_{3}=0.001I$) RES $074.4970e-$ 07 $073.8763e-$ 07 $073.6261e-$ 07 $073.4880e-$ 07 $073.3983e-$ 07 $077.7830e-$ 07
+LPESS-I IT $2$ $2$ $2$ $2$ $2$ $2$
+($s=1,$ $\Lambda_{2}=0.1I,$ CPU $0.3070$ $0.8184$ $4.7244$ $21.6534$ $82.3687$ $288.9701$
+$\Lambda_{3}=0.001I$ ) RES $073.1116e-$ 07 $072.2410e-$ 07 $071.8435e-$ 07 $071.6030e-$ 07 $071.4375e-$ 07 $071.1441e-$ 07
+PESS-II IT $3$ $3$ $3$ $3$ $3$ $3$
+($s=s_{est},$ $\Lambda_{1}=A,$ CPU $0.20373$ $1.05327$ $5.84150$ $29.3914$ $104.9663$ $424.2004336$
+$\Lambda_{2}=\beta_{est}I,$ $\Lambda_{3}=10^{-4}CC^{T}$) RES $082.2013e-$ 08 $083.4850e-$ 08 $084.3689e-$ 08 $085.0705e-$ 08 $085.8005e-$ 08 $087.2906e-$ 08
+LPESS-II IT $3$ $3$ $3$ $3$ $3$ $3$
+($s=s_{est},$ $\Lambda_{2}=\beta_{est}I,$ CPU $0.19215$ $1.04964$ $4.63143$ $24.20924$ $102.84651$ $389.10761$
+$\Lambda_{3}=10^{-4}CC^{T}$) RES $091.8126e-$ 09 $131.0207e-$ 13 $158.0620e-$ 15 $149.1601e-$ 14 $135.7292e-$ 13 $071.1950e-$ 07
+Results using parameters selection strategy in Section 6: To demonstrate the effectiveness of the proposed preconditioners PESS and LPESS using the parameters discussed in Section 6, we present the numerical results by choosing $s=1,$ $\Lambda_{1}=0.01I,\Lambda_{2}=0.1I,\Lambda_{3}=0.001I$ (denoted by PESS-I and LPESS-I) and $s=s_{est},$ $\Lambda_{1}=A,$ $\Lambda_{2}=\beta_{est}I$ and $\Lambda_{3}=10^{-4}C^{T}C$ (denoted by PESS-II and LPESS-II) for the proposed preconditioner.
+These results are summarized in Table 3.
+Convergence curves: Figure 1 illustrates convergence curves pertaining to preconditioners BD, IBD, MAPSS, SL, SS, RSS, EGSS, RPGSS, PESS and LPESS ($s=12$) for Case II with $l=16,32,48,64,80$ and $128$.
+These curves depict the relationship between the relative residue at each iteration step (RES) and IT counts.
+Spectral distributions: To further illustrate the superiority of the PESS preconditioner, spectral distributions of $\mathcal{A}$ and the preconditioned matrices ${\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},$ ${{\mathscr{P}}_{MAPSS}^{-1}\mathcal{A}},$ ${{\mathscr{P}}_{SL}^{-1}\mathcal{A}},$ ${\mathscr{P}}_{SS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{RSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A},$ ${{\mathscr{P}}_{RPGSS}^{-1}\mathcal{A}},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for the Case II with $l=16$ and $s=13$ are displayed in Figure 2.
 (a) $l=16$
@@
 (e) $l=80$
-Figure 1: Convergence curves for IT versus RES of PGMRES process employing BD, IBD, SS, RSS, EGSS, PESS and LPESS $(s=12)$ preconditioners in Case II for Example 1.
+(f) $l=128$
+Figure 1: Convergence curves for IT versus RES of PGMRES processes employing BD, IBD, MAPSS, SL, SS, RSS, EGSS, RPGSS, PESS and LPESS $(s=12)$ preconditioners in Case II for Example 1.
 (a) $\mathcal{A}$
@@
 (c) ${\mathscr{P}}_{IBD}^{-1}\mathcal{A}$
-(d) ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$
-(e) ${\mathscr{P}}_{RSS}^{-1}\mathcal{A}$
-(f) ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$
-(g) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=13$
-(h) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=13$
-Figure 2: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with $l=16$ for Example 1.
-Figure 3: Parameter $s$ vs condition number of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ for Example 1.
-Figure 1 illustrates convergence curves pertaining to preconditioners BD, IBD, SS, RSS, EGSS, PESS and LPESS ($s=12$) for Case II with $l=16,32,48,64$ and $80$ These curves depict the relationship between the relative residue at each iteration step (RES) and IT counts.
-To further illustrate the superiority of the PESS preconditioner, spectral distributions of $\mathcal{A}$ and the preconditioned matrices ${\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{RSS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for the Case II with $l=16$ and $s=13$ are displayed in Figure 2.
-To investigate the robustness of the proposed PESS preconditioner, we measure the condition number of the ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ which is for any nonsingular matrix $A$ defined by $\kappa(A):=\|A^{-1}\|_{2}\|A\|_{2}.$ In Figure 3, the influence of the parameter $s$ on the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ is depicted.
+(d) ${\mathscr{P}}_{MAPSS}^{-1}\mathcal{A}$
+(e) ${\mathscr{P}}_{SL}^{-1}\mathcal{A}$
+(f) ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$
+(g) ${\mathscr{P}}_{RSS}^{-1}\mathcal{A}$
+(h) ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$
+(i) ${\mathscr{P}}_{RPGSS}^{-1}\mathcal{A}$
+(j) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=13$
+(k) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=13$
+Figure 2: Spectral distributions of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{{\mathscr{P}}_{MAPSS}^{-1}\mathcal{A},{\mathscr{P}}_{SL}^{-1}\mathcal{A},}{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{{\mathscr{P}}_{RPGSS}^{-1}\mathcal{A},}{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with $l=16$ for Example 1.
+(a) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=13$
+(b) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=13$
+Figure 3: Spectral bounds for ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with $l=16$ for Example 1.
+Spectral bounds: Furthermore, we draw the eigenvalue bounds of Theorems 4.1 and 4.2 for the PESS preconditioned matrix.
+The $|\lambda-1|=1$ of Theorem 4.1 is drawn by the green unit circle in Figure 3(a) and the points in blue color are the bounds of Theorem 4.2.
+To draw the eigenvalues bounds in Theorem 5.1 for LPESS preconditioned matrix, we define the following circles:
+$C_{1}:=\left\{\lambda\in\mathbb{C}:|\lambda|=\sqrt{\frac{\tilde{\theta}_{\max}}{1+s\vartheta_{\min}+s^{2}\tilde{\theta}_{\max}}}\right\},\leavevmode\nobreak\ C_{2}:=\left\{\lambda\in\mathbb{C}:|\lambda|=\frac{\vartheta_{\min}}{2+s\vartheta_{\min}}\right\},$
+$C_{3}:=\left\{\lambda\in\mathbb{C}:\left|\lambda-\frac{1}{s}\right|=\frac{\vartheta_{\min}}{2+s\vartheta_{\min}}\right\}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ C_{4}:=\left\{\lambda\in\mathbb{C}:\left|\lambda-\frac{1}{s}\right|=\frac{1}{s(1+s\sqrt{\tilde{\theta}_{\max}})}\right\}.$
+In Figure 3(b), $C_{1}$ is in red color, $C_{2}$ is in blue color, $C_{3}$ is in black color and $C_{4}$ is in green color.
+Additionally, the eigenvalue bounds in Theorems 4.2, 4.3 and 5.1 are also presented in Table 4.
+Table 4: Spectral bounds for ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for case II with $l=16$ for Example 1.
+Real eigenvalue $\lambda$ Non-real eigenvalue $\lambda$
+Bounds of Theorem 4.2 Bounds of Theorem 4.3
+${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ contained in $0.0667\leq|\lambda|\leq 0.0739$
+the interval $4.258\times 10^{-5}\leq\Re(\lambda/(1-s\lambda))\leq 0.5$
+$(0,0.071429]$ $|\Im(\lambda/(1-s\lambda))|\leq 31.6386$
+Bounds of Theorem 5.1 Bounds of Theorem 5.1
+${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ contained in $0.0285\leq|\lambda|\leq 0.07693$
+the interval $1.8666\times 10^{-4}\leq|\lambda-\frac{1}{s}|\leq 0.04384$
+$(0.0416,0.0769]$
+Condition number analysis: To investigate the robustness of the proposed PESS preconditioner, we measure the condition number of the ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ which is for any nonsingular matrix $A$ is defined by $\kappa(A):=\|A^{-1}\|_{2}\|A\|_{2}.$ In Figure 4, the influence of the parameter $s$ on the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ is depicted.
 The parameter $s$ is chosen from the interval $[5,50]$ with step size one.
-The data presented in Table 2 and Figures 1-3 allow us to make the following noteworthy observations:
+Discussions: The data presented in Tables 2 and 3 and Figures 1-4 allow us to make the following noteworthy observations:
   • From Table 2, it is observed that the GMRES has a very slow convergence speed.
 In both Cases I and II, our proposed PESS and LPESS preconditioners outperform all other compared existing preconditioners from the aspects of IT and CPU times.
-For example, in Case I with $l=80,$ our proposed PESS preconditioner is $76\%$, $43\%$, $28\%$, $27\%$, and $33\%$ more efficient than the existing BD, IBD, SS, RSS, and EGSS preconditioner, respectively.
-Moreover, LPESS preconditioners performs approximately $78\%,$ $48\%,$ $35\%,$ $34\%,$ and $39\%$ faster than BD, IBD, SS, RSS, and EGSS preconditioner, respectively.
-Similar patterns are noticed for $l=16,32,48$ and $64$.
-  • In Figure 1, it is evident that the PESS and LPESS preconditioners have a faster convergence speed than the other baseline preconditioners when applied to the PGMRES process for all $l=16,32,48,64$ and $80$
+For example, in Case I with $l=80,$ our proposed PESS preconditioner is $76\%$, $43\%$, $36\%,$ $22\%,$ $28\%$, $27\%,$ $33\%$ and $39\%$ more efficient than the existing BD, IBD, MAPSS, SL, SS, RSS, EGSS and RPGSS preconditioner, respectively.
+Moreover, LPESS preconditioners performs approximately $78\%,$ $48\%,$ $38\%,$ $29\%$, $35\%,$ $34\%,$ $39\%$ and $44\%$ faster than BD, IBD, MAPSS, ${\mathscr{P}}_{SL},$ SS, RSS, EGSS and RPGSS preconditioner, respectively.
+Similar patterns are noticed for $l=16,32,48,64$ and $128$.
+  • Comparing the results of Tables 2 and 3, we observe that in both cases, PESS and LPESS preconditioners outperform the existing baseline preconditioners and IT are the same as in the experimentally found optimal parameters.
+This shows that the parameter selection strategy in Section 6 is effective.
+  • In Figure 1, it is evident that the PESS and LPESS preconditioners have a faster convergence speed than the other baseline preconditioners when applied to the PGMRES process for all $l=16,32,48,64,80$ and $128$
+    Figure 4: Parameter $s$ vs condition number of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ for Example 1.
   • According to Figure 2, the spectrum of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LESS}^{-1}\mathcal{A}$ have better clustering properties than the baseline preconditioned matrices, consequently improves the computational efficiency of our proposed PESS and LPESS PGMRES processes.
-Moreover, the real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ are contained in the interval $(0,0.0714]$ and for ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ are in $(0,1.9991]$ and $(0,1],$ respectively, which are consistent with bounds of Theorem 4.2, Corollaries 3.1, and 3.2, respectively.
+Moreover, the real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ are contained in the interval $(0,0.071429]$ and for ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ are in $(0,1.9991]$ and $(0,1],$ respectively, which are consistent with bounds of Theorem 4.2, Corollaries 3.1 and 3.2, respectively.
 For non-real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ obtained bounds are consistent with Theorem 4.3.
-  • For $l=32,$ the computed condition numbers of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},$ and ${\mathscr{P}}_{IBD}^{-1}\mathcal{A}$ are $5.4289e+04,$ $9.5567e+09,$ and $9.5124e+09,$ respectively, which are very large.
-Whereas Figure 3 illustrate a decreasing trend in the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ (for instance, when $s=50$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=3.4221$) with increasing values of $s.$ This observation highlights that ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ emerges as well-conditioned, consequently the solution obtained by the proposed PESS PGMRES process is more reliable and robust.
-The above discussions affirm that the proposed PESS and LPESS preconditioner is efficient, robust, and better well-conditioned with respect to the baseline preconditioners.
-Figure 4: Relationship of norm error of solution with increasing noise percentage, employing proposed ${\mathscr{P}}_{PESS}$ for Case II with $s=12$ for $l=16,32,48,64$ and $80$ of Example 1.
-Sensitivity analysis of the solution by employing the proposed PESS preconditioner
-To study the sensitivity of the solution obtained by employing the proposed PESS PGMRES process to small perturbations on the input data matrices.
+Moreover, from Figure 3(b), we observe that eigenvalues of ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ are contained in $C_{2}\leq\lambda\leq C_{1}\cap C_{4}\leq\lambda\leq C_{4}$, which shows the consistency of the bounds in Theorem 5.1.
+  • For $l=32,$ the computed condition numbers of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A}$ ${\mathscr{P}}_{IBD}^{-1}\mathcal{A},$ ${{\mathscr{P}}_{MAPSS}^{-1}\mathcal{A}}$ and ${\mathscr{P}}_{SL}^{-1}\mathcal{A}$ are $5.4289e+04,$ $9.5567e+09,$ $9.5124e+09,$ $057.2548e+$ 05 and $4.2852e+09,$ respectively, which are very large.
+Whereas Figure 4 illustrates a decreasing trend in the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ (for instance, when $s=50$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=3.4221$) with increasing values of $s.$ This observation highlights that ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ emerges as well-conditioned, consequently, the solution obtained by the proposed PESS PGMRES process is more reliable and robust.
+The above discussions affirm that the proposed PESS and LPESS preconditioner is efficient, robust and better well-conditioned with respect to the baseline preconditioners.
+Figure 5: Relationship of norm error of solution with increasing noise percentage, employing proposed ${\mathscr{P}}_{PESS}$ for Case II with $s=12$ for $l=16,32,48,64$ and $80$ for Example 1.
+Sensitivity analysis of the solution by employing the proposed PESS preconditioner: To study the sensitivity of the solution obtained by employing the proposed PESS PGMRES process to small perturbations on the input data matrices.
 In the following, we consider the perturbed counterpart $(\mathcal{A}+\Delta\mathcal{A})\tilde{\textbf{u}}=\textbf{d}$ of the three-by-three block SPP (1.1), where the perturbation matrix $\Delta\mathcal{A}$ has same structure as of $\mathcal{A}.$ We construct the perturbation matrix $\Delta\mathcal{A}$ by adding noise to the block matrices $B$ and $C$ of $\mathcal{A}$ as follows:
@@
 where $N_{P}$ is the noise percentage and $std(B)$ and $std(C)$ are the standard deviation of $B$ and $C,$ respectively.
-Here, $randn(m,n)$ is the is an $m$-by-$n$ normally distributed random matrix, generated by the MATLAB command $randn$.
+Here, $randn(m,n)$ is an $m$-by-$n$ normally distributed random matrix generated by the MATLAB command $randn$.
 We perform the numerical test for $l=16,32,48,64$ and $80$ for Example 1 using the PESS PGMRES process (Case II, $s=12$).
 The calculated norm error $\|\tilde{\textbf{u}}-\textbf{u}\|_{2}$ among the solution of the perturbed system and the original system with increasing $N_{P}$ from $5\%$ to $40\%$ with step size $5\%$ are displayed in the Figure 5.
-We noticed that with growing $N_{P},$ the norm error of the solution remains consistently less than $10^{-8},$ which demonstrates the robustness of the proposed PESS preconditioner and the solution u is not sensitive to small perturbation on $\mathcal{A}$.
+We noticed that with growing $N_{P},$ the norm error of the solution remains consistently less than $10^{-8},$ which demonstrates the robustness of the proposed PESS preconditioner and the solution u is insensitive to small perturbation on $\mathcal{A}$.
 Example 2.
+Table 5: Numerical results of GMRES, IBD, MAPSS, SL, SS, RSS, EGSS, RPGSS, PESS and LPESS PGMRES processes for Example 2.
+Process h $1/8$ $1/16$ $1/32$ $1/64$ $1/128$
+size($\mathcal{A}$) $288$ $1088$ $4224$ $16640$ $66048$
+[…]
+CPU $0.2325$ $0.6960$ $124.7595$ $4545.5072$ $9396.6958$
+RES $078.0113e-$ 07 $078.8023e-$ 07 $079.0887e-$ 07 $064.0481e-$ 06 $068.3021e-$ 06
+IBD IT $34$ $35$ $25$ $23$ $21$
+CPU $\bf 0.0147$ $\bf 0.1166$ $1.9070$ $32.6213$ $1664.6991$
+RES $079.4111e-$ 07 $076.9932e-$ 07 $079.2544e-$ 07 $079.3340e-$ 07 $077.3550e-$ 07
+MAPSS IT $11$ $7$ $4$ $4$ $4$
+CPU $0.2061$ $0.3453$ $1.8717$ $26.6600$ $482.0390$
+RES $085.2642e-$ 08 $071.1366e-$ 07 $087.1116e-$ 08 $078.2780e-$ 07 $092.1189e-$ 09
+SL IT $6$ $6$ $5$ $5$ $4$
+CPU $0.2130$ $0.2929$ $1.8697$ $27.1124$ $516.3012$
+RES $076.6518e-$ 07 $081.1953e-$ 08 $089.5638e-$ 08 $112.1528e-$ 11 $102.9479e-$ 10
+Case I
+SS IT $5$ $6$ $9$ $11$ $11$
+CPU $0.3047$ $0.4422$ $2.4279$ $51.5825$ $1245.2019$
+RES $074.5523e-$ 07 $077.9331e-$ 07 $072.5307e-$ 07 $074.5327e-$ 07 $079.8497e-$ 07
+RSS IT $\bf 4$ $\bf 4$ $5$ $5$ $6$
+CPU $0.2044$ $0.3436$ $1.6047$ $27.6128$ $720.4234$
+RES $083.4853e-$ 08 $079.2155e-$ 07 $081.7469e-$ 08 $072.7261e-$ 07 $093.2621e-$ 09
+EGSS IT $7$ $9$ $11$ $15$ $5$
+CPU $0.2389$ $0.4051$ $4.4353$ $52.7804$ $744.5983$
+RES $071.6861e-$ 07 $084.0517e-$ 08 $071.3223e-$ 07 $071.774e-$ 07 $079.1207e-$ 07
+RPGSS IT $4$ $4$ $4$ $5$ $5$
+CPU $0.2105$ $0.3883$ $2.2981$ $28.1748$ $544.8828$
+RES $092.8700e-$ 09 $085.1131e-$ 08 $073.3646e-$ 07 $098.8663e-$ 09 $071.4981e-$ 07
+PESS IT $4$ $4$ $5$ $5$ $4$
+CPU $0.2068$ $0.3480$ $1.7633$ $25.2920$ $484.2796$
+$s=30$ RES $082.3499e-$ 08 $074.2761e-$ 07 $082.1247e-$ 08 $072.0987e-$ 07 $074.3859e-$ 07
+LPESS IT $\bf 3$ $\bf{3}$ 4 3 44{4}4
+CPU $\bf 0.1990$ $\bf 0.2578$ $\bf 1.4778$ $\bf 17.9730$ $457.7446$
+$s=30$ RES $071.6606e-$ 07 $078.3803e-$ 07 $104.9289e-$ 10 $075.2658e-$ 07 $083.3191e-$ 08
+Case II
+SS IT $29$ $38$ $60$ $52$ $55$
+CPU $0.2999$ $1.1144$ $13.4580$ $213.2381$ $4980.7387$
+RES $077.3560e-$ 07 $079.5790e-$ 07 $079.2515e-$ 07 $079.6425e-$ 07 $079.6948e-$ 07
+RSS IT $12$ $13$ $14$ $12$ $12$
+CPU $0.3797$ $0.6019$ $3.4491$ $50.3476$ $1361.8902$
+RES $073.2456e-$ 07 $075.3086e-$ 07 $076.6600e-$ 07 $074.2834e-$ 07 $071.2733e-$ 07
+EGSS IT $9$ $9$ $9$ $7$ $4$
+CPU $0.2459$ $0.4180$ $2.4320$ $27.9359$ $419.1294$
+RES $097.9001e-$ 09 $098.6623e-$ 09 $097.4644e-$ 09 $071.2457e-$ 07 $087.6469e-$ 08
+RPGSS IT $6$ $7$ $7$ $5$ $4$
+CPU $0.2154$ $0.4096$ $1.8893$ $23.2802$ $423.0121$
+RES $085.2642e-$ 08 $071.1366e-$ 07 $087.1116e-$ 08 $078.2780e-$ 07 $092.1189e-$ 09
+PESS IT $5$ $5$ $5$ $5$ $\bf 3$
+CPU $0.2091$ $0.4184$ $1.8118$ $25.8048$ $\bf 365.3128$
+$s=26$ RES $082.3801e-$ 08 $081.1965e-$ 08 $098.4923e-$ 09 $085.9356e-$ 08 $074.2039e-$ 07
+LPESS IT $4$ $5$ $\bf 5$ $\bf 5$ $\bf 3$
+CPU $0.2070$ $0.3662$ $\bf 1.5330$ $\bf 23.5997$ $\bf 359.1330$
+$s=26$ RES $072.4250e-$ 07 $128.6182e-$ 12 $115.3930e-$ 11 $108.1253e-$ 10 $075.8345e-$ 07
+[…]
+[…]
 […]
@@
 […]
-Table 3: Numerical results of GMRES, BD, IBD, SS, RSS, EGSS, PESS and LPESS PGMRES processes for Example 2.
-Process h $1/8$ $1/16$ $1/32$ $1/64$
-size($\mathcal{A}$) $288$ $1088$ $4224$ $16640$
-GMRES IT $158$ $548$ $2048$bf{--}- -
-CPU $0.2325$ $0.6960$ $124.7595$ $4545.5072$
-RES $078.0113e-$ 07 $078.8023e-$ 07 $079.0887e-$ 07 $064.0481e-$ 06
-IBD IT $34$ $35$ $25$ $23$
-CPU $\bf 0.0147$ $\bf 0.1166$ $1.9070$ $32.6213$
-RES $079.4111e-$ 07 $076.9932e-$ 07 $079.2544e-$ 07 $079.3340e-$ 07
-Case I
-SS IT $5$ $6$ $9$ $11$
-CPU $0.3047$ $0.4422$ $2.4279$ $51.5825$
-RES $074.5523e-$ 07 $077.9331e-$ 07 $072.5307e-$ 07 $074.5327e-$ 07
-RSS IT $\bf 4$ $\bf 4$ $5$ $5$
-CPU $0.2044$ $0.3436$ $1.6047$ $27.6128$
-RES $083.4853e-$ 08 $079.2155e-$ 07 $081.7469e-$ 08 $072.7261e-$ 07
-EGSS IT $7$ $9$ $11$ $15$
-CPU $0.2389$ $0.4051$ $4.4353$ $52.7804$
-RES $071.6861e-$ 07 $084.0517e-$ 08 $071.3223e-$ 07 $071.774e-$ 07
-PESS IT $4$ $4$ $5$ $5$
-CPU $0.2068$ $0.3480$ $1.7633$ $25.2920$
-$s=30$ RES $082.3499e-$ 08 $074.2761e-$ 07 $082.1247e-$ 08 $072.0987e-$ 07
-LPESS IT $\bf 3$ $\bf{3}$ 4 3
-CPU $\bf 0.1990$ $\bf 0.2578$ $\bf 1.4778$ $\bf 17.9730$
-$s=30$ RES $071.6606e-$ 07 $078.3803e-$ 07 $104.9289e-$ 10 $075.2658e-$ 07
-Case II
-SS IT $29$ $38$ $60$ $52$
-CPU $0.2999$ $1.1144$ $13.4580$ $213.2381$
-RES $077.3560e-$ 07 $079.5790e-$ 07 $079.2515e-$ 07 $079.6425e-$ 07
-RSS IT $12$ $13$ $14$ $12$
-CPU $0.3797$ $0.6019$ $3.4491$ $50.3476$
-RES $073.2456e-$ 07 $075.3086e-$ 07 $076.6600e-$ 07 $074.2834e-$ 07
-EGSS IT $9$ $9$ $9$ $7$
-CPU $0.2459$ $0.4180$ $2.4320$ $27.9359$
-RES $097.9001e-$ 09 $098.6623e-$ 09 $097.4644e-$ 09 $071.2457e-$ 07
-PESS IT $5$ $5$ $5$ $5$
-CPU $0.2091$ $0.4184$ $1.8118$ $25.8048$
-$s=26$ RES $082.3801e-$ 08 $081.1965e-$ 08 $098.4923e-$ 09 $085.9356e-$ 08
-LPESS IT $4$ $5$ $\bf 5$ $\bf 5$
-CPU $0.2070$ $0.3662$ $\bf 1.5330$ $\bf 23.5997$
-$s=26$ RES $072.4250e-$ 07 $128.6182e-$ 12 $115.3930e-$ 11 $108.1253e-$ 10
-[…]
-[…]
+Table 6: Numerical results of PESS-I, LPESS-I, PESS-II and LPESS-II PGMRES processes for Example 2.
+Process h $1/8$ $1/16$ $1/32$ $1/64$ $1/128$
+size($\mathcal{A}$) $288$ $1088$ $4224$ $16640$ $66048$
+PESS-I IT $5$ $5$ $6$ $6$ $5$
+($s=1,$ $\Lambda_{1}=0.001I,$ CPU $0.30867$ $0.46688$ $2.48254$ $46.52887$ $577.9420$
+$\Lambda_{2}=0.1I,$ $\Lambda_{3}=0.001I$) RES $081.4125e-$ 08 $071.3379e-$ 07 $083.8821e-$ 08 $072.6061e-$ 07 $079.7383e-$ 07
+LPESS-I IT $4$ $4$ $4$ $5$ $3$
+($s=1,$ $\Lambda_{2}=0.1I,$ CPU $0.1829$ $0.3319$ $1.2774$ $19.1090$ $349.6728$
+$\Lambda_{3}=0.001I$) RES $093.5507e-$ 09 $083.4257e-$ 08 $077.0375e-$ 07 $092.3073e-$ 09 $077.1453e-$ 07
+PESS-II IT $3$ $4$ $4$ $4$ $4$
+($s=s_{est},$ $\Lambda_{1}=A,$ CPU $0.18470$ $0.39536$ $1.32148$ $17.22840$ $485.01767$
+$\Lambda_{2}=\beta_{est}I,$ $\Lambda_{3}=10^{-4}CC^{T}$) RES $079.6100e-$ 07 $075.6520e-$ 07 $087.8750e-$ 08 $083.9903e-$ 08 $096.9694e-$ 09
+LPESS-II IT $6$ $6$ $5$ $5$ $3$
+($s=s_{est},$ $\Lambda_{2}=\beta_{est}I,$ CPU $0.1656$ $0.3777$ $1.4341$ $19.0658$ $377.2001$
+$\Lambda_{3}=10^{-4}CC^{T}$) RES $072.5227e-$ 07 $075.1622e-$ 07 $084.8924e-$ 08 $083.2751e-$ 08 $077.9950e-$ 07
+To generate the matrices $A\in{\mathbb{R}}^{n\times n}$ and $B\in{\mathbb{R}}^{m\times n}$ of the system (1.1), the discretization task of the Stokes equation (7.1) is accomplished by the IFISS software developed by Elman et al.
+[3].
+Following [3] with grid parameters $\textbf{h}=1/8,1/16,1/32,1/64,{1/128}$, we get $n=2(1+1/\bf{h})^{2}$ and $m=1/\textbf{h}^{2}.$ To make the system of equations in (1.1) not too ill-conditioned and not too sparse, we construct the block matrix $C=\left[\begin{matrix}\Pi&randn(p,m-p)\end{matrix}\right]$, where $\Pi=\mathrm{diag}(1,3,5,\ldots,2p-1)$ and $p=m-2.$ Additionally, to ensure the positive definiteness of the matrix $A$, we add $0.001I$ with $A.$
 (a) $\textbf{h}=1/8$
@@
 (d) $\textbf{h}=1/64$
-Figure 5: Convergence curves for IT versus RES of the PGMRES processes by employing IBD, SS, RSS, EGSS, PESS and LPESS preconditioners in Case II for Example 2.
-[…]
-Table 3 shows that the GMRES process has a very slow convergence speed and also does not converge for $\textbf{h}=1/64$ within $7000$ iterations.
-The proposed preconditioners require almost five times fewer iterations compared to the IBD preconditioner for convergence.
-Moreover, in both cases, the PESS and LPESS preconditioners outperform the SS, RSS, and EGSS preconditioners in terms of IT and CPU times.
-Notably, IT for the PESS preconditioner, in both cases, remains constant, whereas for SS and EGSS, IT increases as the size of $\mathcal{A}$ increases.
-Furthermore, in Case II with $l=1/64$, our proposed LPESS preconditioner is approximately $28\%$, $90\%$, $53\%$, and $16\%$ more efficient than the existing IBD, SS, RSS, and EGSS preconditioner, respectively.
-Similar patterns are observed for other values of $l.$
+(e) $\textbf{h}=1/128$
+Figure 6: Convergence curves for IT versus RES of the PGMRES processes by employing IBD, MAPSS, SL, SS, RSS, EGSS, RPGSS, PESS and LPESS preconditioners in Case II for Example 2.
 (a) $\mathcal{A}$
 (b) ${\mathscr{P}}_{IBD}^{-1}\mathcal{A}.$
-(c) ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$
-(d) ${\mathscr{P}}_{RSS}^{-1}\mathcal{A}$
-(e) ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$
-(f) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=26$
-(g) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=26$
-Figure 6: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf h}=1/8.$
-The convergence curves in Figure 5 demonstrate the rapid convergence of the proposed PESS and LPESS PGMRES processes compared to the baseline IBD, SS, RSS, and EGSS ones in terms of RES versus IT counts.
-Moreover, Figure 6 illustrates the spectral distributions of the original matrix $\mathcal{A},$ and the preconditioned matrices ${\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{RSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf{h}}=1/8.$ According to Figure 6, eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ demonstrate a superior clustering compared to the other preconditioned matrices, which leads to a favorable convergence speed for the proposed PESS and LPESS PGMRES processes.
+(c) ${\mathscr{P}}_{MAPSS}^{-1}\mathcal{A}$
+(d) ${\mathscr{P}}_{SL}^{-1}\mathcal{A}$
+(e) ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$
+(f) ${\mathscr{P}}_{RSS}^{-1}\mathcal{A}$
+(g) ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$
+(h) ${\mathscr{P}}_{RPGSS}^{-1}\mathcal{A}$
+(i) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=26$
+(j) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=26$
+Figure 7: Spectral distributions of $\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{{\mathscr{P}}_{MAPSS}^{-1}\mathcal{A},{\mathscr{P}}_{SL}^{-1}\mathcal{A},}{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{{\mathscr{P}}_{RPGSS}^{-1}\mathcal{A},}{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf h}=1/8.$
+[…]
+Parameter selection: Parameter choices for the proposed preconditioners are made in two cases as in Example 1.
+However, in Case I, we take $\alpha=0.01,\beta=0.1,$ $\Lambda_{1}=0.01I$ and $\Lambda_{2}=0.1I.$ The parameter $\alpha$ in Case I is taken as in [36] and the parameters in Case II are taken as in [39].
+For the IBD preconditioner, the matrices $\widehat{A}$ and $\widehat{S}$ and for MAPSS preconditioner $\alpha$ and $\beta$ are constructed as in Example 1.
+[…]
+(a) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=26$
+(b) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=26$
+Figure 8: Spectral bounds for ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf h}=1/8$ for Example 2.
+Results using parameters selection strategy in Section 6: To demonstrate the effectiveness of the proposed PESS and LPESS preconditioners using the parameters discussed in Section 6, we present the numerical test results for PESS-I, LPESS-I, PESS-II and LPESS-II in Table 6 as in Example 1.
+Comparing the results in Tables 5 and 6, we observe that the parameter selection strategy described in Section 6 is effective.
+Table 7: Spectral bounds for ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for case II with $l=16$ for Example 2
+Real eigenvalue $\lambda$ Non-real eigenvalue $\lambda$
+Bounds of Theorem 4.2 Bounds of Theorem 4.3
+${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ contained in $0.0334\leq|\lambda|\leq 0.375$
+the interval $3.3624\times 10^{-6}\leq\Re(\lambda/(1-s\lambda))\leq 0.5$
+$(0,0.0370]$ $|\Im(\lambda/(1-s\lambda))|\leq 31.6386$
+Bounds of Theorem 5.1 Bounds of Theorem 5.1
+${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ contained in $0.0030\leq|\lambda|\leq 0.0385$
+the interval $4.6722\times 10^{-5}\leq|\lambda-\frac{1}{s}|\leq 0.0355$
+$(0.00552,0.0270]$
+Convergence curves: The convergence curves in Figure 6 demonstrate the rapid convergence of the proposed PESS and LPESS PGMRES processes compared to the IBD, MAPSS, SL, SS, RSS, EGSS and RPGSS in terms of RES versus IT counts.
+Spectral distributions: Figure 7 illustrates the spectral distributions of the original matrix $\mathcal{A},$ and the preconditioned matrices ${\mathscr{P}}_{IBD}^{-1}\mathcal{A},$ ${{\mathscr{P}}_{MAPSS}^{-1}\mathcal{A},}$ ${{\mathscr{P}}_{SL}^{-1}\mathcal{A},}$ ${\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{RSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A},$ ${{\mathscr{P}}_{PRGSS}^{-1}\mathcal{A},}$ ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf{h}}=1/8.$ According to Figure 7, eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ demonstrate a superior clustering compared to the other preconditioned matrices, which leads to a favorable convergence speed for the proposed PESS and LPESS PGMRES processes.
+Spectral bounds: In Figure 8, we draw the spectral bounds of Theorems 4.1, 4.2 and 5.1 for PESS and LPESS preconditioned matrices.
+In Figure 8(a), $|\lambda-1|=1$ of Theorem 4.1 is drawn by the green unit circle, while the points in blue indicate the bounds from Theorem 4.2.
+Moreover, we draw the circles $C_{1}$ (in red), $C_{2}$ (in blue), $C_{3}$ (in black) and $C_{4}$ (in green) in Figure 8(b).
+We can observe that the eigenvalues of the preconditioned matrix ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ lie in the intersection of the annulus of the circles $C_{1}$ and $C_{2},$ and the annulus of the circles $C_{3}$ and $C_{4}.$ Additionally, the eigenvalue bounds derived in Theorems 4.2, 4.3 and 5.1 are listed in Table 7.
 (a) PESS PGMRES process
@@
 Figure 9: Characteristic curves for IT of the proposed PESS (left) and LPESS (right) PGMRES processes by varying $s$ in the interval $[1,100]$ with step size $1$ with $\textbf{h}=1/16$ in Case I for Example 2.
-Furthermore, the system (1.1) exhibits ill-conditioning nature with $\kappa(\mathcal{A})=5.0701e+05.$ While for Case I with $\mathbf{h}=1/16$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=1.3353$ and $\kappa({\mathscr{P}}_{LPESS}^{-1}\mathcal{A})=1.2056,$ indicating that the proposed PESS and LPESS preconditioned systems are well-conditioned, ensuring an efficient and robust solution.
-In addition, to demonstrate the relationship of the parameter $s$ and the convergence speed of the PESS and LPESS preconditioners, we plot graphs of IT counts by varying the parameters $s$ in the interval $[1,100]$ with step size one in Figure 7.
+Condition number analysis: Furthermore, the system (1.1) exhibits ill-conditioning nature with $\kappa(\mathcal{A})=5.0701e+05.$ While for Case I with $\mathbf{h}=1/16$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=1.3353$ and $\kappa({\mathscr{P}}_{LPESS}^{-1}\mathcal{A})=1.2056,$ indicating that the proposed PESS and LPESS preconditioned systems are well-conditioned, ensuring an efficient and robust solution.
+Relationship between $s$ and convergence speed: In addition, to demonstrate the relationship of the parameter $s$ and the convergence speed of the PESS and LPESS preconditioners, we plot graphs of IT counts by varying the parameters $s$ in the interval $[1,100]$ with step size one in Figure 9.
 We consider $\textbf{h}=1/16$ and other choices for $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ as in Case I.
 Figure 9 shows that, with the increasing value of $s,$ decreasing trend in the IT counts for both the PESS and LPESS preconditioner.
-Moreover, for $s>22,$ using LPESS preconditioner, we can solve the three-by-three $SPP$ mentioned only in $3$ iterations.
+Moreover, for $s>22,$ using LPESS preconditioner, we can solve this three-by-three block $SPP$ only in $3$ iterations.
+Example 3.
+[1, 8]
+Table 8: Numerical results of GMRES, IBD, MAPSS, SL, SS, RSS, EGSS, RPGSS, PESS and LPESS PGMRES processes for Example 3.
+Process Problem aug2dc aug3dc listwet12 yao
+size($\mathcal{A}$) $50400$ $8746$ $30004$ $6004$
+GMRES IT $370$ $78$ $2251$ $49$
+CPU $360.1613$ $14.2486$ $422.8209$ $13.9152$
+RES $079.9905e-$ 07 $089.2839e-$ 08 $089.9700e-$ 08 $089.7483e-$ 08
+IBD IT $5$ $5$ $28$ $25$
+CPU $278.7707$ $4.80051$ $278.8961$ $21.9511$
+RES $152.1187e-$ 15 $097.5842e-$ 09 $078.3961e-$ 07 $079.4491e-$ 07
+MAPSS IT $8$ $8$ $124$ $36$
+CPU $324.3828$ $7.6662$ $2078.9607$ $18.6853$
+RES $073.6981e-$ 07 $091.7312e-$ 09 $079.2628e-$ 07 $077.6716e-$ 07
+SL IT $65$ $14$ $15$ $37$
+CPU $2848.2117$ $14.4158$ $270.3901$ $36.7559$
+RES $088.4621e-$ 08 $082.0305e-$ 08 $088.2851e-$ 08 $099.8555e-$ 09
+SS IT $23$ $6$ $7$ $39$
+CPU $1547.7320$ $8.9176$ $200.8205$ $28.9668$
+RES $079.9399e-$ 07 $077.7120e-$ 07 $073.5756e-$ 07 $079.5393e-$ 07
+RSS IT $23$ $6$ $5$ $16$
+CPU $1774.9981$ $8.6199$ $161.8405$ $18.5057$
+RES $075.5282e-$ 07 $072.6941e-$ 07 $074.1147e-$ 07 $078.2434e-$ 07
+EGSS IT $5$ $5$ $4$ $6$
+CPU $350.3081$ $7.61206$ $124.0637$ $6.9551$
+RES $123.9934e-$ 12 $121.1711e-$ 12 $071.5556e-$ 07 $072.0227e-$ 07
+RPGSS IT $4$ $4$ $3$ $5$
+CPU $297.3382$ $6.16249$ $96.5956$ $5.0133$
+RES $093.2268e-$ 09 $102.8160e-$ 10 $087.1294e-$ 08 $089.2318e-$ 08
+PESS IT $3$ $3$ $2$ $3$
+CPU $249.0203$ $5.5716$ $68.5225$ $4.8778$
+$s=30$ RES $088.3682e-$ 08 $085.7233e-$ 08 $078.3159e-$ 07 $073.3146e-$ 07
+LPESS IT $3$ $3$ $2$ $3$
+CPU $237.0945$ $5.4662$ $67.0810$ $3.9688$
+$s=30$ RES $091.0780e-$ 09 $098.4676e-$ 09 $074.3037e-$ 07 $073.7858e-$ 07
+[…]
+[…]
+within the prescribed IT.
+Problem formulation: In this example, we consider the three-by-three block SPP (1.1) arising from the following quadratic programming problem:
+$\min_{x\in{\mathbb{R}}^{n},\,y\in{\mathbb{R}}^{p}}\frac{1}{2}x^{T}Ax+f^{T}x+h^{T}y$ (7.2)
+such $\text{such that}\leavevmode\nobreak\ \leavevmode\nobreak\ Bx+C^{T}y=b,$ (7.3)
+where $f\in{\mathbb{R}}^{n}$ and $h\in{\mathbb{R}}^{p}.$ To solve the minimization problem, we define the Lagrange function:
+$L(x,y,\lambda)=\frac{1}{2}x^{T}Ax+f^{T}x+h^{T}y+\lambda^{T}(Bx+C^{T}y-b),$ (7.4)
+where $\lambda\in{\mathbb{R}}^{m}.$ Then the Karush-Kuhn-Tucker [47] necessary conditions are as follows:
+$\nabla_{x}L(x,y,\lambda)=0,\leavevmode\nobreak\ \nabla_{y}L(x,y,\lambda)=0\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \nabla_{\lambda}L(x,y,\lambda)=0,$
+which can be reformulated as a three-by-three block SPP of the form (1.1).
+Table 9: Numerical results of PESS-I, LPESS-I, PESS-II and LPESS-II PGMRES processes for Example 3.
+Process Problem aug2dc aug3dc listwet12 yao
+size($\mathcal{A}$) $50400$ $8746$ $30004$ $6004$
+PESS-I IT $5$ $3$ $5$ $6$
+($s=1,$ $\Lambda_{1}=0.001I,$ CPU $520.7330$ $10.1460$ $310.5960$ $5.7449$
+$\Lambda_{2}=0.1I,$ $\Lambda_{3}=0.001I$) RES $085.4476e-$ 08 $081.1047e-$ 08 $071.5581e-$ 07 $072.9433e-$ 07
+LPESS-I IT $4$ $3$ $3$ $3$
+($s=1,$ $\Lambda_{2}=0.1I,$ CPU $337.9530$ $8.3194$ $118.2540$ $3.5354$
+$\Lambda_{3}=0.001I$) RES $095.4773e-$ 09 $072.8655e-$ 07 $082.0805e-$ 08 $071.5576e-$ 07
+PESS-II IT $3$ $3$ $2$ $3$
+($s=s_{est},$ $\Lambda_{1}=A,$ CPU $236.4489$ $8.0818$ $74.9301$ $3.7349$
+$\Lambda_{2}=\beta_{est}I,$ $\Lambda_{3}=10^{-4}CC^{T}$) RES $083.0982e-$ 08 $095.8697e-$ 09 $097.8956e-$ 09 $076.3723e-$ 07
+LPESS-II IT $3$ $3$ $2$ $2$
+($s=s_{est},$ $\Lambda_{2}=\beta_{est}I,$ CPU $282.3466$ $8.3194$ $114.7483$ $2.8330$
+$\Lambda_{3}=10^{-4}CC^{T}$) RES $093.7409e-$ 09 $091.0390e-$ 09 $081.0465e-$ 08 $086.8900e-$ 08
+Data matrices $A$, $B$ and $C$ are selected from the CUTEr collection [48]. The numerical results for the problems ‘aud2dc’, ‘aud3dc’, ‘liswet12’ and ‘yao’ are presented in Table 8.
+Parameter selection: The parameter selection for the IBD and MAPSS preconditioners are as in Example 1.
+For the SS, RSS, EGSS, RPGSS, PESS and LPESS preconditioners, the parameters are chosen the same as in Case II of Example 1 with $\alpha=0.1$ and $\Lambda_{1}=0.1I$.
+Results for experimentally found optimal parameter: For the proposed preconditioner, the experimentally found parameter $s$ is $30$ in the interval $[10,30].$ The numerical results presented in Table 8 demonstrate that the proposed PESS and LPESS preconditioners significantly outperform existing baseline preconditioners in terms of IT and CPU time.
+The proposed preconditioners can solve the there-by-three block SPP within $2$ or $3$ iterations.
+Furthermore, for the problem ‘liswet12’, our proposed LPESS preconditioner is approximately $76\%$, $97\%$, $75\%,$ $66\%,$ $59\%,$ $46\%$ and $31\%$ more efficient than the existing IBD, MAPSS, SL, SS, RSS, EGSS and RPGSS preconditioner, respectively.
+Results using parameters selection strategy in Section 6: To further illustrate the effectiveness of the proposed preconditioners under the parameter selection strategy discussed in Section 6, we reported the numerical results in Table 9 for the preconditioners PESS-I, LPESS-I, PESS-II and LPESS-II (as discussed in Example $1$).
+A comparison of the results in Table 9 with the existing baseline preconditioners listed in Table 8 indicates that the proposed preconditioners perform well when the parameter $s$ is selected as outlined in Section 6.
 8 Conclusion
-In this paper, we proposed the PESS iterative process and corresponding PESS preconditioner ${\mathscr{P}}_{PESS}$ and its relaxed variant LPESS preconditioner to solve the three-by-three block SPP (1.1).
+In this paper, we proposed the PESS iterative process and corresponding PESS preconditioner and its relaxed variant LPESS preconditioner to solve the three-by-three block SPP (1.1).
 For the convergence of the proposed PESS iterative process, necessary and sufficient criteria are derived.
@@
 The key observations are as follows: $(i)$ the proposed PESS and LPESS preconditioners are found to outperform the existing baseline preconditioners in terms of IT and CPU times.
-$(ii)$ The proposed preconditioners significantly reduces the condition number of $\mathcal{A}$, consequently showing the its proficiency in solving three-by-three block SPPs.
-$(iii)$ The PESS and LPESS preconditioned matrices have a more clustered spectral distribution than the baseline preconditioned matrices.
+$(ii)$ The proposed preconditioners significantly reduces the condition number of $\mathcal{A}$, consequently showing their proficiency in solving three-by-three block SPPs.
+$(iii)$ The proposed PESS and LPESS preconditioned matrices have better clustered spectral distribution than the baseline preconditioned matrices.
 $(iv)$ Sensitivity analysis conducted by introducing different percentages of noise on the system (1.1) showcases the robustness of the proposed PESS preconditioner.
-Further investigation to determine the optimal parameter $s$ can be done in future research endeavors.
 Declaration of competing interest